### v1

# Language Model Preference Evaluation with Multiple Weak Evaluators

Abstract
Despite the remarkable success of Large Language Models (LLMs), evaluating their outputs’ quality regarding preference remains a critical challenge. Existing works usually leverage a powerful LLM (e.g., GPT4) as the judge for comparing LLMs’ output pairwisely, yet such model-based evaluator is vulnerable to conflicting preference, i.e., output A is better than B, B than C, but C than A, causing contradictory evaluation results. To improve model-based preference evaluation, we introduce GED (Preference Graph Ensemble and Denoise), a novel approach that leverages multiple model-based evaluators to construct preference graphs, and then ensemble and denoise these graphs for better, non-contradictory evaluation results. In particular, our method consists of two primary stages: aggregating evaluations into a unified graph and applying a denoising process to eliminate cyclic inconsistencies, ensuring a directed acyclic graph (DAG) structure. We provide theoretical guarantees for our framework, demonstrating its efficacy in recovering the ground truth preference structure. Extensive experiments across ten benchmark datasets show that GED outperforms baseline methods in model ranking, response selection, and model alignment tasks. Notably, GED combines weaker evaluators like Llama3-8B, Mistral-7B, and Qwen2-7B to surpass the performance of stronger evaluators like Qwen2-72B, highlighting its ability to enhance evaluation reliability and improve model performance.
1 Introduction
Large Language Models (LLMs) have rapidly transformed various fields within artificial intelligence, particularly natural language processing (NLP) and decision-making systems (Wu et al., 2023; Li et al., 2023a). Despite the remarkable success of LLMs, the need for effective evaluation methods becomes paramount (Liu et al., 2023; Desmond et al., 2024; Siska et al., 2024). Preference evaluation, as one of the most important assessment methods, plays an indispensable role in evaluating and optimizing model performance (Rafailov et al., 2024; Yuan et al., 2024; Dubois et al., 2024b). Existing works usually leverage a powerful LLM (e.g., GPT4 (Achiam et al., 2023)) as the judge for comparing LLMs’ output pairwisely (Li et al., 2023b; Chen et al., 2023; Li et al., 2023b; Wang et al., 2022).
Figure 1: The figure is divided into two parts. The upper part shows the construction of a preference graph from pairwise LLM evaluations, with cycles indicating inconsistent preferences highlighted in red. The lower part presents the evaluation of ten responses for HumanEval (Chen et al., 2021) and MATH (Hendrycks et al., 2021) tasks, generated by Llama3-70B (AI@Meta, 2024). These were evaluated by GPT-4-o, GPT-4-o-mini, GPT-3.5 (Achiam et al., 2023), Qwen2-72B (Yang et al., 2024a), and Llama3-8B, showing that even advanced LLMs exhibit significant noise in their evaluations, illustrating their limitations as reliable evaluators.
However, while such model-based pairwise preference evaluations offer a flexible approach, they can lead to contradictory evaluations in the assessment process (Zhang et al., 2024b). For example, an LLM might evaluate three responses and conclude that Response A is better than Response B (A $\succ$ B), Response B is better than Response C (B $\succ$ C), yet paradoxically also rank Response C as better than Response A (C $\succ$ A). These cyclic patterns introduce inconsistencies that undermine the reliability of the evaluation results. We model this conflicting preference via the preference graph. Specifically, a preference graph is constructed with each response as a node and directed edges indicating pairwise preferences—an edge from node A to node B shows that the evaluator preferred response A over B. The noise illustrated by cycles (A $\succ$ B $\succ$ C $\succ$ A) manifests as loops in the preference graph. The whole process we show in the upper part of Figure 1. Ideally, a preference graph should be structured as a directed acyclic graph (DAG) to maintain consistency. As shown in the bottom half of Figure  1, even advanced LLMs like GPT-4-o exhibit significant noise in preference evaluation, highlighting their limitations as weak evaluators.
To address this, we propose a novel framework, GED (Preference Graph Ensemble and Denoise), to address the inconsistencies in preference graphs generated through pairwise evaluations. Our method involves two key steps: (1) ensembling multiple weak evaluators to mitigate noise introduced by individual evaluators and (2) applying a denoising process to the resulting preference graph. By aggregating evaluations from multiple weak evaluators, we ”average out” the noise and biases, resulting in a more robust approximation of the true preference structure. The denoising step further refines this aggregated graph by removing inconsistencies, ensuring the final preference graph is more reliable for downstream tasks. We provide a theoretical analysis demonstrating the soundness of GED, showing that by treating each individual preference graph as a random perturbation of a ground truth DAG, our ensemble and denoising framework can recover the ground truth DAG with high probability.
To validate the practical efficacy of GED, we conduct extensive experiments across model ranking, response selection, and model alignment tasks, utilizing ten widely recognized benchmark datasets, including HumanEval (Chen et al., 2021), AlpacaEval (Li et al., 2023b), MATH (Hendrycks et al., 2021), GSM8k (Chen et al., 2021), GAIA (Mialon et al., 2023), LIMA (Zhou et al., 2023), Vicuna (Chiang et al., 2023), Koala (Vu et al., 2023), WizardLM (Xu et al., 2023), and Self-Instruct (Wang et al., 2022). In these experiments, GED consistently outperformed baseline methods. For example, in the response selection task, applying GED yielded an average improvement of 4.51% compared to baseline methods across multiple benchmarks. Additionally, GED demonstrated substantial gains in scenarios where combining preference graphs from weaker evaluators surpassed the performance of even stronger individual evaluators. For instance, by using Llama3-8B, Mistral-7B, and Qwen2-7B as weaker evaluators, GED exceeded the performance of using the stronger Qwen2-72B model as an evaluator. These results highlight GED’s ability to mitigate preference noise, improve consistency, and enhance model performance across diverse evaluation settings.
2 Methodologies
In this section, we begin by defining a preference graph, which serves as the foundation for representing pairwise preferences among candidates (Section 2.1). Building on this foundation, we introduce GED structured into three key stages (Section 2.2): (1) graph ensemble, where we aggregate individual preference graphs into a unified structure, (2) graph denoising, which removes cycles and inconsistencies to ensure the preference graph is acyclic, and (3) graph-to-ranking, where we extract a reliable ranking of candidates from the denoised graph. Below, we provide detailed descriptions of each step.
2.1 Preference Graph
A preference graph is defined as a directed graph $G_{P}=(V,A,w)$, where $V=\{v_{1},v_{2},\ldots,v_{n}\}$ represents a set of $n$ alternatives or candidates, $A\subseteq V\times V$ is a set of directed arcs representing pairwise preferences between these alternatives, and $w:A\rightarrow\mathbb{R}^{+}$ is a weight function that assigns a positive real value to each arc, indicating the strength of the preference.
For any pair of distinct vertices $u,v\in V$, an arc $(u,v)\in A$ exists if there is a preference for $u$ over $v$, with the weight $w(u,v)$ reflecting the intensity of this preference. Formally, this can be represented as:
$(u,v)\in A\quad\text{if and only if}\quad w(u,v)>0$ (1)
The weight function $w(u,v)$ aggregates individual preferences or scores for the pair $(u,v)$. If multiple preference sources exist, the weight can be expressed as:
$w(u,v)=\sum_{i=1}^{k}(s_{i}(u,v)-s_{i}(v,u))$ (2)
where $s_{i}(u,v)$ is the score or preference from the $i$-th source. The preference graph encapsulates the aggregate preferences among all pairs of alternatives, with the weight of each arc representing the cumulative preference strength derived from underlying data or models.
Figure 2: The GED framework mitigates noise and inconsistencies in preference evaluations by utilizing multiple weak evaluators and applying a graph denoising process. It consists of three stages: (1) Graph Ensemble, which combines individual preference graphs; (2) Graph Denoising, which removes cycles to ensure an acyclic structure; and (3) Graph-to-Ranking, which extracts a reliable ranking from the denoised graph.
2.2 GED: Preference Graph Ensemble and Denoise
As illustrated in Figure 2, our method, GED (Preference Graph Ensemble and Denoise), begins by performing graph ensemble to aggregate a set of preference graphs. It then applies graph denoising to ensure acyclicity, followed by graph-to-ranking to derive the final node ranking. The detailed steps are as follows:
Graph ensemble.
In the context of graph ensemble, given multiple weighted graphs $G_{1}=(V,A_{1},w_{1}),G_{2}=(V,A_{2},w_{2}),\ldots,G_{k}=(V,A_{k},w_{k})$ that share the same set of vertices $V$ but may differ in their arc sets $A_{i}$, the goal is to combine these graphs into a single ensemble graph $G_{E}=(V,A_{E},w_{E})$. The ensemble graph $G_{E}$ is formed by first defining the arc set $A_{E}$ as the union of the arc sets of the individual graphs and the weight function $w_{E}:A_{E}\rightarrow\mathbb{R}^{+}$ for the ensemble graph is then determined by summing the weights of the corresponding arcs in each graph.
Graph denoising.
Graph denoising involves transforming the original graph $G=(V,A,w)$ into a DAG. This transformation is achieved by identifying and removing a set of arcs known as the Feedback Arc Set (FAS) (Gabow, 1995), which is a set of arcs whose removal makes the graph acyclic. The goal is to find a minimum FAS, denoted as $R^{*}(G)$, which is a set of arcs with the smallest total weights that needs to be removed to eliminate all cycles in $G$.
To find this minimum FAS, we can order the vertices of $G$ in a specific sequence $s=\{v_{1},v_{2},\ldots,v_{n}\}$. This vertex sequence induces a FAS $R(s)$, consisting of all arcs that point against the direction of the sequence, i.e., arcs $v_{j}\rightarrow v_{i}$ where $j>i$. The graph denoising problem is thus reframed as finding an optimal vertex sequence $s^{*}$ that induces the minimal FAS, such that $R(s^{*})=R^{*}(G)$. This optimal sequence $s^{*}$ ensures that the total weights of arcs eliminated to achieve a DAG is minimized.
Finding a minimum FAS in general is known to be an NP-complete problem, whose computational complexity can be exponential (Karp, 2010; Bodlaender et al., 2012). Therefore, in our experiment, we apply the well-established approximation algorithm proposed in Eades et al. (1993). Details can be found in Appendix D.
Graph to ranking.
Given a DAG graph $G=(V,A,w)$, our goal is to derive a ranking based on the structure of $G$. For each vertex $v\in V$, we compute the descendant count $\text{desc}(v)$, defined as the number of vertices that are reachable from $v$ through directed arcs:
$\text{desc}(v)=\left|\{u\in V:v\rightsquigarrow u\}\right|,$ (3)
where $v\rightsquigarrow u$ denotes that there is a directed path from $v$ to $u$. Vertices are then ranked based on their descendant counts, with higher descendant counts indicating a higher position in the ranking. Formally, the ranking $\mathcal{R}$ is represented as a sequence of subsets:
=$\mathcal{R}=\{v_{1},v_{2},\ldots,v_{n}\},$ (4)
where each $v_{i}$ represents a set of vertices with the $i$-th highest descendant count. The final ranking is then:
$v_{1}\succ v_{2}\succ\cdots\succ v_{n}.$ (5)
This structured approach ensures that the ranking reflects not only the individual preferences captured in the graph but also the relative strength of these preferences as represented through their descendant connections.
2.3 Applications
We apply GED to three tasks: Response Selection, selecting the best response from LLM-generated candidates; Model Ranking, ranking models based on task performance; and Model Alignment, identifying the best instruction-response pairs for training. The steps are as follows:
Response selection.
In the response selection task, a model $\mathcal{M}$ generates $n$ candidate answers $\{ans_{1},\ldots,ans_{n}\}$ for each question $q\in Q$, with the objective of identifying the optimal answer $ans_{q}^{*}$ for each query. To achieve this, we employ multiple evaluators $\mathcal{A}=\{a_{1},\ldots,a_{k}\}$, who provide pairwise preferences among the candidate answers. For each question $q$, we construct a set of preference graphs $\{G_{a}:a\in\mathcal{A}\}$, where each graph $G_{a}=(V_{q},A_{a},w_{a})$ encapsulates the preferences of evaluator $a$. The vertex set $V_{q}=\{v_{1},\ldots,v_{n}\}$ corresponds to the candidate answers, while the directed arcs $A_{a}$ indicate the pairwise preferences among these responses. Each arc is weighted by $w_{a}$, reflecting the strength of the preference indicated by the evaluator. The construction of preference graphs involves evaluating each pair of candidate answers $ans_{i}$ and $ans_{j}$. Evaluators assess the quality of these answers, assigning a preference that is denoted by a directed arc $(v_{i},v_{j})$ in $A_{a}$, with a corresponding weight $w_{a}(v_{i},v_{j})$ based on the strength of preference. This process is detailed in Appendix B. After collecting the preference graphs $\{G_{a}:a\in\mathcal{A}\}$ for a question $q$, we apply GED to aggregate these graphs. The aggregation begins with merging all preference graphs into a unified graph $G_{q}=(V_{q},A_{q},w_{q})$, which is subsequently processed to remove cycles, resulting in a DAG. From this DAG, we derive the ranking $\mathcal{R}_{q}=\{v_{1}\succ v_{2}\succ\ldots\succ v_{n}\}$ of the candidate answers, where the highest-ranked answer is selected as $ans_{q}^{*}$. This process is repeated for each question $q\in Q$, yielding a final set of selected answers $ans^{*}=\{ans_{1}^{*},\ldots,ans_{t}^{*}\}$. This set reflects a consensus from multiple evaluators in $\mathcal{A}$, ensuring the chosen answers represent the highest quality responses based on rigorous preference evaluation.
Model Ranking.
In model ranking task, the goal is to rank a set of models $M=\{\mathcal{M}_{1},\ldots,\mathcal{M}_{n}\}$ based on their responses to a series of questions $Q=\{q_{1},\ldots,q_{t}\}$. A group of evaluators, denoted as $\mathcal{A}=\{a_{1},a_{2},\ldots,a_{k}\}$, assesses the model outputs by providing preferences for pairs of responses for each question. For each question $q\in Q$, the evaluators generate preference graphs $\{G_{a}:a\in\mathcal{A}\}$, where each graph $G_{a}=(V_{q},A_{a},w_{a})$ encapsulates the preferences of evaluator $a$ over the models. The vertex set $V_{q}=\{v_{1},\ldots,v_{n}\}$ corresponds to the models, while the directed arcs $A_{a}$ indicate pairwise preferences, with weights $w_{a}:A_{a}\rightarrow\mathbb{R}^{+}$ reflecting the strength of these preferences. The preference graph for a given question $q$ is constructed by evaluating each pair of models $\mathcal{M}_{i}$ and $\mathcal{M}_{j}$, represented by nodes $v_{i}$ and $v_{j}$. For their respective answers $ans_{i}=\mathcal{M}_{i}(q)$ and $ans_{j}=\mathcal{M}_{j}(q)$, evaluators provide a preference indicating which answer is favored. This preference is represented by a directed arc $(v_{i},v_{j})$ in $A_{a}$, assigned a weight $w_{a}(v_{i},v_{j})$ based on preference strength. The detailed procedure is outlined in Appendix B. Once the preference graphs $\{G_{a}:a\in\mathcal{A}\}$ are collected for a question $q$, we employ GED to aggregate these graphs. The method begins by merging all preference graphs into a single graph $G_{q}=(V_{q},A_{q},w_{q})$, which is then transformed into a DAG by removing cycles. The final ranking $\mathcal{R}_{q}=\{v_{1}\succ v_{2}\succ\ldots\succ v_{n}\}$ is derived from this DAG. This process is repeated for each question $q\in Q$, yielding a set of rankings $\{\mathcal{R}_{q}:q\in Q\}$. To compute the overall ranking $\mathcal{R}^{*}$ of the models across all questions, we conduct a ranking ensemble on the set $\{\mathcal{R}_{q}:q\in Q\}$, as detailed in Appendix E. This approach culminates in a final ranking that reflects the models’ performance as assessed by multiple evaluators in $\mathcal{A}$.
Model Alignment.
In the model alignment task, we have multiple data pairs of the form $(x,y_{1}),(x,y_{2}),\ldots,(x,y_{n})$ for each instruction $x$. The objective is to identify the best response $y^{*}$ corresponding to each instruction $x$. We utilize multiple evaluators $\mathcal{A}=\{a_{1},\ldots,a_{k}\}$ to provide pairwise preferences among the candidate responses. For each instruction $x$, we construct a set of preference graphs $\{G_{a}:a\in\mathcal{A}\}$, where each graph $G_{a}=(V_{x},A_{a},w_{a})$ represents the preferences of evaluator $a$. The vertices $V_{x}$ correspond to the candidate responses $\{y_{1},\ldots,y_{n}\}$, and the directed arcs in $A_{a}$ indicate preferences, weighted by $w_{a}$. After constructing the preference graphs, we apply GED to aggregate them into a single graph $G_{x}$. This graph undergoes a denoising process to remove cycles, allowing us to derive a ranking $\mathcal{R}_{x}$ of the responses. The highest-ranked response in $\mathcal{R}_{x}$ is selected as $y^{*}$ for that instruction $x$. This process is repeated for all $t$ instructions, resulting in the final training set $\{(x_{1},y_{1}^{*}),\ldots,(x_{t},y_{t}^{*})\}$, which reflects a consensus across the evaluators in $\mathcal{A}$.
3 Theoretical Analysis
In this section, we provide a theoretical foundation for our method, showing that by modeling preference graphs as random perturbations of a ground truth DAG, GED can reliably recover the true structure through graph ensemble and denoising with high probability, demonstrating its robustness in handling noisy evaluations.
Theoretically, we treat each of our preference graph as a random perturbation of some ground truth DAG $G=(V,A)$. Specifically, we consider a random graph generator $\mathcal{G}(G,\delta_{1},\delta_{2})$ with parameters $\delta_{1},\delta_{2}\in[0,1]$ such that $G_{i}=(V_{i},A_{i})\sim\mathcal{G}(G,\delta_{1},\delta_{2})$ satisfies $V_{i}=V$. Furthermore, for each $u,v\in V$ with $u\neq v$,
  1. 1)
    If $(u\rightarrow v)\in A$, then
    $\mathbb{P}((u\rightarrow v)\in A_{i})=1-\delta_{1}\quad\text{and}\quad\mathbb{P}((v\rightarrow u)\in A_{i})=\delta_{1};$
  2. 2)
    If $(u\rightarrow v),(v\rightarrow u)\notin A$, then
    $\mathbb{P}((u\rightarrow v),(v\rightarrow u)\notin A_{i})=1-\delta_{2},\quad\mathbb{P}((u\rightarrow v)\in A_{i})=\frac{\delta_{2}}{2}\quad\text{and}\quad\mathbb{P}((v\rightarrow u)\in A_{i})=\frac{\delta_{2}}{2}.$
That is, each edge in $E$ has probability $\delta_{1}$ of being flipped and each pair of unconnected nodes has probability $\delta_{2}$ of being connected with a random direction.
[…]
Theorem 1
[…]
$\mathbb{P}\left(G\subseteq\text{MAS}(\widehat{G})\right)\geq 1-2|A|\exp\left(-\frac{N\epsilon^{2}}{2}\right)-2U\exp\left(-\frac{N\epsilon^{2}}{6U^{2}\delta_{2}+2U\epsilon}\right),$
where $G\subseteq\text{MAS}(\widehat{G})$ represents that $G$ is a subgraph of $\text{MAS}(\widehat{G})$ and $U=\frac{|V|(|V|-1)}{2}-|A|$ is the number of pairs of unconnected nodes in $G$.
The full proof is given in Appendix C. From the theorem, we can see that the probability of failure decreases exponentially as the number of samples $N$ increases. Meanwhile, this guarantee only requires $\delta_{1}<0.5$ and does not place restrictions on $\delta_{2}$, which are very mild conditions.
4 Experiments on Response Selection
Table 1: Performance comparison of response selection methods across five benchmarks. GED consistently outperforms baseline methods, demonstrating the effectiveness of graph denoising and the aggregation of weaker evaluators.
Method HumanEval AlpacaEval MATH GSM8k GAIA Avg
Llama3-8B 43.90 27.29 22.08 56.67 6.78 31.14
Mistral-7B 23.17 11.80 23.25 39.83 7.03 15.21
Qwen2-7B 48.58 25.71 59.92 76.75 7.70 43.17
Qwen2-72B 57.93 29.58 72.75 84.67 11.52 51.91
ContraSolver 65.42 31.12 74.95 86.84 12.22 54.11
ListPreference 61.52 31.67 71.75 85.0 10.90 52.16
Single model Self-consistency 60.98 29.33 73.58 84.91 8.86 51.13
Llama3-8B 62.19 29.31 74.27 83.16 11.31 52.87
with graph denoising 64.02 30.18 74.73 86.00 11.72 53.53
Mistral-7B 67.24 27.70 74.41 83.83 10.50 52.93
with graph denoising 68.73 29.93 74.77 83.91 10.74 53.21
Qwen2-7B 61.58 28.69 74.50 85.41 11.11 52.84
with graph denoising 65.85 29.44 74.79 86.38 11.25 53.54
Qwen2-72B 60.97 31.04 74.73 86.47 12.14 53.87
Single evaluator with graph denoising 68.90 31.17 75.33 87.45 12.26 55.42
GED (w/o denoising) 69.25 30.98 74.29 87.17 12.68 54.83
Multiple evaluator GED 70.73 32.44 75.58 88.18 13.33 56.41
Figure 3: Comparison of GED with GPT-3.5, GPT-4-o-mini, and GPT-4-o on 100 randomly selected tasks. GED consistently outperforms GPT-3.5 across all tasks and surpasses GPT-4-o-mini on challenging tasks like HumanEval and GSM8k, showcasing the effectiveness of weak evaluator aggregation with graph denoising.
Experiment Setup.
In this section, we evaluate the performance of GED on five benchmarks: HumanEval (Chen et al., 2021), AlpacaEval (Li et al., 2023b), MATH (Hendrycks et al., 2021), GSM8k (Chen et al., 2021), and GAIA (Mialon et al., 2023). The Qwen2-72B (Yang et al., 2024a) model ($\mathcal{M}$) generates ten candidate responses per question, and we assess the effectiveness of different methods in selecting the best response. For further implementation details, see Appendix A. We evaluate performance using three setups. First, in the single model setting, the baselines include ContraSolver(Zhang et al., 2024b), Self-consistency(Wang et al., 2022), and direct evaluation with models (Llama3-8B, Mistral-7B, Qwen2-7B and Qwen2-72B). Additionally, we include a baseline called ListPreference, where instead of pairwise comparisons, all candidate responses are input into Qwen2-72B for selecting the most appropriate response. Then, in the single evaluator setting, individual evaluators (Llama3-8B, Mistral-7B, Qwen2-7B, Qwen2-72B) select the best response from $\mathcal{M}$’s outputs, with and without applying GED’s graph denoising. Finally, in the multiple weak evaluators setup, we combine three weaker evaluators (Llama3-8B, Qwen2-7B, Mistral-7B) to select responses from Qwen2-72B with GED. We present the results of GED and its variant (w/o denoising), which ensembles the preference graphs without the denoising step.
Main results.
Table 1 presents the results of the response selection task across five benchmarks. GED consistently outperforms baseline methods, including both single model evaluations (single model) and direct response selection by individual models (single evaluator). This demonstrates the strength of aggregating weak evaluators with GED, particularly when coupled with graph denoising, which enhances response quality by filtering out noise and biases. This highlights the effectiveness of aggregating weak evaluators and applying graph denoising to improve response quality. Furthermore, by combining preference graphs from weaker models (Llama3-8B, Mistral-7B, Qwen2-7B), GED surpasses the performance of a much stronger evaluator (Qwen2-72B). This underscores the value of ensemble methods in mitigating the limitations of individual evaluators. Then, the denoising process proves to be crucial for improving consistency and overall response quality. The substantial performance gains observed when using GED with denoising, compared to both the single evaluator setup and the ensemble without denoising, highlight its importance in refining response selection. Additionally, we observed that the ListPreference baseline performed worse than Qwen2-72B as single evaluator, likely due to LLM limitations in handling long-text. Lastly, to further evaluate GED, we compared its performance with GPT-3.5, GPT-4-o-mini, and GPT-4-o. Due to computational and API cost constraints, we limited the evaluation to 100 data points for each task. As shown in Figure 3, GED consistently outperformed GPT-3.5 across all tasks and surpassed GPT-4-o-mini on challenging benchmarks like HumanEval and GSM8k. These results highlight the superiority of GED, particularly in leveraging multi-weak evaluators and graph denoising to outperform individual state-of-the-art models.
Alabation study.
Figure 4: Comparison of GED and (w/o ensemble) variants. GED outperforms due to preserving more information by directly ensembling preference graphs, while rank aggregation in the (w/o ensemble) methods leads to performance loss.
We evaluate the impact of removing the ensembling step in GED, referred to as the (w/o ensemble) variant. In this case, individual evaluators’ preference graphs are denoised and converted to rankings, which are then aggregated using methods such as Weight Score, Kemeny, Weighted Kemeny, Pairwise Majority, and Weighted Pairwise Majority (detailed in Appendix E). For simplicity of presentation, we use Weight Score to represent GED (w/o ensemble) (Weight Score). As shown in Figure 4, all (w/o ensemble) methods consistently underperform compared to GED. This performance gap arises because converting graphs to ranks before aggregation leads to information loss. In contrast, GED ensembles the graphs directly, preserving more detailed preference information and resulting in better final rankings.
Table 2: Results of the model ranking task, evaluated using Ranking Correction. Higher correlation values indicate a stronger alignment with the ground truth rankings.
Model Weight Score Kemeny
Weighted
Kemeny
Pairwise
Majority
Weighted
Pairwise
Majority
Avg.
Llama3-70B 50.88 60.80 60.80 62.23 61.85 59.31
with graph denoising 52.44 62.54 62.54 63.92 62.18 60.72
Qwen2-72B 65.34 59.87 67.39 66.05 66.59 65.04
with graph denoising 66.05 70.43 70.43 72.32 72.41 70.32
Qwen1.5-72B 63.64 60.72 60.72 62.65 63.28 62.20
with graph denoising 64.81 61.77 61.77 64.36 64.76 63.49
Mistral-8×7B 64.90 68.74 68.74 73.06 72.87 69.66
Single evaluator with graph denoising 65.47 70.06 69.92 73.39 73.21 70.41
GED (w/o ensemble) 62.82 68.44 68.44 69.34 67.34 67.27
GED (w/o denoising) 64.84 69.23 69.81 75.35 74.37 70.72
Multiple evaluator GED 66.59 71.14 71.14 77.17 76.46 72.50
5 Experiments on Model Ranking
Experiment Setup.
In this section, we evaluate the effectiveness of GED in the model ranking task within a human preference setting, using the AlpacaEval benchmark (Li et al., 2023b). We employ 30 widely used models from the AlpacaEval dataset as our model set $\mathcal{M}$, while the benchmark’s questions form the question set $Q$. The rankings provided by the AlpacaEval benchmark serve as ground truth for evaluating the accuracy of various ranking methods. This is justified by AlpacaEval’s strong correlation with Chatbot Arena rankings, making it a reasonable proxy for human judgments (Dubois et al., 2024a). We adopt Ranking Correction, measured by the Spearman rank correlation coefficient, to evaluate the similarity. To generate rankings, we utilize outputs from the open-source models Llama3-70B, Qwen2-72B, Mistral-8$\times$7B, and Qwen1.5-72B as our evaluators, denoted as set $\mathcal{A}$. For further implementation details, see Appendix A. We investigate two variants of GED: (w/o ensemble) denoises the preference graphs from different evaluators for the same question, converts each into a ranking, and then ensembles these rankings to produce the final output, while (w/o denoising) directly ensembles the preference graphs to obtain the final ranking without denoising.
Main results.
The results, presented in Table 2, show that GED outperforms all single-model baselines, highlighting the significant improvement in ranking accuracy achieved by leveraging preference information from multiple evaluators. Moreover, GED surpasses the (w/o ensemble) variant, indicating that generating rankings through graph ensemble first prevents information loss compared to converting individual graphs into rankings. When the ensemble graph is not denoised (w/o denoising), residual noise can adversely affect the final ranking quality. Additionally, our denoising method also enhances results in single-model settings.
6 Experiments on Instruct Tuning
Experiment Setup.
In this section, we explore the effects of various data selection methods for model alignment on Llama-2-7B (Touvron et al., 2023) and Mistral-7B (Jiang et al., 2023) through instruct tuning. Specifically, we randomly sampled 5000 data points from UltraFeedback (Cui et al., 2023), and utilized Qwen1.5-14B (Yang et al., 2024a) to generate 8 responses for each data point as instruct data. We then applied four different methods—Random, Longest (Zhao et al., 2024), ContraSolver (Zhang et al., 2024b), and our proposed GED—to select a subset of these responses for model alignment training. The Origin refers to the performance of the base model without alignment. The resulting models were evaluated on the HH-RLHF (Bai et al., 2022) benchmark, which consists of four sub-sets: Harmless (base), Helpful (base), Helpful (online), and Helpful (rejection). For evaluation, we adopted the same Reward model, following prior work (Song et al., 2024; Yu et al., 2023), to measure human preference levels gained by the models. These results are summarized in Table 3. To ensure a comprehensive evaluation, we also tested the models, using the Llama-2-7B backbone, across additional benchmarks such as LIMA (Zhou et al., 2023), Vicuna (Chiang et al., 2023), Koala (Vu et al., 2023), WizardLM (Xu et al., 2023) and Self-Instruct (Wang et al., 2022), in line with recent works (Chen et al., 2023; Zhang et al., 2024a; Hu et al., 2024).
Table 3: Performance comparison of different methods (Random, Longest, ContraSolver, and GED) on model alignment task across the HH-RLHF benchmark. The results demonstrate the superiority of GED in consistently selecting high-quality responses, leading to improved model performance compared to baseline methods.
BaseModel Harmless (base) Helpful (base) Helpful (online) Helpful (rejection) Avg.
Origin 69.67 61.12 65.41 64.06 65.07
Random 69.38 62.87 66.75 65.57 66.14
Longest 69.65 63.54 66.99 66.43 66.65
ContraSolver 69.57 63.61 66.87 66.59 66.66
Llama-2-7B GED 69.71 64.10 67.87 67.01 67.17
Origin 61.59 59.51 65.21 63.17 62.37
Random 59.15 59.61 64.06 62.38 61.30
Longest 61.81 60.53 64.52 63.22 62.52
ContraSolver 61.48 59.85 64.66 63.41 62.85
Mistral-7B GED 61.96 60.71 65.49 63.82 63.50
Figure 5: Performance comparison of different methods (Random, Longest, ContraSolver, and GED) across multiple benchmarks, including LIMA, Vicuna, Koala, WizardLM and Self-Instruct. The results show GED effectively filters low-quality responses, improving performance and model alignment over baselines.
Main results.
From Table 3, we observe that GED consistently outperforms all baseline methods, demonstrating its effectiveness in selecting high-quality responses when multiple answers are available for a given instruction. When faced with multiple responses $y_{1},y_{2},\ldots,y_{n}$ for a given instruction $x$, the Random selection method can have a detrimental impact, especially when the quality of the responses is inconsistent. This effect is most evident with the Mistral-7B, where Random selection actually performs worse than the Origin, indicating that randomly chosen data points can introduce noise and degrade the model’s performance. Moreover, we find that simply selecting the longest response does not always lead to the best outcomes. While longer responses may provide more detailed answers, they are not necessarily better in terms of quality, particularly when both high-quality and low-quality answers exist for the same question. This is reflected in the results where the Longest method underperforms compared to both ContraSolver and GED, emphasizing that response length alone is not always a reliable criterion. From Figure 5, we can draw similar conclusions as Table 3. Specifically, we observe that GED consistently outperforms all baselines, demonstrating its effectiveness across all datasets. Particularly in AlpacaEval and Self-Instruct, the Random baseline performs worse than the Origin model, indicating that when response quality varies significantly, poor selection can lead to negative performance. In contrast, GED excels by aggregating preference graphs and applying denoising, effectively filtering out low-quality responses. This ensures robust performance, especially in cases where response quality is inconsistent. The denoising step is crucial for removing noisy evaluations, leading to improved model alignment. The denoising process in GED proves essential, particularly in settings with inconsistent responses, as it removes evaluation noise and leads to more robust performance.
7 Related Work
Preference evaluation of LLMs.
Reference-free evaluation metrics have a long history (Louis & Nenkova, 2013), which evaluates the generated text based on intrinsic properties and coherence with the context. Although they achieve high accuracy on matching inner-evaluator, the achievement suffers from spurious correlations such as perplexity and length (Durmus et al., 2022). Recently, people have started using a strong model (e.g., GPT-4) as an evaluator to perform a zero-shot reference-free evaluation on the weak models (Shen et al., 2023; Dubois et al., 2024b; Chen et al., 2023). However, using LLM-based preference evaluations can introduce inconsistencies in preference graphs, often resulting in cyclic preferences or contradictions when comparing multiple outputs.
Weak supervision.
The concept of weak-to-strong supervision originates from the need to leverage noisy or partial labels in machine learning tasks, enabling the development of more robust models from imperfect data (Ratner et al., 2016; Zhang et al., 2023b; 2022). In LLMs, weak-to-strong supervision aids AI alignment by allowing weaker models to improve strong ones, enhancing performance without extensive data and supporting scalable oversight (Zheng et al., 2024; Guo & Yang, 2024; Tong et al., 2024). Similarly, in task-oriented LLMs, weak-to-strong learning improves LLM’s ability by enabling strong models to refine their data autonomously, boosting performance without extensive high-quality input (Zhang et al., 2023a; Yang et al., 2024b). Through weak-to-strong supervision, LLM performance can be significantly improved by iteratively transforming low-quality labels into more reliable ones, leading to more effective model training and robust outputs (Zakershahrak & Ghodratnama, 2024; Lang et al., 2024).
8 Conclusion
In this paper, we presented GED, a framework designed to address inconsistencies in pairwise preference evaluations by LLMs. By employing graph ensemble techniques and denoising, GED reduces cyclic patterns and enhances the reliability of evaluation outcomes. Our theoretical analysis shows that GED can recover the ground truth DAG under reasonable conditions, improving consistency in preference rankings. Extensive experiments across response selection, model ranking, and instruct tuning demonstrate the efficacy of our method. GED consistently outperformed baseline methods in both single-evaluator and multi-evaluator settings, particularly in scenarios where combining weak evaluators led to superior results over stronger individual evaluators. Future work will explore extending GED to broader evaluation frameworks and applying its principles to more complex decision-making tasks, including multi-agent systems and human-AI interaction.
Appendix A Implementation Details
A.1 Experimental setup
All experimental procedures were conducted on a machine equipped with an AMD EPYC 7543 32-Core Processor, 512GB memory, 128 CPUs, and four 80GB NVIDIA A800 GPUs. The code is available at https://github.com/ppsmk388/GED. The references to Llama-2-7B, Llama3-70B, Llama3-8B, Mistral-7B, Qwen2-7B, and Qwen2-72B in the main text refer to the specific models: Llama-2-7b-chat-hf, Meta-Llama-3-70B-Instruct, Meta-Llama-3-8B-Instruct, Mixtral-8x7B-Instruct-v0.1, Qwen2-7B-Instruct, and Qwen2-72B-Instruct. We utilized the reward model oasst-rm-2-pythia-6.9b-epoch-1 following prior works (Song et al., 2024; Yu et al., 2023). Each experiment was repeated three times, and the average performance was reported as the final result. The training was configured with a batch size of 1 per device, gradient accumulation steps of 4, a learning rate of 1e-5, and he model was trained for 3 epochs, with warmup over 20 steps and a cosine learning rate scheduler. For tasks such as AlpacaEval (Dubois et al., 2024b), we used GPT-4-o unless stated otherwise.
A.2 Dataset
In this appendix, we provide detailed information about the datasets used in main text.
  • UltraFeedback (Cui et al., 2023): UltraFeedback is a large-scale, fine-grained, diverse preference dataset, used for training powerful reward models and critic models. We collect about 64k prompts from diverse resources (including UltraChat, ShareGPT, Evol-Instruct, TruthfulQA, FalseQA, and FLAN). We then use these prompts to query multiple LLMs (see Table for model lists) and generate 4 different responses for each prompt, resulting in a total of 256k samples.
  • HH-RLHF (Bai et al., 2022): The HH-RLHF dataset contains human preference data for training language models to be helpful and harmless, as well as red teaming data to identify harmful model outputs. The preference data includes pairs of chosen and rejected responses, while the red teaming data includes transcripts of adversarial interactions with AI assistants, rated for harmfulness. We strictly follow prior works (Song et al., 2024; Yu et al., 2023) and used the code from this repository  [^1] https://github.com/AlibabaResearch/DAMO-ConvAI/tree/main/PRO for testing.
  • MATH (Hendrycks et al., 2021): The MATH dataset consists of 12,500 challenging competition-level math problems, each with a detailed step-by-step solution. It is designed to teach models to generate answer derivations and explanations, aiding in mathematical reasoning. Despite progress in improving accuracy, the dataset highlights the limitations of large Transformer models in solving complex math problems without new algorithmic advancements. Due to the high resource cost of using the full test set, we randomly sampled 400 problems from the test set for evaluation.
  • GSM8k (Chen et al., 2021): GSM8K (Grade School Math 8K) is a collection of 8.5K high-quality math word problems designed for grade school students. It supports the task of multi-step reasoning and question answering in basic math. The problems require 2 to 8 steps, focusing on elementary arithmetic operations (addition, subtraction, multiplication, and division). The solutions are provided in natural language, making it accessible for evaluation of language models’ internal reasoning. GSM8K has been widely used to test logic and mathematical capabilities in language models, especially for benchmarks like the LLM Leaderboard. Due to the high computational cost of using the entire test set, we randomly sampled 400 data points from the test set for our evaluation.
  • GAIA (Mialon et al., 2023): The GAIA dataset is a benchmark designed to evaluate next-generation LLMs with augmented capabilities like tooling and search access. It consists of over 450 complex questions with unambiguous answers, requiring various levels of autonomy and tooling. The dataset is divided into three levels, each increasing in difficulty, with a public dev set for validation and a private test set for evaluation. We used the entire test set for our evaluation.
  • HumanEval (Chen et al., 2021): The OpenAI HumanEval dataset contains 164 programming problems, each with a function signature, docstring, body, and unit tests. These problems are handwritten to ensure they were not included in the training sets of code generation models. The dataset is designed to evaluate the performance of models in Python code generation. We used the entire test set for evaluation.
  • AlpacaEval. (Dubois et al., 2024b): AlpacaEval consists of 805 instructions, including 252 from the self-instruct test set (Wang et al., 2022), 188 from the Open Assistant (OASST) test set, 129 from Anthropic’s helpful test set (Zhou et al., 2023), 80 from the Vicuna test set (Chiang et al., 2023), and 156 from the Koala test set (Vu et al., 2023).
  • LIMA. (Zhou et al., 2023): LIMA collects a training dataset of 1000 prompts and responses, curated to ensure stylistic consistency while accommodating diverse input types. It also includes an open-source test set of 300 prompts and a development set of 50. The data is primarily sourced from community-driven Q&A platforms like Stack Exchange, wikiHow, and the Pushshift Reddit Dataset (Baumgartner et al., 2020), along with manually curated examples. The inclusion of human-authored examples further increases dataset diversity. In our experiments, we use the LIMA test set to evaluate our models.
  • Vicuna. (Chiang et al., 2023): Vicuna organizes its 80 test instructions into eight distinct categories: Fermi problems, commonsense, roleplaying, coding/math/writing tasks, counterfactuals, knowledge, and general questions. This categorization aims to comprehensively assess different facets of chatbot performance. Prior work suggests that Vicuna’s instructions are generally of lower complexity and difficulty (Xu et al., 2023). We utilize the Vicuna test set to assess the performance of large language models across these varied categories of instructions.
  • Self-Instruct. (Wang et al., 2022): Self-Instruct contains 252 human-authored test instructions, each paired with a well-constructed output. This dataset is curated to simulate real-world use cases of instruction-following models, spanning various domains such as email composition, social media, productivity tools, and coding tasks. The instructions differ in format and complexity, featuring diverse task lengths and output types such as bullet points, tables, code snippets, and mathematical expressions. In our research, we utilized the Self-Instruct test set to rigorously evaluate our model’s ability to follow detailed instructions across multiple domains.
  • Wizardlm. (Xu et al., 2023): Wizardlm consists of a training set of 70k examples derived from 52k instructions initially provided by Alpaca. The test set includes 218 instructions sourced from various open-source projects and online communities, covering 29 distinct skills derived from real-world tasks. These skills range from Code Generation & Debugging to Reasoning, Mathematics, Writing, Complex Format Handling, and Mastery of Extensive Domains. In our study, we employed the Wizardlm test set to evaluate the model’s ability to adhere to detailed instructions comprehensively.
  • Koala. (Vu et al., 2023): Koala comprises 180 real-world user queries sourced from the web, spanning diverse topics and typically reflecting a conversational tone. These queries are especially relevant for evaluating models intended for chat-based applications. To ensure no overlap with training data, any query yielding a BLEU score above 20% compared to examples from our training set is excluded. Additionally, queries involving programming or non-English languages are omitted, as our evaluation team, composed of crowd-sourced raters, lacks the expertise to assess such content effectively. We exclusively used the Koala test set to gauge our model’s proficiency in handling authentic conversational queries.
Appendix B Construction of the Preference Graph
In this section, we provide a detailed explanation of the process used to construct the preference graph sets for both the model ranking and response selection tasks, as outlined in Algorithms 1 and 2. These algorithms form the backbone of our method, enabling the representation of pairwise preferences as directed graphs, which are essential for downstream aggregation and ranking.
Algorithm 1 describes the construction process for generating a set of preference graphs for the model ranking task. The procedure is as follows: Initialization: For each question $q\in Q$, we begin by initializing a vertex set $V_{q}$, where each vertex $v_{i}$ corresponds to a model $\mathcal{M}_{i}$ in the set of models $M$. We also initialize an empty set of edges $A_{a}$ and a weight function $w_{a}$, which will be used to track the strength of the preferences between model pairs. Pairwise Comparisons: For each pair of models $\mathcal{M}_{i}$ and $\mathcal{M}_{j}$, the assigned evaluator $a\in\mathcal{A}$ assesses their responses to the given question $q$. If evaluator $a$ prefers $\mathcal{M}_{i}$ over $\mathcal{M}_{j}$, a directed edge $(v_{i}\rightarrow v_{j})$ is added to the edge set $A_{a}$, and its corresponding weight is incremented. Conversely, if $\mathcal{M}_{j}$ is preferred, the edge $(v_{j}\rightarrow v_{i})$ is added or its weight is updated. Graph Storage: Once all pairwise comparisons have been processed for a given evaluator, the resulting graph $G_{a}=(V_{q},A_{a},w_{a})$ is stored for that evaluator. This process is repeated for all evaluators in $\mathcal{A}$ and for all questions in $Q$, generating a set of preference graphs for each evaluator and question.
Algorithm 1 Construction of the Preference Graph for Model Ranking
1:Set of models $M=\{\mathcal{M}_{1},\mathcal{M}_{2},\ldots,\mathcal{M}_{n}\}$, set of questions $Q=\{q_{1},q_{2},\ldots,q_{t}\}$, set of evaluators $\mathcal{A}=\{a_{1},a_{2},\ldots,a_{k}\}$
2:Set of preference graph sets $\{G_{a}:a\in\mathcal{A}\}$ for each question $q\in Q$
3:for each question $q\in Q$ do
4:     for each evaluator $a\in\mathcal{A}$ do
5:         Initialize vertex set $V_{q}=\{v_{1},v_{2},\ldots,v_{n}\}$, where each $v_{i}$ corresponds to model $\mathcal{M}_{i}$
6:         Initialize edge set $A_{a}=\emptyset$, and weight function $w_{a}:A_{a}\rightarrow\mathbb{R}^{+}$
7:         for each pair of models $(\mathcal{M}_{i},\mathcal{M}_{j})$ with $i\neq j$ do
8:              Let $ans_{i}=\mathcal{M}_{i}(q)$ and $ans_{j}=\mathcal{M}_{j}(q)$
9:              if $a(ans_{i},ans_{j})>0$ then $\triangleright$ evaluator $a$ prefers $\mathcal{M}_{i}$ over $\mathcal{M}_{j}$
10:                  if $(v_{i},v_{j})\notin A_{a}$ then
11:                       Add directed edge $(v_{i}\rightarrow v_{j})$ to $A_{a}$
12:                       Set $w_{a}(v_{i},v_{j})=1$
13:                  else
14:                       Increment $w_{a}(v_{i},v_{j})$ by 1
15:                  end if
16:              else
17:                  if $(v_{j},v_{i})\notin A_{a}$ then
18:                       Add directed edge $(v_{j}\rightarrow v_{i})$ to $A_{a}$
19:                       Set $w_{a}(v_{j},v_{i})=1$
20:                  else
21:                       Increment $w_{a}(v_{j},v_{i})$ by 1
22:                  end if
23:              end if
24:         end for
25:         Store the preference graph $G_{a}=(V_{q},A_{a},w_{a})$ for evaluator $a$
26:     end for
27:end for
Algorithm 2 follows a similar structure but applies to the response selection task, where the objective is to rank a set of candidate answers for each question: Initialization: For each question $q\in Q$, we initialize a vertex set $V_{q}$, where each vertex corresponds to a candidate answer $ans_{i}$. As in the model ranking task, we also initialize an edge set $A_{a}$ and a weight function $w_{a}$ for each evaluator $a\in\mathcal{A}$. Pairwise Comparisons: Evaluators compare the quality of pairs of candidate answers $ans_{i}$ and $ans_{j}$ for each question. A directed edge is added based on the evaluator’s preference, with the weight reflecting the strength of preference. As before, if evaluator $a$ prefers $ans_{i}$ over $ans_{j}$, an edge $(v_{i}\rightarrow v_{j})$ is added or its weight incremented, and vice versa. Graph Storage: After all pairwise comparisons are complete, the preference graph $G_{a}=(V_{q},A_{a},w_{a})$ for evaluator $a$ is stored. This procedure is repeated for all evaluators and questions, resulting in a set of preference graphs for each evaluator and each question.
Both algorithms ensure that the preference graphs are constructed in a consistent manner, forming the basis for the aggregation and denoising processes used later in our framework. These graphs encapsulate the evaluators’ preferences and provide a structured representation of pairwise comparisons, facilitating further analysis.
Algorithm 2 Construction of the Preference Graph for Response Selection
1:Set of candidate answers $\{ans_{1},ans_{2},\ldots,ans_{n}\}$ for each question $q\in Q$, set of evaluators $\mathcal{A}=\{a_{1},a_{2},\ldots,a_{k}\}$
2:Set of preference graph sets $\{G_{a}:a\in\mathcal{A}\}$ for each question $q\in Q$
3:for each question $q\in Q$ do
4:     for each evaluator $a\in\mathcal{A}$ do
5:         Initialize vertex set $V_{q}=\{v_{1},v_{2},\ldots,v_{n}\}$, where each $v_{i}$ corresponds to $ans_{i}$
6:         Initialize edge set $A_{a}=\emptyset$, and weight function $w_{a}:A_{a}\rightarrow\mathbb{R}^{+}$
7:         for each pair of answers $(ans_{i},ans_{j})$ with $i\neq j$ do
8:              if $a(ans_{i},ans_{j})>0$ then
9:                  if $(v_{i},v_{j})\notin A_{a}$ then
10:                       Add directed edge $(v_{i}\rightarrow v_{j})$ to $A_{a}$, set $w_{a}(v_{i},v_{j})=1$
11:                  else
12:                       Increment $w_{a}(v_{i},v_{j})$ by 1
13:                  end if
14:              else
15:                  if $(v_{j},v_{i})\notin A_{a}$ then
16:                       Add directed edge $(v_{j}\rightarrow v_{i})$ to $A_{a}$, set $w_{a}(v_{j},v_{i})=1$
17:                  else
18:                       Increment $w_{a}(v_{j},v_{i})$ by 1
19:                  end if
20:              end if
21:         end for
22:         Store the preference graph $G_{a}=(V_{q},A_{a},w_{a})$ for evaluator $a$
23:     end for
24:end for
Appendix C Proof of Theorem 1
See 1
Proof C.1
For brevity, we consider all edges in $G$ have weights equal to 1 and all weights in $\widehat{G}$ are divided by $N$. By construction, we can notice that for each $(u\rightarrow v)\in A$, the weight $w_{\widehat{G}}(v\rightarrow u)$ can be viewed as an empirical estimate of $\delta_{1}$. Then, we claim that the following two events can imply $G\subseteq\mathrm{MAS}(\widehat{G})$:
  • $(\mathcal{E}_{1})$ For any $(u\rightarrow v)\in A$, it holds $\left|w_{\widehat{G}}(v\rightarrow u)-\delta_{1}\right|\leq\frac{\epsilon}{2}$.
  • $(\mathcal{E}_{2})$ For any pair of nodes $(u,v)$ such that $(u\rightarrow v),(v\rightarrow u)\notin A$, it holds $\left|w_{\widehat{G}}(u\rightarrow v)-w_{\widehat{G}}(v\rightarrow u)\right|\leq\frac{\epsilon}{U}$.
To see this, first by Lemma 1, we know that for any pair of nodes $(u,v)$, $\mathrm{MAS}(\widehat{G})$ will contain exactly one of $(u\rightarrow v)$ and $(v\rightarrow u)$. [^2] There is non-zero probability that some edges in $\widehat{G}$ will have zero weight, but we treat them as existing for the ease of argument. That is, we allow only $\widehat{G}$ to contain zero-weight edges. Therefore, for any $(u\rightarrow v)\in A$, $\mathrm{MAS}(\widehat{G})$ will contain exactly one of $(u\rightarrow v)$ and $(v\rightarrow u)$. Then, since $\mathcal{E}_{1}$ holds, $\delta_{1}=0.5-\epsilon<0.5$ and $w_{\widehat{G}}(u\rightarrow v)+w_{\widehat{G}}(v\rightarrow u)=1$, we have $w_{\widehat{G}}(u\rightarrow v)-w_{\widehat{G}}(v\rightarrow u)\geq\epsilon$. Furthermore, since $\mathcal{E}_{2}$ holds, for $(u,v)$ such that $(u\rightarrow v),(v\rightarrow u)\notin A$, arbitrary way of edge removing among these nodes can influence the total edge weights by at most $\epsilon$. Therefore, when applying the denoising operation to $\widehat{G}$, for any $(u\rightarrow v)\in A$, only $(u\rightarrow v)$ will be kept in $\mathrm{MAS}(\widehat{G})$, which makes $G\subseteq\mathrm{MAS}(\widehat{G})$. As a result, we have $\mathbb{P}(\mathcal{E}_{1}\cap\mathcal{E}_{2})\leq\mathbb{P}\left(G\subseteq\mathrm{MAS}(\widehat{G})\right)$.
Then, we can now bound the probability of $\mathcal{E}_{1}\cap\mathcal{E}_{2}$. In particular, for fixed $(u\rightarrow v)\in A$, since $w_{\widehat{G}}(v\rightarrow u)$ is an empirical mean estimate of $\delta_{1}$, by Hoeffding’s inequality, we have
$\mathbb{P}\left(\left|w_{\widehat{G}}(v\rightarrow u)-\delta_{1}\right|\leq\frac{\epsilon}{2}\right)\geq$ $1-2\exp\left(-N\epsilon^{2}/2\right)$
$\implies\mathbb{P}(\mathcal{E}_{1})\geq$ $1-2|A|\exp\left(-N\epsilon^{2}/2\right),$
where the second inequality comes from the union bound over all edges in $A$. Similarly, for fixed node pair $(u,v)$ that is unconnected in $G$, $w_{\widehat{G}}(u\rightarrow v)-w_{\widehat{G}}(v\rightarrow u)$ can be viewed as $\frac{1}{N}\sum_{i=1}^{N}X_{i}$, where $X_{i}$’s are i.i.d. and
$X_{i}=\begin{cases}1,&\text{with probability }\frac{\delta_{2}}{2}\\ -1,&\text{with probability }\frac{\delta_{2}}{2}\\ 0,&\text{with probability }1-\delta_{2}\end{cases}.$
Therefore, by Bernstein’s inequality, we have
$\mathbb{P}\left(\left|w_{\widehat{G}}(u\rightarrow v)-w_{\widehat{G}}(v\rightarrow u)\right|\leq\frac{\epsilon}{U}\right)\geq$ $1-2\exp\left(-\frac{N\epsilon^{2}}{6U^{2}\delta_{2}+2U\epsilon}\right)$
$\implies\mathbb{P}(\mathcal{E}_{2})\geq$ $1-2U\exp\left(-\frac{N\epsilon^{2}}{6U^{2}\delta_{2}+2U\epsilon}\right),$
where the second inequality is an union bound over all unconnected node pairs in $G$. As a result, we eventually have
$\mathbb{P}\left(G\subseteq\mathrm{MAS}(\widehat{G})\right)\geq$ $\mathbb{P}\left(\mathcal{E}_{1}\cap\mathcal{E}_{2}\right)\geq 1-2|A|\exp\left(-\frac{N\epsilon^{2}}{2}\right)-2U\exp\left(-\frac{N\epsilon^{2}}{6U^{2}\delta_{2}+2U\epsilon}\right).$
Lemma 1
For a weighted directed graph $G=(V,A,w)$, if $(u\rightarrow v),(v\rightarrow u)\in A$, then $\mathrm{MAS}(G)$ contains exactly one of $(u\rightarrow v)$ and $(v\rightarrow u)$.
Proof C.2
Recall that $\mathrm{MAS}(G)$ gives an acyclic subgraph of $G$ with the maximum weight. Since it has to be acyclic, it is obvious that $\mathrm{MAS}(G)$ cannot contain both $(u\rightarrow v)$ and $(v\rightarrow u)$.
We will then use contradiction to show it is impossible for $\mathrm{MAS}(G)$ to contain neither $(u\rightarrow v)$ nor $(v\rightarrow u)$. Suppose this is true instead. Then, since $\mathrm{MAS}(G)$ is a maximum acyclic subgraph, adding either $(u\rightarrow v)$ or $(v\rightarrow u)$ to $\mathrm{MAS}(G)$ will make it cyclic. That is, $\mathrm{MAS}(G)$ contains a path that goes from $v$ to $u$; meanwhile, it also contains a path from $u$ to $v$. As a result, it contains a cycle that goes from $v$ to $u$ and then from $u$ to $v$, which contradicts with the fact that $\mathrm{MAS}(G)$ is a maximum acyclic subgraph. Therefore, $\mathrm{MAS}(G)$ must contain exactly one of $(u\rightarrow v)$ and $(v\rightarrow u)$.
Appendix D Denoising of Preference Graphs for GED
In this section, we describe the denoising procedure used in the GED framework, specifically for transforming the aggregated preference graph into a DAG. Let $G=(V,A)$ be a simple connected directed graph, where $n=|V|$ and $m=|A|$, with $n-1\leq m\leq\binom{n}{2}$. A feedback arc set (FAS) of $G$, denoted as $R(G)$, is a set of arcs whose removal transforms $G$ into a DAG. The minimum feedback arc set $R^{*}(G)$ is the FAS of minimum cardinality, and finding $R^{*}(G)$ is the well-known FAS problem.
Consider a scenario where the vertices of $G$ are arranged sequentially along a horizontal line and labeled as $v_{1},v_{2},\ldots,v_{n}$ from left to right. This arrangement is referred to as a vertex sequence and is denoted by $s=v_{1}v_{2}\ldots v_{n}$. Each vertex sequence $s$ induces a feedback arc set $R(s)$, consisting of all leftward arcs $v_{j}\rightarrow v_{i}$ for $j>i$. The FAS problem can therefore be reformulated as finding a vertex sequence $s^{*}$ such that $R(s^{*})=R^{*}(G)$. Our proposed algorithm, GED, computes a vertex sequence $s$ that corresponds to a minimal feedback arc set $R(s)$. The algorithm iteratively removes vertices (and their incident arcs) from $G$, focusing on sinks, sources, and vertices that maximize a specific property. For any vertex $u\in V$, let $d(u)$ denote its degree, $d^{+}(u)$ its outdegree, and $d^{-}(u)$ its indegree, such that $d(u)=d^{+}(u)+d^{-}(u)$. At each step, after removing sinks and sources, the algorithm selects a vertex $u$ for which $\delta(u)=d^{+}(u)-d^{-}(u)$ is maximized. If the removed vertex $u$ is a sink, it is concatenated with a vertex sequence $s_{2}$; otherwise, it is concatenated with $s_{1}$. Once $G$ is reduced to an empty graph, the final vertex sequence $s$ is obtained by concatenating $s_{1}$ and $s_{2}$. The detailed steps of the algorithm are shown in Algorithm 3. In our GED framework, this denoising step is essential for ensuring that the aggregated preference graph becomes acyclic, thus enabling a reliable ranking to be extracted. By iteratively removing vertices based on their structural properties, we minimize the feedback arc set and ensure that the remaining graph is a DAG, which can be directly used to generate rankings in subsequent steps.
Algorithm 3 Preference Graphs Denoising for GED
1:$G$: Directed graph, var $s$: Vertex sequence
2:$s_{1}\leftarrow\emptyset$
3:$s_{2}\leftarrow\emptyset$
4:while $G\neq\emptyset$ do
5:     while $G$ contains a sink do
6:         Choose a sink $u$
7:         $s_{2}\leftarrow u\cdot s_{2}$
8:         $G\leftarrow G-u$
9:     end while
10:     while $G$ contains a source do
11:         Choose a source $u$
12:         $s_{1}\leftarrow s_{1}\cdot u$
13:         $G\leftarrow G-u$
14:     end while
15:     Choose a vertex $u$ with maximum $\delta(u)$
16:     $s_{1}\leftarrow s_{1}\cdot u$
17:     $G\leftarrow G-u$
18:end while
19:$s\leftarrow s_{1}\cdot s_{2}$
Appendix E Rank ensemble method
Weight score (Adler et al., 2002)
: The Weight Score method assigns a score to each vertex $v$ based on its position in each ranking $\mathcal{R}_{i}$. For a vertex $v$ in ranking $\mathcal{R}_{i}$, the score is given by:
$S_{i}(v)=l_{i}-r_{i}(v)+1$ (6)
where $l_{i}$ is the length of ranking $\mathcal{R}_{i}$ and $r_{i}(v)$ is the rank of vertex $v$ in $\mathcal{R}_{i}$. If $v$ is not present in $\mathcal{R}_{i}$, $S_{i}(v)=0$. The total score for each vertex across all rankings is:
$T(v)=\sum_{i=1}^{k}S_{i}(v)$ (7)
The final consensus ranking $\mathcal{R}^{*}$ is obtained by sorting the vertices $v$ in descending order of $T(v)$.
Kemeny and weighted Kemeny (Kemeny, 1959)
: The Kemeny method seeks a consensus ranking $\mathcal{R}^{*}$ that minimizes the total pairwise disagreements between $\mathcal{R}^{*}$ and the input rankings, measured using the Kendall $\tau$-distance:
$\mathcal{R}^{*}=\arg\min_{\mathcal{R}}\sum_{i=1}^{k}\tau(\mathcal{R},\mathcal{R}_{i})$ (8)
The Weighted Kemeny method introduces a weight $\alpha_{i}$ for each ranking $\mathcal{R}_{i}$, reflecting its importance or reliability:
$\mathcal{R}^{*}=\arg\min_{\mathcal{R}}\sum_{i=1}^{k}\alpha_{i}\cdot\tau(\mathcal{R},\mathcal{R}_{i})$ (9)
Here, the goal is to minimize the weighted Kendall tau distance, emphasizing rankings with higher weights.
Pairwise majority and weighted pairwise majority (Caragiannis et al., 2016)
: The Pairwise Majority (PM) method determines a consensus ranking $\mathcal{R}^{*}$ by maximizing the number of pairwise agreements with the input rankings. For each pair of vertices $(v_{i},v_{j})$, the goal is to ensure that the majority of rankings agree with their relative order in $\mathcal{R}^{*}$:
$\mathcal{R}^{*}=\arg\max_{\mathcal{R}}\sum_{i<j}\left(\sum_{p=1}^{k}\mathds{1}(\mathcal{R}_{p}(v_{i})<\mathcal{R}_{p}(v_{j}))\right)\cdot\mathds{1}(\mathcal{R}(v_{i})<\mathcal{R}(v_{j}))$ (10)
The Weighted Pairwise Majority method incorporates weights $\alpha_{p}$ to account for the reliability of each ranking $\mathcal{R}_{p}$:
$\mathcal{R}^{*}=\arg\max_{\mathcal{R}}\sum_{i<j}\left(\sum_{p=1}^{k}\alpha_{p}\cdot\mathds{1}(\mathcal{R}_{p}(v_{i})<\mathcal{R}_{p}(v_{j}))\right)\cdot\mathds{1}(\mathcal{R}(v_{i})<\mathcal{R}(v_{j}))$ (11)
In both methods, the objective is to maximize the (weighted) pairwise agreement between the consensus ranking and the input rankings.
Appendix F Visualization
(a) Raw Preference Graph 1
(b) Denoised Preference Graph 1
(c) Raw Preference Graph 2
(d) Denoised Preference Graph 2
Figure 6: Case studies showcasing the raw and denoised preference graphs. In both Case 1 and Case 2, the raw preference graphs (a, c) contain cyclic inconsistencies, which are resolved by GED into directed acyclic graphs (b, d). The dashed lines in the denoised graphs represent the edges that were removed by GED to eliminate cycles. The nodes labeled 0-9 correspond to the ten generated responses by Qwen2-72B. These examples illustrate the effectiveness of GED in eliminating noise and restoring consistency in preference evaluations.
In this section, we present two case studies that demonstrate the effectiveness of our proposed GED method in denoising preference graphs. Figure 6 illustrates the raw preference graphs, which are generated by multiple evaluators, i.e., Llama3-8B, Mistral-7B, and Qwen2-7B, through the responses produced by Qwen2-72B on the HumanEval benchmark. The nodes labeled 0-9 in the graphs correspond to the ten generated responses. The comparison between the raw (left) and denoised (right) graphs shows how our method successfully resolves cyclic inconsistencies, transforming noisy graphs into DAGs. In the denoised graphs, the dashed lines represent the edges that were removed by GED to eliminate cycles.
Case Study 1.
In the first case study, we showcase the impact of our denoising approach. The raw graph (Figure 6 (a)) contains multiple cyclic inconsistencies, such as $9\succ 4\succ 3\succ 9$, $9\succ 4\succ 3\succ 2\succ 9$, and $9\succ 4\succ 3\succ 7\succ 9$. By applying GED, we identify that removing a single edge (4 $\succ$ 3) eliminates all cycles, converting the noisy preference graph into a consistent DAG, as shown in Figure 6 (b).
Case Study 2.
The second case study (Figure 6 (c)) presents another scenario with conflicting preferences, where cycles like $7\succ 0\succ 6\succ 7$ and $7\succ 1\succ 0\succ 6\succ 7$ indicate noisy judgments. Here, removing just one edge (0 $\succ$ 6) using GED is sufficient to eliminate all cycles and convert the graph into a DAG, as depicted in Figure 6 (d).
Conclusion from Case Studies.
In both case studies, removing any other edge would not fully resolve all cyclic inconsistencies without requiring additional deletions, which would result in more information loss. GED effectively minimizes edge removals while maintaining the integrity of the original preference graph, making it a highly efficient solution for improving the consistency of preference evaluations.

### v2

@@
 Preference evaluation, as one of the most important assessment methods, plays an indispensable role in evaluating and optimizing model performance (Rafailov et al., 2024; Yuan et al., 2024; Dubois et al., 2024b).
-Existing works usually leverage a powerful LLM (e.g., GPT4 (Achiam et al., 2023)) as the judge for comparing LLMs’ output pairwisely (Li et al., 2023b; Chen et al., 2023; Li et al., 2023b; Wang et al., 2022).
+Existing works usually leverage a powerful LLM (e.g., GPT4 (Achiam et al., 2023)) as the judge for comparing LLMs’ output pairwisely (Li et al., 2023b; Chen et al., 2023b; Wang et al., 2022).
 Figure 1: The figure is divided into two parts.
@@
 The lower part presents the evaluation of ten responses for HumanEval (Chen et al., 2021) and MATH (Hendrycks et al., 2021) tasks, generated by Llama3-70B (AI@Meta, 2024).
-These were evaluated by GPT-4-o, GPT-4-o-mini, GPT-3.5 (Achiam et al., 2023), Qwen2-72B (Yang et al., 2024a), and Llama3-8B, showing that even advanced LLMs exhibit significant noise in their evaluations, illustrating their limitations as reliable evaluators.
-However, while such model-based pairwise preference evaluations offer a flexible approach, they can lead to contradictory evaluations in the assessment process (Zhang et al., 2024b).
+These were evaluated by GPT-4-o, GPT-4-o-mini, GPT-3.5 (Achiam et al., 2023), Qwen2-72B (Yang et al., 2024a), and Llama3-8B, revealing a significant cycle rate (details in Appendix A.5) in their evaluations.
+This highlights the limitations of even advanced LLMs as reliable evaluators.
+However, while such model-based pairwise preference evaluations offer a flexible approach, they can lead to contradictory evaluations in the assessment process (Naresh et al., 2024; Zhang et al., 2024b).
 For example, an LLM might evaluate three responses and conclude that Response A is better than Response B (A $\succ$ B), Response B is better than Response C (B $\succ$ C), yet paradoxically also rank Response C as better than Response A (C $\succ$ A).
@@
 Ideally, a preference graph should be structured as a directed acyclic graph (DAG) to maintain consistency.
-As shown in the bottom half of Figure  1, even advanced LLMs like GPT-4-o exhibit significant noise in preference evaluation, highlighting their limitations as weak evaluators.
+As shown in the bottom half of Figure  1, even advanced LLMs like GPT-4-o exhibit significant noise in preference evaluation, highlighting their limitations as weak evaluators (Refer to Appendix A.3 for further details).
 To address this, we propose a novel framework, GED (Preference Graph Ensemble and Denoise), to address the inconsistencies in preference graphs generated through pairwise evaluations.
@@
 (1993).
-Details can be found in Appendix D.
+Details can be found in Appendix L.
 Graph to ranking.
@@
 Evaluators assess the quality of these answers, assigning a preference that is denoted by a directed arc $(v_{i},v_{j})$ in $A_{a}$, with a corresponding weight $w_{a}(v_{i},v_{j})$ based on the strength of preference.
-This process is detailed in Appendix B.
+This process is detailed in Appendix D.
 After collecting the preference graphs $\{G_{a}:a\in\mathcal{A}\}$ for a question $q$, we apply GED to aggregate these graphs.
@@
 This preference is represented by a directed arc $(v_{i},v_{j})$ in $A_{a}$, assigned a weight $w_{a}(v_{i},v_{j})$ based on preference strength.
-The detailed procedure is outlined in Appendix B.
+The detailed procedure is outlined in Appendix D.
 Once the preference graphs $\{G_{a}:a\in\mathcal{A}\}$ are collected for a question $q$, we employ GED to aggregate these graphs.
@@
 This process is repeated for each question $q\in Q$, yielding a set of rankings $\{\mathcal{R}_{q}:q\in Q\}$.
-To compute the overall ranking $\mathcal{R}^{*}$ of the models across all questions, we conduct a ranking ensemble on the set $\{\mathcal{R}_{q}:q\in Q\}$, as detailed in Appendix E.
+To compute the overall ranking $\mathcal{R}^{*}$ of the models across all questions, we conduct a ranking ensemble on the set $\{\mathcal{R}_{q}:q\in Q\}$, as detailed in Appendix M.
 This approach culminates in a final ranking that reflects the models’ performance as assessed by multiple evaluators in $\mathcal{A}$.
@@
 where $G\subseteq\text{MAS}(\widehat{G})$ represents that $G$ is a subgraph of $\text{MAS}(\widehat{G})$ and $U=\frac{|V|(|V|-1)}{2}-|A|$ is the number of pairs of unconnected nodes in $G$.
-The full proof is given in Appendix C.
+The full proof is given in Appendix E.
 From the theorem, we can see that the probability of failure decreases exponentially as the number of samples $N$ increases.
@@
 Method HumanEval AlpacaEval MATH GSM8k GAIA Avg
-Llama3-8B 43.90 27.29 22.08 56.67 6.78 31.14
-Mistral-7B 23.17 11.80 23.25 39.83 7.03 15.21
-Qwen2-7B 48.58 25.71 59.92 76.75 7.70 43.17
-Qwen2-72B 57.93 29.58 72.75 84.67 11.52 51.91
+Llama3-8B 43.90 27.29 22.08 56.67 6.78 31.34
+Mistral-7B 23.17 11.80 23.25 39.83 7.03 21.01
+Qwen2-7B 48.58 25.71 59.92 76.75 7.70 43.73
+Qwen2-72B 57.93 29.58 72.75 84.67 11.52 51.29
 ContraSolver 65.42 31.12 74.95 86.84 12.22 54.11
 ListPreference 61.52 31.67 71.75 85.0 10.90 52.16
-Single model Self-consistency 60.98 29.33 73.58 84.91 8.86 51.13
-Llama3-8B 62.19 29.31 74.27 83.16 11.31 52.87
-with graph denoising 64.02 30.18 74.73 86.00 11.72 53.53
-Mistral-7B 67.24 27.70 74.41 83.83 10.50 52.93
-with graph denoising 68.73 29.93 74.77 83.91 10.74 53.21
-Qwen2-7B 61.58 28.69 74.50 85.41 11.11 52.84
+Single model Self-consistency 60.98 29.33 73.58 84.91 8.86 51.53
+Llama3-8B 62.19 29.31 74.27 83.16 11.31 52.04
+with graph denoising 64.02 30.18 74.73 86.00 11.72 53.33
+Mistral-7B 67.24 27.70 74.41 83.83 10.50 52.73
+with graph denoising 68.73 29.93 74.77 83.91 10.74 53.61
+Qwen2-7B 61.58 28.69 74.50 85.41 11.11 52.25
 with graph denoising 65.85 29.44 74.79 86.38 11.25 53.54
-Qwen2-72B 60.97 31.04 74.73 86.47 12.14 53.87
-Single evaluator with graph denoising 68.90 31.17 75.33 87.45 12.26 55.42
-GED (w/o denoising) 69.25 30.98 74.29 87.17 12.68 54.83
-Multiple evaluator GED 70.73 32.44 75.58 88.18 13.33 56.41
+Qwen2-72B 60.97 31.04 74.73 86.47 12.14 53.07
+Single evaluator with graph denoising 68.90 31.17 75.33 87.45 12.26 55.02
+Multi-MV 66.18 29.57 74.77 86.42 11.72 53.73
+GED (w/o denoising) 69.25 30.98 74.29 87.17 12.68 54.87
+Multiple evaluator GED 70.73 32.44 75.58 88.18 13.33 56.05
 Figure 3: Comparison of GED with GPT-3.5, GPT-4-o-mini, and GPT-4-o on 100 randomly selected tasks.
@@
 Finally, in the multiple weak evaluators setup, we combine three weaker evaluators (Llama3-8B, Qwen2-7B, Mistral-7B) to select responses from Qwen2-72B with GED.
+We also include a new baseline referred to as Multi-MV, which just takes the majority for each pairwise comparison as the preferred answer.
 We present the results of GED and its variant (w/o denoising), which ensembles the preference graphs without the denoising step.
@@
 The substantial performance gains observed when using GED with denoising, compared to both the single evaluator setup and the ensemble without denoising, highlight its importance in refining response selection.
+We also evaluated Multi-MV, which aggregates pairwise comparisons using majority voting across evaluators.
+While Multi-MV offers improvements over individual evaluators, it underperforms compared to GED, highlighting GED’s ability to capture more nuanced evaluation signals and reduce inconsistencies.
 Additionally, we observed that the ListPreference baseline performed worse than Qwen2-72B as single evaluator, likely due to LLM limitations in handling long-text.
@@
 These results highlight the superiority of GED, particularly in leveraging multi-weak evaluators and graph denoising to outperform individual state-of-the-art models.
-Alabation study.
+Ablation study.
 Figure 4: Comparison of GED and (w/o ensemble) variants.
@@
 We evaluate the impact of removing the ensembling step in GED, referred to as the (w/o ensemble) variant.
-In this case, individual evaluators’ preference graphs are denoised and converted to rankings, which are then aggregated using methods such as Weight Score, Kemeny, Weighted Kemeny, Pairwise Majority, and Weighted Pairwise Majority (detailed in Appendix E).
+In this case, individual evaluators’ preference graphs are denoised and converted to rankings, which are then aggregated using methods such as Weight Score, Kemeny, Weighted Kemeny, Pairwise Majority, and Weighted Pairwise Majority (detailed in Appendix M).
 For simplicity of presentation, we use Weight Score to represent GED (w/o ensemble) (Weight Score).
@@
 These results are summarized in Table 3.
-To ensure a comprehensive evaluation, we also tested the models, using the Llama-2-7B backbone, across additional benchmarks such as LIMA (Zhou et al., 2023), Vicuna (Chiang et al., 2023), Koala (Vu et al., 2023), WizardLM (Xu et al., 2023) and Self-Instruct (Wang et al., 2022), in line with recent works (Chen et al., 2023; Zhang et al., 2024a; Hu et al., 2024).
+To ensure a comprehensive evaluation, we also tested the models, using the Llama-2-7B backbone, across additional benchmarks such as LIMA (Zhou et al., 2023), Vicuna (Chiang et al., 2023), Koala (Vu et al., 2023), WizardLM (Xu et al., 2023) and Self-Instruct (Wang et al., 2022), in line with recent works (Chen et al., 2023b; Zhang et al., 2024a; Hu et al., 2024).
 Table 3: Performance comparison of different methods (Random, Longest, ContraSolver, and GED) on model alignment task across the HH-RLHF benchmark.
@@
 Although they achieve high accuracy on matching inner-evaluator, the achievement suffers from spurious correlations such as perplexity and length (Durmus et al., 2022).
-Recently, people have started using a strong model (e.g., GPT-4) as an evaluator to perform a zero-shot reference-free evaluation on the weak models (Shen et al., 2023; Dubois et al., 2024b; Chen et al., 2023).
+Recently, people have started using a strong model (e.g., GPT-4) as an evaluator to perform a zero-shot reference-free evaluation on the weak models (Shen et al., 2023; Dubois et al., 2024b; Chen et al., 2023b).
 However, using LLM-based preference evaluations can introduce inconsistencies in preference graphs, often resulting in cyclic preferences or contradictions when comparing multiple outputs.
@@
 The concept of weak-to-strong supervision originates from the need to leverage noisy or partial labels in machine learning tasks, enabling the development of more robust models from imperfect data (Ratner et al., 2016; Zhang et al., 2023b; 2022).
-In LLMs, weak-to-strong supervision aids AI alignment by allowing weaker models to improve strong ones, enhancing performance without extensive data and supporting scalable oversight (Zheng et al., 2024; Guo & Yang, 2024; Tong et al., 2024).
+In LLMs, weak-to-strong supervision aids AI alignment by allowing weaker models to improve strong ones, enhancing performance without extensive data and supporting scalable oversight (Zheng et al., 2024a; Guo & Yang, 2024; Tong et al., 2024).
 Similarly, in task-oriented LLMs, weak-to-strong learning improves LLM’s ability by enabling strong models to refine their data autonomously, boosting performance without extensive high-quality input (Zhang et al., 2023a; Yang et al., 2024b).
@@
 Future work will explore extending GED to broader evaluation frameworks and applying its principles to more complex decision-making tasks, including multi-agent systems and human-AI interaction.
+Acknowledgement
+This project was partially supported by an OpenAI “Superalignment Fast” grant.
 Appendix A Implementation Details
@@
 All experimental procedures were conducted on a machine equipped with an AMD EPYC 7543 32-Core Processor, 512GB memory, 128 CPUs, and four 80GB NVIDIA A800 GPUs.
-The code is available at https://github.com/ppsmk388/GED.
-The references to Llama-2-7B, Llama3-70B, Llama3-8B, Mistral-7B, Qwen2-7B, and Qwen2-72B in the main text refer to the specific models: Llama-2-7b-chat-hf, Meta-Llama-3-70B-Instruct, Meta-Llama-3-8B-Instruct, Mixtral-8x7B-Instruct-v0.1, Qwen2-7B-Instruct, and Qwen2-72B-Instruct.
+The references to Llama-2-7B, Llama3-70B, Llama3-8B, Mistral-7B, Mistral-8×7B, Qwen2-7B, and Qwen2-72B in the main text refer to the specific models: Llama-2-7b-chat-hf, Meta-Llama-3-70B-Instruct, Meta-Llama-3-8B-Instruct, Mixtral-7B-Instruct-v0.3, Mixtral-8x7B-Instruct-v0.1, Qwen2-7B-Instruct, and Qwen2-72B-Instruct.
 We utilized the reward model oasst-rm-2-pythia-6.9b-epoch-1 following prior works (Song et al., 2024; Yu et al., 2023).
 Each experiment was repeated three times, and the average performance was reported as the final result.
+Our training script was adapted from the example provided in LlamaFactory (Zheng et al., 2024b), specifically https://github.com/hiyouga/LLaMA-Factory/blob/main/examples/train_full/llama3_full_sft_ds3.yaml.
 The training was configured with a batch size of 1 per device, gradient accumulation steps of 4, a learning rate of 1e-5, and he model was trained for 3 epochs, with warmup over 20 steps and a cosine learning rate scheduler.
+For generating diverse responses from LLMs, we followed the configuration in Yuan et al.
+(2024), setting $T=0.7$ and $p=0.9$.
 For tasks such as AlpacaEval (Dubois et al., 2024b), we used GPT-4-o unless stated otherwise.
-A.2 Dataset
+A.2 Details of Evaluator Selection Across Different Tasks
+In this subsection, we provide more detailed information about the selection of evaluators across different tasks.
+Response Selection.
+In the response selection task, we evaluated both single models and ensembles of evaluators using the GED method.
+For single model evaluation, we assessed the standalone performance of Llama3-8B, Mistral-7B, Qwen2-7B, and Qwen2-72B on benchmarks such as HumanEval, AlpacaEval, MATH, GSM8k, and GAIA.
+This served two purposes: first, to establish the baseline performance of smaller models (Llama3-8B, Mistral-7B, Qwen2-7B) that are later combined using GED, and second, to compare against Qwen2-72B, where we tested a baseline approach by randomly selecting one response from the 10 it generated for each question.
+For single evaluator evaluation, each model (Llama3-8B, Mistral-7B, Qwen2-7B, and Qwen2-72B) was also used as an evaluator to rank responses generated by Qwen2-72B.
+Notably, Qwen2-72B acted as a self-evaluator, selecting the best response from its own generated outputs.
+Finally, for multiple evaluator evaluation, our GED method combined the evaluations of Llama3-8B, Mistral-7B, and Qwen2-7B to rank responses.
+Model Ranking.
+For the model ranking task, we selected larger models as evaluators to ensure alignment with rankings produced by GPT-4.
+Specifically, we used Llama3-70B, Qwen2-72B, Qwen1.5-72B, and Mistral-8×7B as single evaluators.
+This choice was guided by two factors: performance considerations and practical feasibility.
+Larger models generally produce more reliable rankings, closely aligning with GPT-4, and the AlpacaEval benchmark, containing 805 tasks, makes the computational cost of using larger models acceptable.
+In the multiple evaluator setting, our GED method aggregated the evaluations from these four larger models to produce robust and consistent rankings.
+The combination of these high-capacity models ensures that our approach yields rankings that are both accurate and consistent across tasks.
+Instruction Tuning.
+In the instruction tuning task, the objective was to perform data selection for model training.
+For this task, we employed Llama3-8B, Mistral-7B, and Qwen2-7B as evaluators.
+These models were selected because they balance computational efficiency and evaluation performance, making them suitable for iterative instruction tuning processes.
+The evaluators’ pairwise preferences were aggregated using GED to identify the most appropriate responses for training.
+This approach ensured that the selected data pairs reflected a consensus among the evaluators while keeping computational costs manageable.
+A.3 Definition of Weak Evaluators
+In this work, ”weak evaluators” are defined by their tendency to produce noisy or inconsistent pairwise preferences, not by their overall model capacity.
+Even advanced models like GPT-4-o can generate preference graphs with cycles, indicating that preference inconsistency is a universal challenge, regardless of model scale.
+Evaluators such as Llama3-8B, Mistral-7B, Qwen2-7B, and even GPT-4-o are considered weak if their pairwise evaluations exhibit significant noise or conflicts.
+A.4 Evaluation Settings: Single Model vs. Single Evaluator
+The ”single model” setting evaluates each model’s (Llama3-8B, Mistral-7B, Qwen2-7B, Qwen2-72B) outputs directly on benchmarks like HumanEval, AlpacaEval, MATH, GSM8k, and GAIA, without selection or modification.
+In contrast, the ”single evaluator” setting uses these models to select the best response from ten candidates generated by Qwen2-72B, assessing their evaluation capability.
+The key difference is that the single model setting focuses on generation quality, while the single evaluator setting assesses evaluation ability.
+A.5 Definition of Cycle Rate
+The Cycle Rate is the percentage of preference graphs with at least one cycle, indicating inconsistency in pairwise comparisons.
+For example, if an evaluator produces cycles in 100 out of 164 graphs for the HumanEval dataset, the Cycle Rate is $\frac{100}{164}\times 100=60.97\%$.
+A lower Cycle Rate indicates greater consistency, while a higher rate suggests evaluator biases or difficulties with ambiguous comparisons.
+This metric helps assess the reliability of evaluators, such as GPT-4-o and GPT-4-o-mini, across different datasets.
+A.6 Dataset
 In this appendix, we provide detailed information about the datasets used in main text.
@@
 We exclusively used the Koala test set to gauge our model’s proficiency in handling authentic conversational queries.
-Appendix B Construction of the Preference Graph
+A.7 Aggregation Process in GED Across Different Tasks
+The GED implementation differs between response selection and model ranking tasks due to their objectives.
+In response selection, GED aggregates preference graphs from multiple evaluators for each question into a single DAG.
+From this DAG, a final ranking is derived, and the top-ranked answer is selected as the output.
+Notably, aggregation is performed only at the per-question level, without a rank aggregation across questions.
+In model ranking, GED involves two steps: first, aggregating the evaluators’ preference graphs into a DAG for each question to rank the models, and second, employing a rank ensemble method to aggregate these per-question rankings into a final overall model ranking across all questions.
+For details on the aggregation modeling, see Section 2.3, covering Response Selection, Model Ranking, and Model Alignment.
+Appendix B Alternative Evaluator Configurations for Model Ranking
+To address concerns about the computational cost and fairness of using 70B-level models as weak evaluators in the model ranking task, we conducted additional experiments using smaller, more comparable models.
+Table 4 summarizes the performance of these models, both individually and when combined using GED.
+The results show that GED outperforms individual evaluators even when using smaller 7B-scale models, achieving an average score of 62.70 compared to the best individual performance of 57.92 (Qwen2-7B with graph denoising).
+This demonstrates that the aggregation of smaller models through GED effectively enhances performance while reducing computational costs.
+These findings validate the versatility of GED, showing that it can provide robust and accurate rankings without relying solely on large-scale models.
+Table 4: Performance comparison in the model ranking task using 7B-scale models as evaluators on the AlpacaEval dataset.
+Higher values indicate better performance.
+GED achieves robust performance even with smaller evaluators.
+Model Weight Score Kemeny
+Weighted
+Kemeny
+Pairwise
+Majority
+Weighted
+Pairwise
+Majority
+Avg.
+Llama3-8B 35.88 45.80 45.80 47.23 46.85 44.31
+with graph denoising 37.44 47.54 47.54 48.92 48.18 45.92
+Qwen2-7B 55.34 52.87 52.87 56.05 56.59 54.74
+with graph denoising 56.05 57.43 57.43 59.32 59.41 57.92
+Mistral-7B 49.90 53.74 53.74 58.06 57.87 54.66
+Single evaluator with graph denoising 50.47 55.06 54.92 61.39 61.21 56.61
+Multiple evaluator GED 57.59 61.14 61.14 67.17 66.46 62.70
+Appendix C Impact of Evaluator Quantity on Denoising Quality
+We investigated how the number of evaluators affects the denoising quality in GED by conducting experiments using different numbers of evaluators in the response selection setting.
+Specifically, we evaluated the performance of GED when aggregating preferences from two, three, and four evaluators.
+The evaluators used were Llama3-8B, Mistral-7B, Qwen2-7B, and Gemma-9B [^2] https://huggingface.co/google/gemma-2-9b-it.
+Table 5: Performance comparison of GED with varying numbers of evaluators across five benchmarks.
+Increasing the number of evaluators enhances denoising quality, reflected in improved performance metrics.
+Evaluators Set HumanEval AlpacaEval MATH GSM8k GAIA Avg.
+Llama3-8B, Mistral-7B 69.21 31.87 74.97 86.92 12.51 55.10
+Llama3-8B, Mistral-7B, Qwen2-7B 70.73 32.44 75.58 88.18 13.33 56.05
+Llama3-8B, Mistral-7B, Qwen2-7B, Gemma-9B 70.98 32.87 75.91 88.75 13.46 56.39
+As shown in Table 5, increasing the number of evaluators consistently improves the denoising quality of GED, as reflected by higher performance across all benchmarks.
+For instance, the average performance improves from 55.10% with two evaluators to 56.39% with four evaluators.
+This enhancement can be attributed to the diverse perspectives and complementary strengths of multiple evaluators, contributing to a more robust and accurate aggregation of preferences.
+These results highlight the importance of selecting a diverse and capable set of evaluators to enhance GED’s effectiveness.
+Incorporating more evaluators allows the denoising process to better identify and mitigate inconsistencies in the preference graphs, leading to improved overall performance in response selection tasks.
+Appendix D Construction of the Preference Graph
 In this section, we provide a detailed explanation of the process used to construct the preference graph sets for both the model ranking and response selection tasks, as outlined in Algorithms 1 and 2.
@@
 24:end for
-Appendix C Proof of Theorem 1
+Appendix E Proof of Theorem 1
 See 1
-Proof C.1
+Proof E.1
 For brevity, we consider all edges in $G$ have weights equal to 1 and all weights in $\widehat{G}$ are divided by $N$.
@@
 For a weighted directed graph $G=(V,A,w)$, if $(u\rightarrow v),(v\rightarrow u)\in A$, then $\mathrm{MAS}(G)$ contains exactly one of $(u\rightarrow v)$ and $(v\rightarrow u)$.
-Proof C.2
+Proof E.2
 Recall that $\mathrm{MAS}(G)$ gives an acyclic subgraph of $G$ with the maximum weight.
@@
 Therefore, $\mathrm{MAS}(G)$ must contain exactly one of $(u\rightarrow v)$ and $(v\rightarrow u)$.
-Appendix D Denoising of Preference Graphs for GED
+Appendix F Impact of Evaluator Weighting on GED Performance
+In our theoretical analysis, GED assumes equal weighting of edges in the preference graphs.
+However, in practical scenarios, evaluators may have varying levels of reliability or expertise.
+Incorporating evaluator-specific confidence scores or performance metrics could enhance the effectiveness of GED.
+To investigate this, we conducted experiments using a weighted version of GED, referred to as WeightGED, under the Response Selection setting.
+In WeightGED, we assigned weights to evaluators based on their individual performance on specific datasets.
+For example, on the GSM8k dataset, the response selection accuracies for Llama3-8B, Mistral-7B, and Qwen2-7B were 62.19%, 67.24%, and 61.58%, respectively.
+These accuracies were normalized to compute evaluator weights:
+[…]
+[…]
+[…]
+These weights were used to scale the contributions of each evaluator’s preferences during graph construction.
+We compared the performance of GED and WeightGED across multiple benchmarks, as presented in Table 6.
+Table 6: Performance comparison of GED and WeightGED using 7B-scale evaluators across five benchmarks.
+WeightGED marginally outperforms GED, demonstrating the potential benefits of incorporating evaluator-specific weights.
+Method HumanEval AlpacaEval MATH GSM8k GAIA Avg.
+GED 70.73 32.44 75.58 88.18 13.33 56.05
+7B Evaluators WeightGED 70.97 32.67 75.71 88.24 13.56 56.23
+GED 73.21 59.87 82.49 86.43 16.27 63.65
+GPT Evaluators WeightGED 74.52 61.71 83.93 87.84 17.32 65.06
+As shown in Table 6, WeightGED achieves marginal but consistent improvements over GED across all benchmarks.
+For instance, using 7B-scale evaluators, WeightGED improves the average performance from 56.05% to 56.23%.
+Similarly, with GPT evaluators, the average performance increases from 63.65% to 65.06%.
+These results suggest that incorporating evaluator-specific weights based on performance metrics can enhance the effectiveness of GED.
+Furthermore, we conducted additional experiments using stronger evaluators such as GPT-3.5, GPT-4-o-mini, and GPT-4-o.
+The weights were computed based on their respective performance accuracies, following the same normalization procedure.
+The improvements observed with these evaluators reinforce the potential of weighting schemes in enhancing GED.
+In summary, our preliminary findings indicate that integrating evaluator-specific confidence scores or performance metrics is a promising direction for future work.
+Systematically designing and optimizing these weighting schemes could lead to more robust and accurate evaluation frameworks.
+Appendix G Evaluating GED on Diverse Metrics
+To address the need for more comprehensive assessments beyond accuracy-based metrics, we expanded our evaluation of GED to include nuanced quality aspects of LLM-generated outputs, such as factuality, relevance, coherence, informativeness, helpfulness, and validity.
+This evaluation was motivated by the understanding that tasks involving LLMs often require subtle judgments beyond simple accuracy, and GED’s adaptability to these scenarios is crucial.
+Building upon this foundation, we conducted experiments following (Chen et al., 2023a).
+Specifically, we used Llama3-70B to generate ten candidate responses for each query, and GED was applied in the response selection setting with Llama3-8B, Mistral-7B, and Qwen2-7B serving as evaluators.
+Each metric—factuality, relevance, coherence, informativeness, helpfulness, and validity—was assessed to measure GED’s ability to enhance the overall quality of LLM outputs.
+The results, shown in Table 7, highlight GED’s effectiveness.
+GED consistently outperformed individual evaluators and random selection across all metrics.
+For instance, GED improved factuality by approximately 5 percentage points over the best individual evaluator.
+Similarly, it enhanced relevance and coherence, indicating better alignment with the query and logically consistent responses.
+GED also demonstrated notable improvements in informativeness and helpfulness, suggesting that the responses selected by GED provide more valuable and user-centric information.
+These findings confirm GED’s adaptability to tasks requiring nuanced quality assessments.
+By aggregating preferences from multiple evaluators, GED captures subtle qualities of generated content, enhancing not only accuracy but also the overall reliability and utility of LLM outputs.
+This adaptability reinforces GED’s value in scenarios demanding comprehensive evaluations of language models.
+Table 7: Performance comparison on nuanced quality metrics (in %).
+GED outperforms individual evaluators and random selection across Factuality, Relevance, Coherence, Inform., Helpful.
+and Validity metrics.
+Method Factuality Relevance Coherence Inform. Helpful. Validity Avg.
+Random 86.73 87.91 92.47 77.62 17.48 48.92 68.52
+Llama3-8B 88.59 89.91 94.41 79.77 18.48 50.92 70.35
+Mistral-7B 89.10 90.29 94.85 79.95 18.55 51.13 70.65
+Qwen2-7B 89.25 90.44 95.03 80.09 18.58 51.21 70.77
+GED 94.73 95.91 97.36 86.62 19.48 55.92 75.00
+Appendix H Expanded Related Work on Preference Denoising for LLMs
+This work situates itself within the broader field of denoising preference evaluations for LLMs, addressing inconsistencies and noise in preference graphs.
+We acknowledge that existing literature has explored two primary approaches to this challenge: within-model preference modeling and modular pre-processing.
+Below, we provide a detailed comparison of GED with representative methods from these approaches.
+Within-model approaches, such as robust DPO (rDPO) (Chowdhury et al., 2024) and conservative DPO (cDPO) (Mitchell, 2023), focus on integrating denoising mechanisms directly within the preference modeling process.
+These methods incorporate regularization techniques to mitigate the effects of noisy or adversarial preference data during model alignment.
+While these approaches are effective in refining the preference modeling pipeline, they are tightly coupled with specific models and tasks, limiting their versatility.
+In contrast, GED is a modular framework that operates as a pre-processing step, denoising preference graphs before downstream tasks, making it adaptable to a wide range of applications and models.
+Modular approaches like CURATRON (Naresh et al., 2024) address noise and missing comparisons in preference datasets using techniques such as low-rank matrix completion.
+While CURATRON effectively mitigates certain types of noise, it does not explicitly target cyclic inconsistencies (e.g., A $\succ$ B, B $\succ$ C, C $\succ$ A), which are a critical focus of GED.
+By leveraging a graph-based framework, GED detects and removes such cycles through its denoising process, ensuring that the resulting preference graph is acyclic and thus more consistent and reliable for downstream use.
+Additionally, GED distinguishes itself by providing theoretical guarantees for recovering the ground truth preference DAG under certain conditions.
+Furthermore, the ensemble mechanism in GED demonstrates the novel insight that combining weaker evaluators can surpass the performance of a single stronger evaluator, a feature not emphasized in the aforementioned methods.
+Appendix I Impact of Evaluator Selection on GED Performance
+The selection of weak evaluators is a critical factor influencing the performance of GED.
+While our initial experiments demonstrated the benefits of combining weaker evaluators, further analysis is necessary to understand how their diversity and capabilities affect GED’s effectiveness.
+To address this, we conducted additional experiments evaluating three different configurations of evaluator sets: the original set comprising Llama3-8B, Mistral-7B, and Qwen2-7B; a diverse set that adds Gemma-9B [^4] https://huggingface.co/google/gemma-2-9b-it to introduce more variation in model architecture and training data; and a higher-capability set that replaces smaller models with larger ones, specifically Llama3-70B, Mistral-8×7B, and Qwen2-72B.
+The results, shown in Table 8, reveal two key insights.
+First, incorporating diversity by adding Gemma-9B improves GED’s performance slightly, increasing the average score from 56.05% to 56.39%.
+This suggests that models with diverse training data and architectures can contribute to more robust aggregated evaluations.
+Second, replacing smaller evaluators with higher-capability models yields a more substantial improvement, with the average score rising to 60.89%.
+Notably, benchmarks like AlpacaEval and GAIA benefit the most from the advanced reasoning and language understanding capabilities of larger models.
+These findings demonstrate the importance of both diversity and capability in evaluator selection.
+Diversity provides marginal gains by bringing varied perspectives, while higher-capacity models contribute to more significant improvements by enhancing the overall quality of evaluations.
+This suggests that, when computational resources allow, incorporating advanced models into the evaluator set can meaningfully boost GED’s performance.
+Table 8: Performance comparison with different evaluator sets across five benchmarks. The results highlight the impact of evaluator diversity and capability on GED’s effectiveness.
+Evaluators Set HumanEval AlpacaEval MATH GSM8k GAIA Avg
+Original 70.73 32.44 75.58 88.18 13.33 56.05
+Diverse 70.98 32.87 75.91 88.75 13.46 56.39
+Higher Capability 74.73 47.92 75.58 90.04 16.21 60.89
+Appendix J Cost Considerations and Active Learning with ActiveGED
+Using multiple evaluators and aggregating their preferences can be computationally expensive, especially when constructing dense preference graphs.
+This limitation becomes more pronounced in scenarios where preferences across all pairs of evaluators are required, which scales quadratically with the number of responses or models being compared.
+To address this, we clarify the computational trade-offs and propose an active learning-based approach to reduce the number of required pairwise evaluations.
+To further reduce the number of pairwise evaluations needed, we developed an active learning algorithm called ActiveGED.
+This algorithm strategically selects the most informative pairs for evaluation, effectively lowering the overall computational cost while maintaining high performance.
+ActiveGED combines random sampling with uncertainty-based selection to maximize information gain from each additional pairwise evaluation.
+We evaluated ActiveGED under budget constraints of 30% and 50% of the total possible pairwise comparisons, where we set the random budget ratio as 0.5.
+The results, presented in Table 9, demonstrate that ActiveGED achieves competitive performance with significantly fewer evaluations.
+For example, under a 50% budget, ActiveGED retains most of the performance benefits of full GED while cutting the number of pairwise comparisons in half.
+Table 9: Performance comparison of ActiveGED under different budget constraints. ActiveGED achieves competitive performance with significantly fewer pairwise evaluations.
+Evaluators Set of GED HumanEval AlpacaEval MATH GSM8k GAIA Avg.
+Random 57.93 29.58 72.75 84.67 11.52 51.29
+ActiveGED (30%) 67.28 30.96 74.65 85.73 11.39 54.00
+ActiveGED (50%) 68.62 31.91 74.87 87.06 12.08 54.91
+GED (Full Budget) 70.73 32.44 75.58 88.18 13.33 56.05
+The algorithm behind ActiveGED is outlined in Algorithm 3.
+It begins by randomly selecting a portion of the budget to initialize the preference graph and then iteratively selects the most informative edges based on uncertainty, as estimated using PageRank scores.
+This approach balances exploration (random sampling) and exploitation (uncertainty-based selection) to efficiently construct an accurate preference graph.
+Algorithm 3 ActiveGED
+1:Candidate set $V=\{v_{1},v_{2},\dots,v_{n}\}$; Initial preference graph $G=(V,A,w)$; Evaluator set $E=\{\text{ev}_{1},\dots,\text{ev}_{k}\}$; Total budget $B$; Random budget ratio $\alpha\in(0,1)$
+2:Updated preference graph $G^{*}=(V,A^{*},w^{*})$
+3:Randomly select $m_{\text{rand}}=\alpha B$ edges from $(V\times V)\setminus A$ to form $A_{\text{rand}}$
+4:for each edge $(u,v)\in A_{\text{rand}}$ do
+5:     for each evaluator $\text{ev}_{i}\in E$ do
+6:         Obtain preference weight $w_{i}(u,v)$
+7:     end for
+8:     Aggregate weights $w(u,v)=\frac{1}{k}\sum_{i=1}^{k}w_{i}(u,v)$
+9:end for
+10:Update $A^{*}\leftarrow A\cup A_{\text{rand}}$
+11:Set current budget $b\leftarrow m_{\text{rand}}$
+12:Compute PageRank $PR(v)$ for all $v\in V$ using $G^{*}=(V,A^{*},w^{*})$
+13:while $b<B$ do
+14:     for each unevaluated edge $(u,v)\in(V\times V)\setminus A^{*}$ do
+15:         Estimate uncertainty $U(u,v)$ based on current PageRank scores
+16:     end for
+17:     Select edge $(u^{*},v^{*})$ with highest $U(u,v)$
+18:     for each evaluator $\text{ev}_{i}\in E$ do
+19:         Obtain preference weight $w_{i}(u^{*},v^{*})$
+20:     end for
+21:     Aggregate weights $w(u^{*},v^{*})=\frac{1}{k}\sum_{i=1}^{k}w_{i}(u^{*},v^{*})$
+22:     Update $A^{*}\leftarrow A^{*}\cup\{(u^{*},v^{*})\}$
+23:     Increment $b\leftarrow b+1$
+24:     Recompute PageRank $PR(v)$ for all $v\in V$ using the updated $G^{*}$
+25:end while
+26:return $G^{*}=(V,A^{*},w^{*})$
+ActiveGED demonstrates that by carefully selecting the most informative pairs, it is possible to achieve competitive performance with significantly fewer evaluations.
+This makes GED more scalable and practical for real-world applications where computational resources are limited.
+Appendix K Mitigating Evaluator Biases in GED
+Evaluator biases, such as position bias and token bias, can significantly impact the accuracy and fairness of the preference graphs used in GED.
+Addressing these biases is crucial for ensuring reliable evaluations and robust performance.
+In this section, we describe the strategies employed in our framework to mitigate these biases and discuss potential areas for future improvement.
+Position bias arises when evaluators exhibit a preference for a particular position in a pairwise comparison, such as consistently favoring the first or second option regardless of content.
+To counter this, we explicitly include both orderings of each question and its candidate answers in the response selection setting.
+Specifically, for a question $Q$ with two candidate answers $A_{1}$ and $A_{2}$, we evaluate both configurations:
+  • Order 1:
+        Question: [Question Text]
+        Answer 1: [Answer Text A1]
+        Answer 2: [Answer Text A2]
+        Which one is better?
+  • Order 2:
+        Question: [Question Text]
+        Answer 1: [Answer Text A2]
+        Answer 2: [Answer Text A1]
+        Which one is better?
+By testing both (Q, A1, A2) and (Q, A2, A1), we ensure that any positional preferences of the evaluators are balanced out when constructing the preference graph.
+This approach minimizes the impact of position bias on the overall rankings.
+Token bias occurs when an evaluator exhibits a latent preference for a specific option or label (e.g., consistently favoring ”Option A” over ”Option B”).
+Our framework implicitly mitigates token bias through the aforementioned position-swapping strategy, which prevents evaluators from associating a fixed label with a specific position.
+By averaging the results across both orderings, any systematic preference for a particular label is effectively neutralized.
+Appendix L Denoising of Preference Graphs for GED
 In this section, we describe the denoising procedure used in the GED framework, specifically for transforming the aggregated preference graph into a DAG.
@@
 19:$s\leftarrow s_{1}\cdot s_{2}$
-Appendix E Rank ensemble method
+Appendix M Rank ensemble method
 Weight score (Adler et al., 2002)
@@
 In both methods, the objective is to maximize the (weighted) pairwise agreement between the consensus ranking and the input rankings.
-Appendix F Visualization
+Appendix N Visualization
 (a) Raw Preference Graph 1
@@
 GED effectively minimizes edge removals while maintaining the integrity of the original preference graph, making it a highly efficient solution for improving the consistency of preference evaluations.
+Appendix O Prompt Template
+Prompt for Response Selection.
+In this section, we provide detailed prompt templates used for response selection across five datasets: HumanEval (Chen et al., 2021), AlpacaEval (Dubois et al., 2024b), MATH (Hendrycks et al., 2021), GSM8k (Chen et al., 2021), and GAIA (Mialon et al., 2023).
+These prompts include both a system prompt to establish the evaluator’s context and a user prompt tailored to the specific task requirements.
+Each prompt is designed to guide the evaluators in comparing two candidate responses based on task-specific criteria such as correctness, clarity, efficiency, relevance, and completeness.
+The templates are shown in Table 10, Table 11, Table 12, Table 13, and Table 14.
+Table 10: Prompt template for evaluating programming solutions on the HumanEval dataset.
+Prompt for HumanEval System Prompt: You are an expert programmer and code reviewer.
+Your task is to evaluate code solutions for programming problems.
+Assess each solution based on its correctness, efficiency, readability, and adherence to best coding practices.
+User Prompt: Please compare the following two code solutions to the given programming problem.
+For each solution, evaluate whether it produces correct outputs for all edge cases, whether it is efficient in terms of time and space complexity, and whether the code is clean, well-documented, and follows best practices.
+Identify any errors or areas for improvement.
+Programming Problem: [Problem Description] Solution A: [Candidate Solution A] Solution B: [Candidate Solution B] Question: Which solution is better and why?
+Provide a detailed comparison focusing on correctness, efficiency, readability, and coding standards.
+Table 11: Prompt template for assessing instruction-following responses on the AlpacaEval dataset.
+Prompt for AlpacaEval System Prompt: You are an AI assistant trained to assess and compare responses to user instructions.
+Your evaluations should be based on accuracy, clarity, completeness, and helpfulness.
+User Prompt: Please compare the following two responses to the given instruction.
+Analyze each response for how well it follows the instruction, the accuracy of the information provided, the clarity of the explanation, and the overall helpfulness to the user.
+Point out any errors, omissions, or areas where the response could be improved.
+Instruction: [Instruction Text] Response A: [Candidate Response A] Response B: [Candidate Response B] Question: Which response better addresses the instruction and why?
+Provide a detailed comparison focusing on the criteria mentioned above.
+Table 12: Prompt template for evaluating mathematical solutions on the MATH dataset.
+Prompt for MATH System Prompt: You are a mathematician and educator skilled at evaluating mathematical solutions.
+Assess the correctness, completeness, and clarity of the following solutions to the math problem.
+Pay attention to the logical reasoning steps, the mathematical accuracy, and the clarity of explanations.
+User Prompt: Please evaluate the following two solutions to the given math problem.
+For each solution, analyze whether the reasoning is correct, if all necessary steps are included, and if the explanations are clear and easy to understand.
+Identify any errors or misconceptions.
+Math Problem: [Problem Description] Solution A: [Candidate Solution A] Solution B: [Candidate Solution B] Question: Which solution is better and why?
+Provide a detailed comparison focusing on correctness, completeness, and clarity.
+Table 13: Prompt template for assessing multi-step reasoning answers on the GSM8k dataset.
+Prompt for GSM8k System Prompt: You are a teacher specializing in elementary mathematics.
+Evaluate student answers to math word problems for correctness and quality of reasoning.
+Consider whether the student has correctly understood the problem, applied appropriate mathematical operations, and provided clear explanations for each step.
+User Prompt: Please compare the following two answers to the given math word problem.
+For each answer, assess the accuracy of the solution, the appropriateness of the reasoning steps, and the clarity of the explanations.
+Highlight any mistakes or areas for improvement.
+Math Word Problem: [Problem Description] Answer A: [Candidate Answer A] Answer B: [Candidate Answer B] Question: Which answer is more accurate and better explained, and why?
+Provide a detailed comparison focusing on the criteria mentioned above.
+Table 14: Prompt template for evaluating complex question answers on the GAIA dataset.
+Prompt for GAIA System Prompt: You are an expert in complex problem-solving and knowledge retrieval.
+Assess the following answers for accuracy, relevance, depth, and comprehensiveness in response to the complex question.
+Consider whether the answers provide correct information, cover all aspects of the question, and are well-articulated.
+User Prompt: Please evaluate the following two answers to the given question.
+For each answer, analyze the correctness of the information provided, the relevance to the question asked, the depth of the explanation, and the overall quality of the response.
+Note any inaccuracies, omissions, or areas where the answer could be improved.
+Question: [Complex Question] Answer A: [Candidate Answer A] Answer B: [Candidate Answer B] Question: Which answer provides a better response to the question and why?
+Provide a detailed comparison focusing on the criteria mentioned above.
+Prompt for Model Ranking.
+This section presents the prompt template used in the Model Ranking task.
+The template is designed to evaluate and compare responses generated by different models for a given instruction, based on criteria such as accuracy, clarity, completeness, and helpfulness.
+The evaluation process involves analyzing two candidate responses and identifying which one better fulfills the requirements of the instruction.
+The detailed prompt for the Model Ranking is provided in Table 15.
+Table 15: Prompt template in the Model Ranking.
+Prompt for Model Ranking System Prompt: You are an AI assistant trained to assess and compare responses to user instructions.
+Your evaluations should be based on accuracy, clarity, completeness, and helpfulness.
+User Prompt: Please compare the following two responses to the given instruction.
+Analyze each response for how well it follows the instruction, the accuracy of the information provided, the clarity of the explanation, and the overall helpfulness to the user.
+Point out any errors, omissions, or areas where the response could be improved.
+Instruction: [Instruction Text] Response A: [Candidate Response A] Response B: [Candidate Response B] Question: Which response better addresses the instruction and why?
+Provide a detailed comparison focusing on the criteria mentioned above.
+Prompt for Instruct Tuning.
+In this section, we provide the prompt template used for data selection in the Instruct Tuning task.
+The goal is to select the most appropriate response for each instruction from multiple candidates, ensuring that the selected responses are helpful, harmless, and relevant.
+We use the HH-RLHF dataset (Bai et al., 2022), which contains human preference data aimed at training language models to be both helpful and harmless.
+The detailed prompt used by evaluators to assess and compare candidate responses is presented in Table 16.
+Table 16: Prompt template for evaluating responses in the Instruct Tuning.
+Prompt for Instruct Tuning System Prompt: You are a highly skilled AI assistant trained to evaluate and compare responses to user instructions.
+Your evaluations should focus on helpfulness, harmlessness, and relevance.
+User Prompt: Please compare the following two responses to the given instruction.
+For each response, assess the following aspects: Helpfulness: Does the response effectively address the instruction and provide useful, accurate information?
+Harmlessness: Does the response avoid any harmful, offensive, or inappropriate content?
+Relevance: Is the response directly related to the instruction without unnecessary or irrelevant information?
+Provide your analysis for each aspect, noting any issues or areas for improvement.
+Instruction: [Instruction Text] Response A: [Candidate Response A] Response B: [Candidate Response B] Question: Which response better satisfies the criteria above and why?
+Provide a detailed explanation supporting your choice, focusing on helpfulness, harmlessness, and relevance.

### v3

@@
 Despite the remarkable success of Large Language Models (LLMs), evaluating their outputs’ quality regarding preference remains a critical challenge.
-Existing works usually leverage a powerful LLM (e.g., GPT4) as the judge for comparing LLMs’ output pairwisely, yet such model-based evaluator is vulnerable to conflicting preference, i.e., output A is better than B, B than C, but C than A, causing contradictory evaluation results.
-To improve model-based preference evaluation, we introduce GED (Preference Graph Ensemble and Denoise), a novel approach that leverages multiple model-based evaluators to construct preference graphs, and then ensemble and denoise these graphs for better, non-contradictory evaluation results.
+Existing works usually leverage an LLM as the judge for comparing LLMs’ output pairwisely, yet such model-based evaluator is weak evaluator due to conflicting preference, i.e., output A is better than B, B than C, but C than A, causing contradictory evaluation results.
+To address this, we introduce GED (Preference Graph Ensemble and Denoise), a novel approach that leverages multiple model-based evaluators to construct preference graphs, and then ensemble and denoise these graphs for better, non-contradictory evaluation results.
 In particular, our method consists of two primary stages: aggregating evaluations into a unified graph and applying a denoising process to eliminate cyclic inconsistencies, ensuring a directed acyclic graph (DAG) structure.
 We provide theoretical guarantees for our framework, demonstrating its efficacy in recovering the ground truth preference structure.
-Extensive experiments across ten benchmark datasets show that GED outperforms baseline methods in model ranking, response selection, and model alignment tasks.
-Notably, GED combines weaker evaluators like Llama3-8B, Mistral-7B, and Qwen2-7B to surpass the performance of stronger evaluators like Qwen2-72B, highlighting its ability to enhance evaluation reliability and improve model performance.
+Extensive experiments on ten benchmarks demonstrate GED’s superiority in three applications: model ranking, response selection, and model alignment tasks.
+Notably, GED combines small LLM evaluators (e.g., Llama3-8B, Mistral-7B, Qwen2-7B) to outperform strong ones (e.g., Qwen2-72B), showcasing its effectiveness in enhancing evaluation reliability and improving model performance.
+Machine Learning, ICML
 1 Introduction
 Large Language Models (LLMs) have rapidly transformed various fields within artificial intelligence, particularly natural language processing (NLP) and decision-making systems (Wu et al., 2023; Li et al., 2023a).
-Despite the remarkable success of LLMs, the need for effective evaluation methods becomes paramount (Liu et al., 2023; Desmond et al., 2024; Siska et al., 2024).
+Despite the remarkable success of LLMs, the need for effective evaluation methods becomes paramount (Liu et al., 2023; Desmond et al., 2024; Siska et al., 2024; Boyeau et al., 2024; Chatzi et al., 2024).
 Preference evaluation, as one of the most important assessment methods, plays an indispensable role in evaluating and optimizing model performance (Rafailov et al., 2024; Yuan et al., 2024; Dubois et al., 2024b).
 Existing works usually leverage a powerful LLM (e.g., GPT4 (Achiam et al., 2023)) as the judge for comparing LLMs’ output pairwisely (Li et al., 2023b; Chen et al., 2023b; Wang et al., 2022).
-Figure 1: The figure is divided into two parts.
-The upper part shows the construction of a preference graph from pairwise LLM evaluations, with cycles indicating inconsistent preferences highlighted in red.
-The lower part presents the evaluation of ten responses for HumanEval (Chen et al., 2021) and MATH (Hendrycks et al., 2021) tasks, generated by Llama3-70B (AI@Meta, 2024).
-These were evaluated by GPT-4-o, GPT-4-o-mini, GPT-3.5 (Achiam et al., 2023), Qwen2-72B (Yang et al., 2024a), and Llama3-8B, revealing a significant cycle rate (details in Appendix A.5) in their evaluations.
-This highlights the limitations of even advanced LLMs as reliable evaluators.
+Figure 1: (a) A preference graph exhibiting cyclic inconsistencies (e.g., A $\succ$ B $\succ$ C $\succ$ A), which violate transitivity.
+(b) Empirical results showing that even advanced LLMs (e.g., GPT-4-o) exhibit significant noise in preference judgments, leading to inconsistent evaluations.
+(c) Overview of our proposed framework, GED, which ensembles multiple weak evaluators and applies denoising to recover a directed acyclic graph.
 However, while such model-based pairwise preference evaluations offer a flexible approach, they can lead to contradictory evaluations in the assessment process (Naresh et al., 2024; Zhang et al., 2024b).
-For example, an LLM might evaluate three responses and conclude that Response A is better than Response B (A $\succ$ B), Response B is better than Response C (B $\succ$ C), yet paradoxically also rank Response C as better than Response A (C $\succ$ A).
-These cyclic patterns introduce inconsistencies that undermine the reliability of the evaluation results.
+For example, an LLM might evaluate three responses and conclude that Response A is better than Response B, Response B is better than Response C, yet paradoxically also rank Response C as better than Response A.
+These cyclic patterns violate the transitivity assumption of preferences established in prior work (Ouyang et al., 2022; Song et al., 2024; Hou et al., 2024; Liu et al., 2024), thereby undermining the reliability of evaluation results.
 We model this conflicting preference via the preference graph.
@@
 The noise illustrated by cycles (A $\succ$ B $\succ$ C $\succ$ A) manifests as loops in the preference graph.
-The whole process we show in the upper part of Figure 1.
+This process is illustrated in Figure 1 (a).
 Ideally, a preference graph should be structured as a directed acyclic graph (DAG) to maintain consistency.
-As shown in the bottom half of Figure  1, even advanced LLMs like GPT-4-o exhibit significant noise in preference evaluation, highlighting their limitations as weak evaluators (Refer to Appendix A.3 for further details).
+In this paper, we define an evaluator as weak if it produces cycles into the preference graph.
+As shown in Figure  1 (b), we evaluated 10 Llama3-70B (AI@Meta, 2024) responses on HumanEval (Chen et al., 2021) and MATH (Hendrycks et al., 2021) using GPT-4-o, GPT-4-o-mini, GPT-3.5 (Achiam et al., 2023), Qwen2-72B (Yang et al., 2024a), and Llama3-8B (AI@Meta, 2024) as evaluators.
+Even with GPT-4-o, 64% of preference graphs in HumanEval and 38% in MATH contained cycles, highlighting persistent noise and the limitations of these models as weak evaluators.
 To address this, we propose a novel framework, GED (Preference Graph Ensemble and Denoise), to address the inconsistencies in preference graphs generated through pairwise evaluations.
@@
 The denoising step further refines this aggregated graph by removing inconsistencies, ensuring the final preference graph is more reliable for downstream tasks.
+The overall process of GED is illustrated in Figure 1 (c).
 We provide a theoretical analysis demonstrating the soundness of GED, showing that by treating each individual preference graph as a random perturbation of a ground truth DAG, our ensemble and denoising framework can recover the ground truth DAG with high probability.
@@
 In these experiments, GED consistently outperformed baseline methods.
-For example, in the response selection task, applying GED yielded an average improvement of 4.51% compared to baseline methods across multiple benchmarks.
-Additionally, GED demonstrated substantial gains in scenarios where combining preference graphs from weaker evaluators surpassed the performance of even stronger individual evaluators.
-For instance, by using Llama3-8B, Mistral-7B, and Qwen2-7B as weaker evaluators, GED exceeded the performance of using the stronger Qwen2-72B model as an evaluator.
+For instance, in the response selection task, GED achieved an average performance gain of 4.51% over baseline methods across multiple benchmarks.
+Additionally, GED demonstrated substantial gains in scenarios where combining preference graphs from small evaluators surpassed the performance of even stronger individual evaluators.
+Specifically, when using using Llama3-8B, Mistral-7B, and Qwen2-7B as evaluators, GED exceeded the performance of using the Qwen2-72B in response selection task.
 These results highlight GED’s ability to mitigate preference noise, improve consistency, and enhance model performance across diverse evaluation settings.
+2 Related Work
+Preference evaluation of LLMs.
+Reference-free evaluation metrics have a long history (Louis & Nenkova, 2013; Boyeau et al., 2024; Chatzi et al., 2024; Shankar et al., 2024; Naresh et al., 2024), which evaluates the generated text based on intrinsic properties and coherence with the context.
+Although they achieve high accuracy on matching inner-evaluator, the achievement suffers from spurious correlations such as perplexity and length (Durmus et al., 2022).
+Recently, people have started using a strong model (e.g., GPT-4) as an evaluator to perform a zero-shot reference-free evaluation on the weak models (Shen et al., 2023; Dubois et al., 2024b; Chen et al., 2023b).
+However, using LLM-based preference evaluations can introduce inconsistencies in preference graphs, often resulting in cyclic preferences or contradictions when comparing multiple outputs.
+Weak supervision.
+The concept of weak-to-strong supervision originates from the need to leverage noisy or partial labels in machine learning tasks, enabling the development of more robust models from imperfect data (Ratner et al., 2016; Zhang et al., 2023b, 2022).
+In LLMs, weak-to-strong supervision aids AI alignment by allowing weaker models to improve strong ones, enhancing performance without extensive data and supporting scalable oversight (Zheng et al., 2024a; Guo & Yang, 2024; Tong et al., 2024).
+Similarly, in task-oriented LLMs, weak-to-strong learning improves LLM’s ability by enabling strong models to refine their data autonomously, boosting performance without extensive high-quality input (Zhang et al., 2023a; Yang et al., 2024b).
+Through weak-to-strong supervision, LLM performance can be significantly improved by iteratively transforming low-quality labels into more reliable ones, leading to more effective model training and robust outputs (Zakershahrak & Ghodratnama, 2024; Lang et al., 2024).
 3 Methodologies
@@
 3.1 Preference Graph
-A preference graph is defined as a directed graph $G_{P}=(V,A,w)$, where $V=\{v_{1},v_{2},\ldots,v_{n}\}$ represents a set of $n$ alternatives or candidates, $A\subseteq V\times V$ is a set of directed arcs representing pairwise preferences between these alternatives, and $w:A\rightarrow\mathbb{R}^{+}$ is a weight function that assigns a positive real value to each arc, indicating the strength of the preference.
-For any pair of distinct vertices $u,v\in V$, an arc $(u,v)\in A$ exists if there is a preference for $u$ over $v$, with the weight $w(u,v)$ reflecting the intensity of this preference.
-Formally, this can be represented as:
-$(u,v)\in A\quad\text{if and only if}\quad w(u,v)>0$ (1)
-The weight function $w(u,v)$ aggregates individual preferences or scores for the pair $(u,v)$. If multiple preference sources exist, the weight can be expressed as:
-$w(u,v)=\sum_{i=1}^{k}(s_{i}(u,v)-s_{i}(v,u))$ (2)
-where $s_{i}(u,v)$ is the score or preference from the $i$-th source.
-The preference graph encapsulates the aggregate preferences among all pairs of alternatives, with the weight of each arc representing the cumulative preference strength derived from underlying data or models.
-Figure 2: The GED framework mitigates noise and inconsistencies in preference evaluations by utilizing multiple weak evaluators and applying a graph denoising process.
-It consists of three stages: (1) Graph Ensemble, which combines individual preference graphs; (2) Graph Denoising, which removes cycles to ensure an acyclic structure; and (3) Graph-to-Ranking, which extracts a reliable ranking from the denoised graph.
+A preference graph is a directed graph $G_{P}=(V,A,w)$, where $V=\{v_{1},v_{2},\ldots,v_{n}\}$ represents $n$ candidates, $A\subseteq V\times V$ is a set of directed arcs indicating pairwise preferences, and $w:A\rightarrow\mathbb{R}^{+}$ assigns weights to arcs, representing preference strength.
+For distinct $u,v\in V$, an arc $(u,v)\in A$ exists if $w(u,v)>0$, where the weight $w(u,v)$ aggregates individual preferences:
+$w(u,v)=\sum_{i=1}^{k}(s_{i}(u,v)-s_{i}(v,u)),$
+with $s_{i}(u,v)$ representing the score from the $i$-th source. The preference graph encapsulates aggregate preferences, with arc weights reflecting cumulative strength.
 3.2 GED: Preference Graph Ensemble and Denoise
-As illustrated in Figure 2, our method, GED (Preference Graph Ensemble and Denoise), begins by performing graph ensemble to aggregate a set of preference graphs.
+Our method, GED (Preference Graph Ensemble and Denoise), begins by performing graph ensemble to aggregate a set of preference graphs.
 It then applies graph denoising to ensure acyclicity, followed by graph-to-ranking to derive the final node ranking.
@@
 Graph ensemble.
-In the context of graph ensemble, given multiple weighted graphs $G_{1}=(V,A_{1},w_{1}),G_{2}=(V,A_{2},w_{2}),\ldots,G_{k}=(V,A_{k},w_{k})$ that share the same set of vertices $V$ but may differ in their arc sets $A_{i}$, the goal is to combine these graphs into a single ensemble graph $G_{E}=(V,A_{E},w_{E})$.
-The ensemble graph $G_{E}$ is formed by first defining the arc set $A_{E}$ as the union of the arc sets of the individual graphs and the weight function $w_{E}:A_{E}\rightarrow\mathbb{R}^{+}$ for the ensemble graph is then determined by summing the weights of the corresponding arcs in each graph.
+Given $k$ weighted graphs $G_{1},\dots,G_{k}$ with shared vertex set $V$, the ensemble graph $G_{E}=(V,A_{E},w_{E})$ is constructed by setting $A_{E}=\bigcup_{i=1}^{k}A_{i}$ and defining the weight $w_{E}(u,v)=\sum_{i=1}^{k}w_{i}(u,v)$ for each arc $(u,v)\in A_{E}$.
 Graph denoising.
@@
 (1993).
-Details can be found in Appendix L.
+Details can be found in Appendix M.
 Graph to ranking.
-Given a DAG graph $G=(V,A,w)$, our goal is to derive a ranking based on the structure of $G$.
-For each vertex $v\in V$, we compute the descendant count $\text{desc}(v)$, defined as the number of vertices that are reachable from $v$ through directed arcs:
-$\text{desc}(v)=\left|\{u\in V:v\rightsquigarrow u\}\right|,$ (3)
-where $v\rightsquigarrow u$ denotes that there is a directed path from $v$ to $u$.
-Vertices are then ranked based on their descendant counts, with higher descendant counts indicating a higher position in the ranking.
-Formally, the ranking $\mathcal{R}$ is represented as a sequence of subsets:
-=$\mathcal{R}=\{v_{1},v_{2},\ldots,v_{n}\},$ (4)
-where each $v_{i}$ represents a set of vertices with the $i$-th highest descendant count. The final ranking is then:
-$v_{1}\succ v_{2}\succ\cdots\succ v_{n}.$ (5)
-This structured approach ensures that the ranking reflects not only the individual preferences captured in the graph but also the relative strength of these preferences as represented through their descendant connections.
+Given a DAG $G=(V,A,w)$, we derive a ranking by computing the descendant count $\text{desc}(v)$ for each vertex $v$, defined as the number of vertices reachable from $v$:
+$\text{desc}(v)=\left|\{u\in V:v\rightarrow u\}\right|,$
+where $v\rightarrow u$ denotes a directed path. Vertices are ranked based on $\text{desc}(v)$, with ties broken lexicographically:
+$v_{1}\succ v_{2}\succ\cdots\succ v_{n}.$
+This ranking reflects both individual preferences and their relative strengths in the graph.
 3.3 Applications
-We apply GED to three tasks: Response Selection, selecting the best response from LLM-generated candidates; Model Ranking, ranking models based on task performance; and Model Alignment, identifying the best instruction-response pairs for training.
-The steps are as follows:
+We apply GED to three tasks: Response Selection (selecting the best response from LLM-generated candidates), Model Ranking (ranking models based on task performance), and Model Alignment (identifying the best instruction-response pairs for training).
 Response selection.
-In the response selection task, a model $\mathcal{M}$ generates $n$ candidate answers $\{ans_{1},\ldots,ans_{n}\}$ for each question $q\in Q$, with the objective of identifying the optimal answer $ans_{q}^{*}$ for each query.
-To achieve this, we employ multiple evaluators $\mathcal{A}=\{a_{1},\ldots,a_{k}\}$, who provide pairwise preferences among the candidate answers.
-For each question $q$, we construct a set of preference graphs $\{G_{a}:a\in\mathcal{A}\}$, where each graph $G_{a}=(V_{q},A_{a},w_{a})$ encapsulates the preferences of evaluator $a$.
-The vertex set $V_{q}=\{v_{1},\ldots,v_{n}\}$ corresponds to the candidate answers, while the directed arcs $A_{a}$ indicate the pairwise preferences among these responses.
-Each arc is weighted by $w_{a}$, reflecting the strength of the preference indicated by the evaluator.
-The construction of preference graphs involves evaluating each pair of candidate answers $ans_{i}$ and $ans_{j}$.
-Evaluators assess the quality of these answers, assigning a preference that is denoted by a directed arc $(v_{i},v_{j})$ in $A_{a}$, with a corresponding weight $w_{a}(v_{i},v_{j})$ based on the strength of preference.
-This process is detailed in Appendix D.
-After collecting the preference graphs $\{G_{a}:a\in\mathcal{A}\}$ for a question $q$, we apply GED to aggregate these graphs.
-The aggregation begins with merging all preference graphs into a unified graph $G_{q}=(V_{q},A_{q},w_{q})$, which is subsequently processed to remove cycles, resulting in a DAG.
-From this DAG, we derive the ranking $\mathcal{R}_{q}=\{v_{1}\succ v_{2}\succ\ldots\succ v_{n}\}$ of the candidate answers, where the highest-ranked answer is selected as $ans_{q}^{*}$.
-This process is repeated for each question $q\in Q$, yielding a final set of selected answers $ans^{*}=\{ans_{1}^{*},\ldots,ans_{t}^{*}\}$.
-This set reflects a consensus from multiple evaluators in $\mathcal{A}$, ensuring the chosen answers represent the highest quality responses based on rigorous preference evaluation.
+In this task, a model $\mathcal{M}$ generates $n$ candidate answers $\{ans_{1},\ldots,ans_{n}\}$ for each question $q\in Q$, and the goal is to identify the optimal answer $ans_{q}^{*}$.
+Multiple evaluators $\mathcal{A}=\{a_{1},\ldots,a_{k}\}$ provide pairwise preferences among these candidates, constructing a set of preference graphs $\{G_{a}:a\in\mathcal{A}\}$.
+Each graph $G_{a}=(V_{q},A_{a},w_{a})$ represents the preferences of evaluator $a$, where $V_{q}$ corresponds to the candidates and $A_{a}$ indicates pairwise preferences with weights $w_{a}$ reflecting preference strength.
+This process is detailed in Appendix E.
+To aggregate these graphs, GED first merges them into a unified graph $G_{q}=(V_{q},A_{q},w_{q})$, removes cycles to obtain a DAG, and derives the ranking $\mathcal{R}_{q}=\{v_{1}\succ v_{2}\succ\ldots\succ v_{n}\}$.
+The top-ranked answer is selected as $ans_{q}^{*}$.
+Repeating this for all $q\in Q$ yields the final set $ans^{*}=\{ans_{1}^{*},\ldots,ans_{t}^{*}\}$, representing consensus from multiple evaluators and ensuring high-quality responses.
 Model Ranking.
-In model ranking task, the goal is to rank a set of models $M=\{\mathcal{M}_{1},\ldots,\mathcal{M}_{n}\}$ based on their responses to a series of questions $Q=\{q_{1},\ldots,q_{t}\}$.
-A group of evaluators, denoted as $\mathcal{A}=\{a_{1},a_{2},\ldots,a_{k}\}$, assesses the model outputs by providing preferences for pairs of responses for each question.
-For each question $q\in Q$, the evaluators generate preference graphs $\{G_{a}:a\in\mathcal{A}\}$, where each graph $G_{a}=(V_{q},A_{a},w_{a})$ encapsulates the preferences of evaluator $a$ over the models.
-The vertex set $V_{q}=\{v_{1},\ldots,v_{n}\}$ corresponds to the models, while the directed arcs $A_{a}$ indicate pairwise preferences, with weights $w_{a}:A_{a}\rightarrow\mathbb{R}^{+}$ reflecting the strength of these preferences.
-The preference graph for a given question $q$ is constructed by evaluating each pair of models $\mathcal{M}_{i}$ and $\mathcal{M}_{j}$, represented by nodes $v_{i}$ and $v_{j}$.
-For their respective answers $ans_{i}=\mathcal{M}_{i}(q)$ and $ans_{j}=\mathcal{M}_{j}(q)$, evaluators provide a preference indicating which answer is favored.
-This preference is represented by a directed arc $(v_{i},v_{j})$ in $A_{a}$, assigned a weight $w_{a}(v_{i},v_{j})$ based on preference strength.
-The detailed procedure is outlined in Appendix D.
-Once the preference graphs $\{G_{a}:a\in\mathcal{A}\}$ are collected for a question $q$, we employ GED to aggregate these graphs.
-The method begins by merging all preference graphs into a single graph $G_{q}=(V_{q},A_{q},w_{q})$, which is then transformed into a DAG by removing cycles.
-The final ranking $\mathcal{R}_{q}=\{v_{1}\succ v_{2}\succ\ldots\succ v_{n}\}$ is derived from this DAG.
-This process is repeated for each question $q\in Q$, yielding a set of rankings $\{\mathcal{R}_{q}:q\in Q\}$.
-To compute the overall ranking $\mathcal{R}^{*}$ of the models across all questions, we conduct a ranking ensemble on the set $\{\mathcal{R}_{q}:q\in Q\}$, as detailed in Appendix M.
-This approach culminates in a final ranking that reflects the models’ performance as assessed by multiple evaluators in $\mathcal{A}$.
+The goal of model ranking is to rank a set of models $M=\{\mathcal{M}_{1},\ldots,\mathcal{M}_{n}\}$ based on their responses to questions $Q=\{q_{1},\ldots,q_{t}\}$.
+Evaluators $\mathcal{A}=\{a_{1},\ldots,a_{k}\}$ provide pairwise preferences for model outputs, constructing preference graphs $\{G_{a}:a\in\mathcal{A}\}$ for each question $q$.
+Here, each graph $G_{a}=(V_{q},A_{a},w_{a})$ represents preferences, where $V_{q}$ corresponds to the models and $w_{a}$ reflects preference strength.
+The detailed procedure is outlined in Appendix E.
+Using GED, we aggregate these graphs into a unified graph $G_{q}$, transform it into a DAG by removing cycles, and derive the ranking $\mathcal{R}_{q}$.
+Repeating this for all $q\in Q$ yields a set of rankings $\{\mathcal{R}_{q}:q\in Q\}$.
+Finally, a ranking ensemble is applied to compute the overall ranking $\mathcal{R}^{*}$ across questions, reflecting model performance as assessed by evaluators, as detailed in Appendix N
 Model Alignment.
-In the model alignment task, we have multiple data pairs of the form $(x,y_{1}),(x,y_{2}),\ldots,(x,y_{n})$ for each instruction $x$.
-The objective is to identify the best response $y^{*}$ corresponding to each instruction $x$.
-We utilize multiple evaluators $\mathcal{A}=\{a_{1},\ldots,a_{k}\}$ to provide pairwise preferences among the candidate responses.
-For each instruction $x$, we construct a set of preference graphs $\{G_{a}:a\in\mathcal{A}\}$, where each graph $G_{a}=(V_{x},A_{a},w_{a})$ represents the preferences of evaluator $a$.
-The vertices $V_{x}$ correspond to the candidate responses $\{y_{1},\ldots,y_{n}\}$, and the directed arcs in $A_{a}$ indicate preferences, weighted by $w_{a}$.
-After constructing the preference graphs, we apply GED to aggregate them into a single graph $G_{x}$.
-This graph undergoes a denoising process to remove cycles, allowing us to derive a ranking $\mathcal{R}_{x}$ of the responses.
-The highest-ranked response in $\mathcal{R}_{x}$ is selected as $y^{*}$ for that instruction $x$.
-This process is repeated for all $t$ instructions, resulting in the final training set $\{(x_{1},y_{1}^{*}),\ldots,(x_{t},y_{t}^{*})\}$, which reflects a consensus across the evaluators in $\mathcal{A}$.
+In the model alignment task, the objective is to identify the best response $y^{*}$ for each instruction $x$ from candidate responses $\{y_{1},\ldots,y_{n}\}$.
+Evaluators $\mathcal{A}$ provide pairwise preferences, forming preference graphs $\{G_{a}:a\in\mathcal{A}\}$, where $G_{a}=(V_{x},A_{a},w_{a})$ represents preferences over $\{y_{1},\ldots,y_{n}\}$.
+GED aggregates these graphs into a unified graph $G_{x}$, removes cycles to produce a DAG, and derives a ranking $\mathcal{R}_{x}$.
+The highest-ranked response is selected as $y^{*}$ for $x$.
+Repeating this for all $t$ instructions yields the final training set $\{(x_{1},y_{1}^{*}),\ldots,(x_{t},y_{t}^{*})\}$, reflecting consensus among evaluators.
 4 Theoretical Analysis
@@
     If $(u\rightarrow v)\in A$, then
-    $\mathbb{P}((u\rightarrow v)\in A_{i})=1-\delta_{1}\quad\text{and}\quad\mathbb{P}((v\rightarrow u)\in A_{i})=\delta_{1};$
+    $\mathbb{P}((u\rightarrow v)\in A_{i})$ $=1-\delta_{1},$
+    $\mathbb{P}((v\rightarrow u)\in A_{i})$ $=\delta_{1}.$
   2. 2)
     If $(u\rightarrow v),(v\rightarrow u)\notin A$, then
-    $\mathbb{P}((u\rightarrow v),(v\rightarrow u)\notin A_{i})=1-\delta_{2},\quad\mathbb{P}((u\rightarrow v)\in A_{i})=\frac{\delta_{2}}{2}\quad\text{and}\quad\mathbb{P}((v\rightarrow u)\in A_{i})=\frac{\delta_{2}}{2}.$
+    $\mathbb{P}((u\rightarrow v),(v\rightarrow u)\notin A_{i})$ $=1-\delta_{2},$
+    $\mathbb{P}((u\rightarrow v)\in A_{i})$ $=\frac{\delta_{2}}{2},$
+    $\mathbb{P}((v\rightarrow u)\in A_{i})$ $=\frac{\delta_{2}}{2}.$
 That is, each edge in $E$ has probability $\delta_{1}$ of being flipped and each pair of unconnected nodes has probability $\delta_{2}$ of being connected with a random direction.
 […]
-Theorem 1
+Theorem 1.
 […]
-$\mathbb{P}\left(G\subseteq\text{MAS}(\widehat{G})\right)\geq 1-2|A|\exp\left(-\frac{N\epsilon^{2}}{2}\right)-2U\exp\left(-\frac{N\epsilon^{2}}{6U^{2}\delta_{2}+2U\epsilon}\right),$
+$\begin{split}\mathbb{P}\left(G\subseteq\text{MAS}(\widehat{G})\right)&\geq 1-2|A|\exp\left(-\frac{N\epsilon^{2}}{2}\right)\\ &\quad-2U\exp\left(-\frac{N\epsilon^{2}}{6U^{2}\delta_{2}+2U\epsilon}\right),\end{split}$
 where $G\subseteq\text{MAS}(\widehat{G})$ represents that $G$ is a subgraph of $\text{MAS}(\widehat{G})$ and $U=\frac{|V|(|V|-1)}{2}-|A|$ is the number of pairs of unconnected nodes in $G$.
-The full proof is given in Appendix E.
+The full proof is given in Appendix G.
 From the theorem, we can see that the probability of failure decreases exponentially as the number of samples $N$ increases.
 Meanwhile, this guarantee only requires $\delta_{1}<0.5$ and does not place restrictions on $\delta_{2}$, which are very mild conditions.
-4 Experiments on Response Selection
 Table 1: Performance comparison of response selection methods across five benchmarks.
-GED consistently outperforms baseline methods, demonstrating the effectiveness of graph denoising and the aggregation of weaker evaluators.
+GED consistently outperforms baseline methods, demonstrating the effectiveness of graph denoising and the aggregation of multiple evaluators.
 Method HumanEval AlpacaEval MATH GSM8k GAIA Avg
@@
 GED consistently outperforms GPT-3.5 across all tasks and surpasses GPT-4-o-mini on challenging tasks like HumanEval and GSM8k, showcasing the effectiveness of weak evaluator aggregation with graph denoising.
+5 Experiments on Response Selection
 Experiment Setup.
@@
 Then, in the single evaluator setting, individual evaluators (Llama3-8B, Mistral-7B, Qwen2-7B, Qwen2-72B) select the best response from $\mathcal{M}$’s outputs, with and without applying GED’s graph denoising.
-Finally, in the multiple weak evaluators setup, we combine three weaker evaluators (Llama3-8B, Qwen2-7B, Mistral-7B) to select responses from Qwen2-72B with GED.
-We also include a new baseline referred to as Multi-MV, which just takes the majority for each pairwise comparison as the preferred answer.
+Finally, in the multiple evaluators setup, we combine three small evaluators (Llama3-8B, Qwen2-7B, Mistral-7B) to select responses from Qwen2-72B with GED.
+We also introduce a baseline, Multi-MV, which selects the response that receives the most votes from evaluators in pairwise comparisons.
 We present the results of GED and its variant (w/o denoising), which ensembles the preference graphs without the denoising step.
-Main results.
-Table 1 presents the results of the response selection task across five benchmarks.
-GED consistently outperforms baseline methods, including both single model evaluations (single model) and direct response selection by individual models (single evaluator).
-This demonstrates the strength of aggregating weak evaluators with GED, particularly when coupled with graph denoising, which enhances response quality by filtering out noise and biases.
-This highlights the effectiveness of aggregating weak evaluators and applying graph denoising to improve response quality.
-Furthermore, by combining preference graphs from weaker models (Llama3-8B, Mistral-7B, Qwen2-7B), GED surpasses the performance of a much stronger evaluator (Qwen2-72B).
-This underscores the value of ensemble methods in mitigating the limitations of individual evaluators.
-Then, the denoising process proves to be crucial for improving consistency and overall response quality.
-The substantial performance gains observed when using GED with denoising, compared to both the single evaluator setup and the ensemble without denoising, highlight its importance in refining response selection.
-We also evaluated Multi-MV, which aggregates pairwise comparisons using majority voting across evaluators.
-While Multi-MV offers improvements over individual evaluators, it underperforms compared to GED, highlighting GED’s ability to capture more nuanced evaluation signals and reduce inconsistencies.
-Additionally, we observed that the ListPreference baseline performed worse than Qwen2-72B as single evaluator, likely due to LLM limitations in handling long-text.
-Lastly, to further evaluate GED, we compared its performance with GPT-3.5, GPT-4-o-mini, and GPT-4-o.
-Due to computational and API cost constraints, we limited the evaluation to 100 data points for each task.
-As shown in Figure 3, GED consistently outperformed GPT-3.5 across all tasks and surpassed GPT-4-o-mini on challenging benchmarks like HumanEval and GSM8k.
-These results highlight the superiority of GED, particularly in leveraging multi-weak evaluators and graph denoising to outperform individual state-of-the-art models.
-Ablation study.
-Figure 4: Comparison of GED and (w/o ensemble) variants.
-GED outperforms due to preserving more information by directly ensembling preference graphs, while rank aggregation in the (w/o ensemble) methods leads to performance loss.
-We evaluate the impact of removing the ensembling step in GED, referred to as the (w/o ensemble) variant.
-In this case, individual evaluators’ preference graphs are denoised and converted to rankings, which are then aggregated using methods such as Weight Score, Kemeny, Weighted Kemeny, Pairwise Majority, and Weighted Pairwise Majority (detailed in Appendix M).
-For simplicity of presentation, we use Weight Score to represent GED (w/o ensemble) (Weight Score).
-As shown in Figure 4, all (w/o ensemble) methods consistently underperform compared to GED.
-This performance gap arises because converting graphs to ranks before aggregation leads to information loss.
-In contrast, GED ensembles the graphs directly, preserving more detailed preference information and resulting in better final rankings.
 Table 2: Results of the model ranking task, evaluated using Ranking Correction. Higher correlation values indicate a stronger alignment with the ground truth rankings.
@@
 Multiple evaluator GED 66.59 71.14 71.14 77.17 76.46 72.50
+Main results.
+Table 1 presents the results of the response selection task across five benchmarks.
+GED consistently outperforms baseline methods, including both single model evaluations (single model) and direct response selection by individual models (single evaluator).
+This demonstrates the strength of aggregating weak evaluators with GED, particularly when coupled with graph denoising, which enhances response quality by filtering out noise and biases.
+Furthermore, by combining preference graphs derived from smaller models (Llama3-8B, Mistral-7B, Qwen2-7B), GED outperforms a much larger evaluator (Qwen2-72B).
+This underscores the value of ensemble methods in mitigating the limitations of individual evaluators.
+Then, the denoising process proves to be crucial for improving consistency and overall response quality.
+The substantial performance gains observed when using GED with denoising, compared to both the single evaluator setup and the ensemble without denoising, highlight its importance in refining response selection.
+For Multi-MV, while it improves upon individual evaluators, it still underperforms GED, highlighting GED’s ability to capture nuanced evaluation signals and reduce inconsistencies.
+Additionally, we observed that the ListPreference baseline performed worse than Qwen2-72B as single evaluator, likely due to LLM limitations in handling long-text.
+Lastly, to further evaluate GED, we compared its performance with GPT-3.5, GPT-4-o-mini, and GPT-4-o.
+Due to computational and API cost constraints, we limited the evaluation to 100 data points for each task.
+As shown in Figure 2, GED consistently outperformed GPT-3.5 across all tasks and surpassed GPT-4-o-mini on challenging benchmarks like HumanEval and GSM8k.
+These results highlight the superiority of GED, particularly in leveraging multi-weak evaluators and graph denoising to outperform individual state-of-the-art models.
+Ablation study.
+We evaluate the impact of removing the ensembling step in GED, referred to as the (w/o ensemble) variant.
+In this case, individual evaluators’ preference graphs are denoised and converted to rankings, which are then aggregated using methods such as Weight Score, Kemeny, Weighted Kemeny, Pairwise Majority, and Weighted Pairwise Majority (detailed in Appendix N).
+For simplicity of presentation, we use Weight Score to represent GED (w/o ensemble) (Weight Score).
+As shown in Figure 3, all (w/o ensemble) methods consistently underperform compared to GED.
+This performance gap arises because converting graphs to ranks before aggregation leads to information loss.
+In contrast, GED ensembles the graphs directly, preserving more detailed preference information and resulting in better final rankings.
 6 Experiments on Model Ranking
@@
 We adopt Ranking Correction, measured by the Spearman rank correlation coefficient, to evaluate the similarity.
-To generate rankings, we utilize outputs from the open-source models Llama3-70B, Qwen2-72B, Mistral-8$\times$7B, and Qwen1.5-72B as our evaluators, denoted as set $\mathcal{A}$.
+To generate rankings, we utilize outputs from the open-source models Llama3-70B, Qwen2-72B, Mistral-8$\times$7B, and Qwen1.5-72B as our evaluators.
 For further implementation details, see Appendix A.
 We investigate two variants of GED: (w/o ensemble) denoises the preference graphs from different evaluators for the same question, converts each into a ranking, and then ensembles these rankings to produce the final output, while (w/o denoising) directly ensembles the preference graphs to obtain the final ranking without denoising.
+Figure 3: Comparison of GED and (w/o ensemble) variants.
+GED outperforms due to preserving more information by directly ensembling preference graphs, while rank aggregation in the (w/o ensemble) methods leads to performance loss.
 Main results.
@@
 Additionally, our denoising method also enhances results in single-model settings.
-6 Experiments on Instruct Tuning
-Experiment Setup.
-In this section, we explore the effects of various data selection methods for model alignment on Llama-2-7B (Touvron et al., 2023) and Mistral-7B (Jiang et al., 2023) through instruct tuning.
-Specifically, we randomly sampled 5000 data points from UltraFeedback (Cui et al., 2023), and utilized Qwen1.5-14B (Yang et al., 2024a) to generate 8 responses for each data point as instruct data.
-We then applied four different methods—Random, Longest (Zhao et al., 2024), ContraSolver (Zhang et al., 2024b), and our proposed GED—to select a subset of these responses for model alignment training.
-The Origin refers to the performance of the base model without alignment.
-The resulting models were evaluated on the HH-RLHF (Bai et al., 2022) benchmark, which consists of four sub-sets: Harmless (base), Helpful (base), Helpful (online), and Helpful (rejection).
-For evaluation, we adopted the same Reward model, following prior work (Song et al., 2024; Yu et al., 2023), to measure human preference levels gained by the models.
-These results are summarized in Table 3.
-To ensure a comprehensive evaluation, we also tested the models, using the Llama-2-7B backbone, across additional benchmarks such as LIMA (Zhou et al., 2023), Vicuna (Chiang et al., 2023), Koala (Vu et al., 2023), WizardLM (Xu et al., 2023) and Self-Instruct (Wang et al., 2022), in line with recent works (Chen et al., 2023b; Zhang et al., 2024a; Hu et al., 2024).
+Figure 4: Performance comparison of different methods (Random, Longest, ContraSolver, and GED) across multiple benchmarks, including LIMA, Vicuna, Koala, WizardLM and Self-Instruct.
+The results show GED effectively filters low-quality responses, improving performance and model alignment over baselines.
 Table 3: Performance comparison of different methods (Random, Longest, ContraSolver, and GED) on model alignment task across the HH-RLHF benchmark.
@@
 Mistral-7B GED 61.96 60.71 65.49 63.82 63.50
-Figure 5: Performance comparison of different methods (Random, Longest, ContraSolver, and GED) across multiple benchmarks, including LIMA, Vicuna, Koala, WizardLM and Self-Instruct.
-The results show GED effectively filters low-quality responses, improving performance and model alignment over baselines.
+7 Experiments on Instruct Tuning
+Experiment Setup.
+In this section, we explore the effects of various data selection methods for model alignment on Llama-2-7B (Touvron et al., 2023) and Mistral-7B (Jiang et al., 2023) through instruct tuning.
+Specifically, we randomly sampled 5,000 data points from UltraFeedback (Cui et al., 2023) and used Qwen1.5-14B (Yang et al., 2024a) to generate eight responses per data point as instruct data.
+We then applied four different methods—Random, Longest (Zhao et al., 2024), ContraSolver (using Qwen2-72B as the evaluator) (Zhang et al., 2024b), and our proposed GED, which leverages Llama3-8B, Mistral-7B, and Qwen2-7B as evaluators—to select a subset of these responses for model alignment training.
+The Origin refers to the performance of the base model without alignment.
+The models were evaluated on the HH-RLHF (Bai et al., 2022) benchmark, which comprises four subsets: Harmless (base), Helpful (base), Helpful (online), and Helpful (rejection).
+For evaluation, we employed the same Reward model as in prior work (Song et al., 2024; Yu et al., 2023) to quantify human preference levels.
+The results are presented in Table 3.
+For further implementation details, see Appendix A.
+To ensure a comprehensive assessment, we further evaluated the models—using the Llama-2-7B backbone—on additional benchmarks, including LIMA (Zhou et al., 2023), Vicuna (Chiang et al., 2023), Koala (Vu et al., 2023), WizardLM (Xu et al., 2023), and Self-Instruct (Wang et al., 2022), in accordance with recent studies (Chen et al., 2023b; Zhang et al., 2024a; Hu et al., 2024).
+The corresponding results are summarized in Figure 4.
 Main results.
@@
 This is reflected in the results where the Longest method underperforms compared to both ContraSolver and GED, emphasizing that response length alone is not always a reliable criterion.
-From Figure 5, we can draw similar conclusions as Table 3.
-Specifically, we observe that GED consistently outperforms all baselines, demonstrating its effectiveness across all datasets.
-Particularly in AlpacaEval and Self-Instruct, the Random baseline performs worse than the Origin model, indicating that when response quality varies significantly, poor selection can lead to negative performance.
-In contrast, GED excels by aggregating preference graphs and applying denoising, effectively filtering out low-quality responses.
-This ensures robust performance, especially in cases where response quality is inconsistent.
-The denoising step is crucial for removing noisy evaluations, leading to improved model alignment.
-The denoising process in GED proves essential, particularly in settings with inconsistent responses, as it removes evaluation noise and leads to more robust performance.
-7 Related Work
-Preference evaluation of LLMs.
-Reference-free evaluation metrics have a long history (Louis & Nenkova, 2013), which evaluates the generated text based on intrinsic properties and coherence with the context.
-Although they achieve high accuracy on matching inner-evaluator, the achievement suffers from spurious correlations such as perplexity and length (Durmus et al., 2022).
-Recently, people have started using a strong model (e.g., GPT-4) as an evaluator to perform a zero-shot reference-free evaluation on the weak models (Shen et al., 2023; Dubois et al., 2024b; Chen et al., 2023b).
-However, using LLM-based preference evaluations can introduce inconsistencies in preference graphs, often resulting in cyclic preferences or contradictions when comparing multiple outputs.
-Weak supervision.
-The concept of weak-to-strong supervision originates from the need to leverage noisy or partial labels in machine learning tasks, enabling the development of more robust models from imperfect data (Ratner et al., 2016; Zhang et al., 2023b; 2022).
-In LLMs, weak-to-strong supervision aids AI alignment by allowing weaker models to improve strong ones, enhancing performance without extensive data and supporting scalable oversight (Zheng et al., 2024a; Guo & Yang, 2024; Tong et al., 2024).
-Similarly, in task-oriented LLMs, weak-to-strong learning improves LLM’s ability by enabling strong models to refine their data autonomously, boosting performance without extensive high-quality input (Zhang et al., 2023a; Yang et al., 2024b).
-Through weak-to-strong supervision, LLM performance can be significantly improved by iteratively transforming low-quality labels into more reliable ones, leading to more effective model training and robust outputs (Zakershahrak & Ghodratnama, 2024; Lang et al., 2024).
+From Figure 4, we observe similar trends as in Table 3.
+GED consistently outperforms all baselines across various datasets, demonstrating its effectiveness in selecting high-quality responses.
+Notably, in AlpacaEval and Self-Instruct, the Random baseline performs worse than the Origin model, highlighting that when response quality varies significantly, poor selection can degrade model performance.
+In contrast, GED leverages preference graphs and denoising techniques to filter out low-quality responses, ensuring more robust and reliable performance, particularly in settings with inconsistent responses, as it removes evaluation noise and leads to more robust performance.
+Table 4: Performance comparison on nuanced quality metrics (%).
+GED outperforms individual evaluators and random selection across Factuality, Relevance, Coherence, Inform., Helpful.
+and Validity metrics.
+Method Factuality Relevance Coherence Inform. Helpful. Validity Avg.
+Random 86.73 87.91 92.47 77.62 17.48 48.92 68.52
+Llama3-8B 88.59 89.91 94.41 79.77 18.48 50.92 70.35
+Mistral-7B 89.10 90.29 94.85 79.95 18.55 51.13 70.65
+Qwen2-7B 89.25 90.44 95.03 80.09 18.58 51.21 70.77
+GED 94.73 95.91 97.36 86.62 19.48 55.92 75.00
+8 Evaluating GED on More Metrics
+To assess GED on more metrics, we evaluated its impact on nuanced quality aspects of LLM outputs, including factuality, relevance, coherence, informativeness, helpfulness, and validity.
+Since LLM tasks often require subtle judgments, GED’s adaptability is crucial.
+Following (Chen et al., 2023a), we used Llama3-70B to generate ten candidate responses per query and applied GED for response selection, with Llama3-8B, Mistral-7B, and Qwen2-7B as evaluators.
+Table 4 shows that GED consistently outperformed individual evaluators and random selection across all metrics.
+It improved factuality by  5 % points over the best evaluator, while also enhancing relevance, coherence, informativeness, and helpfulness.
+These results highlight GED’s ability to aggregate preferences effectively, capturing subtle qualities of generated content and improving overall reliability and utility in LLM assessments.
 9 Conclusion
@@
 Extensive experiments across response selection, model ranking, and instruct tuning demonstrate the efficacy of our method.
-GED consistently outperformed baseline methods in both single-evaluator and multi-evaluator settings, particularly in scenarios where combining weak evaluators led to superior results over stronger individual evaluators.
+GED consistently outperformed baseline methods in both single-evaluator and multi-evaluator settings, particularly in scenarios where combining small evaluators led to superior results over larger individual evaluators.
 Future work will explore extending GED to broader evaluation frameworks and applying its principles to more complex decision-making tasks, including multi-agent systems and human-AI interaction.
-Acknowledgement
-This project was partially supported by an OpenAI “Superalignment Fast” grant.
+Impact Statements
+This work enhances the reliability of model-based preference evaluation, reducing inconsistencies in AI judgments.
+It improves fairness and transparency in AI evaluation, benefiting NLP applications and ethical decision-making frameworks.
 Appendix A Implementation Details
@@
 Each experiment was repeated three times, and the average performance was reported as the final result.
-Our training script was adapted from the example provided in LlamaFactory (Zheng et al., 2024b), specifically https://github.com/hiyouga/LLaMA-Factory/blob/main/examples/train_full/llama3_full_sft_ds3.yaml.
+Our training script was adapted from the example provided in LlamaFactory (Zheng et al., 2024b)  [^1] https://github.com/hiyouga/LLaMA-Factory/blob/main/examples/train_full/llama3_full_sft_ds3.yaml.
 The training was configured with a batch size of 1 per device, gradient accumulation steps of 4, a learning rate of 1e-5, and he model was trained for 3 epochs, with warmup over 20 steps and a cosine learning rate scheduler.
@@
 This metric helps assess the reliability of evaluators, such as GPT-4-o and GPT-4-o-mini, across different datasets.
-A.6 Dataset
+A.6 Models Used for Ranking
+We evaluate model ranking using 30 widely used models from the AlpacaEval dataset, covering diverse architectures and capabilities.
+Our model set $\mathcal{M}$ includes families such as Llama (Touvron et al., 2023), Vicuna (Zheng et al., 2023), GPT (Achiam et al., 2023), Claude (Anthropic, 2024), Qwen (Yang et al., 2024a), Mistral (Jiang et al., 2023), Yi (Young et al., 2024), and WizardLM (Xu et al., 2023), ensuring a broad assessment.
+Proprietary models include OpenAI’s GPT series (e.g., gpt-3.5-turbo-0301, gpt-4o-2024-05-13) and Anthropic’s Claude models (e.g., claude-2, claude-3-opus-20240229), serving as strong baselines.
+Open-source models include multiple versions of Llama (e.g., Meta-Llama-3-8B-Instruct, llama-2-70b-chat-hf) and Qwen (e.g., Qwen1.5-72B-Chat, Qwen2-72B).
+We also incorporate Mistral models (e.g., Mistral-7B-Instruct-v0.2, Mixtral-8x22B-Instruct-v0.1), reflecting advances in mixture-of-experts architectures.
+Additional selections include gemini-pro, tulu-2-dpo-70b, oasst-sft-llama-33b, dbrx-instruct, wizardlm-13b, and Yi-34B-Chat, contributing to ranking diversity.
+This diverse model set enables a comprehensive evaluation of ranking accuracy using AlpacaEval’s ground-truth rankings.
+A.7 Dataset
 In this appendix, we provide detailed information about the datasets used in main text.
@@
 We exclusively used the Koala test set to gauge our model’s proficiency in handling authentic conversational queries.
-A.7 Aggregation Process in GED Across Different Tasks
+A.8 Aggregation Process in GED Across Different Tasks
 The GED implementation differs between response selection and model ranking tasks due to their objectives.
@@
 Multiple evaluator GED 57.59 61.14 61.14 67.17 66.46 62.70
-Appendix C Impact of Evaluator Quantity on Denoising Quality
+Appendix C Comparison with Point-Wise Scoring Methods
+To further investigate the difference between point-wise and preference-based scoring methods, we conducted additional experiments under the response selection setting.
+Specifically, we evaluated a point-wise baseline where evaluators assigned scores to responses on a scale of 1 to 5, selecting the highest-rated response.
+The results are shown in Table 6.
+Table 6: Comparison of point-wise methods and GED on response selection tasks.
+Model HumanEval AlpacaEval MATH GSM8k GAIA Avg
+Point-wise - Llama3-8B 60.97 28.77 74.06 82.65 10.97 51.48
+Point-wise - Mistral-7B 62.83 26.93 74.02 83.21 10.22 51.44
+Point-wise - Qwen2-7B 60.41 28.30 74.32 84.27 10.82 51.62
+Point-wise - Majority Voting 63.72 29.42 74.71 84.36 11.32 52.70
+GED (Llama3-8B, Mistral-7B, Qwen2-7B) 70.73 32.44 75.58 88.18 13.33 56.05
+From the results, we observe that point-wise methods consistently underperform compared to GED across all tasks.
+This can be attributed to the fact that point-wise scoring assesses each response independently, without considering the relative quality of different responses.
+As a result, it lacks the global information available in preference-based ranking, where responses are directly compared.
+While majority voting improves point-wise performance slightly by aggregating scores from multiple evaluators, it remains inferior to GED, which leverages pairwise preferences to construct a more informed global ranking.
+Despite its limitations, the point-wise approach has an advantage in terms of computational efficiency, as it requires fewer comparisons than pairwise methods.
+This trade-off suggests that the choice between point-wise and preference-based scoring should depend on the specific application: point-wise scoring may be preferable in large-scale settings where efficiency is critical, while GED is more effective for achieving higher-quality rankings.
+Appendix D Impact of Evaluator Quantity on Denoising Quality
 We investigated how the number of evaluators affects the denoising quality in GED by conducting experiments using different numbers of evaluators in the response selection setting.
@@
 Incorporating more evaluators allows the denoising process to better identify and mitigate inconsistencies in the preference graphs, leading to improved overall performance in response selection tasks.
-Appendix D Construction of the Preference Graph
-In this section, we provide a detailed explanation of the process used to construct the preference graph sets for both the model ranking and response selection tasks, as outlined in Algorithms 1 and 2.
-These algorithms form the backbone of our method, enabling the representation of pairwise preferences as directed graphs, which are essential for downstream aggregation and ranking.
-Algorithm 1 describes the construction process for generating a set of preference graphs for the model ranking task.
-The procedure is as follows: Initialization: For each question $q\in Q$, we begin by initializing a vertex set $V_{q}$, where each vertex $v_{i}$ corresponds to a model $\mathcal{M}_{i}$ in the set of models $M$.
-We also initialize an empty set of edges $A_{a}$ and a weight function $w_{a}$, which will be used to track the strength of the preferences between model pairs.
-Pairwise Comparisons: For each pair of models $\mathcal{M}_{i}$ and $\mathcal{M}_{j}$, the assigned evaluator $a\in\mathcal{A}$ assesses their responses to the given question $q$.
-If evaluator $a$ prefers $\mathcal{M}_{i}$ over $\mathcal{M}_{j}$, a directed edge $(v_{i}\rightarrow v_{j})$ is added to the edge set $A_{a}$, and its corresponding weight is incremented.
-Conversely, if $\mathcal{M}_{j}$ is preferred, the edge $(v_{j}\rightarrow v_{i})$ is added or its weight is updated.
-Graph Storage: Once all pairwise comparisons have been processed for a given evaluator, the resulting graph $G_{a}=(V_{q},A_{a},w_{a})$ is stored for that evaluator.
-This process is repeated for all evaluators in $\mathcal{A}$ and for all questions in $Q$, generating a set of preference graphs for each evaluator and question.
 Algorithm 1 Construction of the Preference Graph for Model Ranking
@@
 27:end for
+Algorithm 2 Construction of the Preference Graph for Response Selection
+Set of candidate answers $\{ans_{1},ans_{2},\ldots,ans_{n}\}$ for each question $q\in Q$, set of evaluators $\mathcal{A}=\{a_{1},a_{2},\ldots,a_{k}\}$
+Set of preference graph sets $\{G_{a}:a\in\mathcal{A}\}$ for each question $q\in Q$
+for each question $q\in Q$ do
+     for each evaluator $a\in\mathcal{A}$ do
+         Initialize vertex set $V_{q}=\{v_{1},v_{2},\ldots,v_{n}\}$, where each $v_{i}$ corresponds to $ans_{i}$
+         Initialize edge set $A_{a}=\emptyset$, and weight function $w_{a}:A_{a}\rightarrow\mathbb{R}^{+}$
+         for each pair of answers $(ans_{i},ans_{j})$ with $i\neq j$ do
+              if $a(ans_{i},ans_{j})>0$ then
+                  if $(v_{i},v_{j})\notin A_{a}$ then
+                       Add directed edge $(v_{i}\rightarrow v_{j})$ to $A_{a}$, set $w_{a}(v_{i},v_{j})=1$
+                  else
+                       Increment $w_{a}(v_{i},v_{j})$ by 1
+                  end if
+              else
+                  if $(v_{j},v_{i})\notin A_{a}$ then
+                       Add directed edge $(v_{j}\rightarrow v_{i})$ to $A_{a}$, set $w_{a}(v_{j},v_{i})=1$
+                  else
+                       Increment $w_{a}(v_{j},v_{i})$ by 1
+                  end if
+              end if
+         end for
+         Store the preference graph $G_{a}=(V_{q},A_{a},w_{a})$ for evaluator $a$
+     end for
+end for
+Appendix E Construction of the Preference Graph
+In this section, we provide a detailed explanation of the process used to construct the preference graph sets for both the model ranking and response selection tasks, as outlined in Algorithms 1 and 2.
+These algorithms form the backbone of our method, enabling the representation of pairwise preferences as directed graphs, which are essential for downstream aggregation and ranking.
+Algorithm 1 describes the construction process for generating a set of preference graphs for the model ranking task.
+The procedure is as follows: Initialization: For each question $q\in Q$, we begin by initializing a vertex set $V_{q}$, where each vertex $v_{i}$ corresponds to a model $\mathcal{M}_{i}$ in the set of models $M$.
+We also initialize an empty set of edges $A_{a}$ and a weight function $w_{a}$, which will be used to track the strength of the preferences between model pairs.
+Pairwise Comparisons: For each pair of models $\mathcal{M}_{i}$ and $\mathcal{M}_{j}$, the assigned evaluator $a\in\mathcal{A}$ assesses their responses to the given question $q$.
+If evaluator $a$ prefers $\mathcal{M}_{i}$ over $\mathcal{M}_{j}$, a directed edge $(v_{i}\rightarrow v_{j})$ is added to the edge set $A_{a}$, and its corresponding weight is incremented.
+Conversely, if $\mathcal{M}_{j}$ is preferred, the edge $(v_{j}\rightarrow v_{i})$ is added or its weight is updated.
+Graph Storage: Once all pairwise comparisons have been processed for a given evaluator, the resulting graph $G_{a}=(V_{q},A_{a},w_{a})$ is stored for that evaluator.
+This process is repeated for all evaluators in $\mathcal{A}$ and for all questions in $Q$, generating a set of preference graphs for each evaluator and question.
 Algorithm 2 follows a similar structure but applies to the response selection task, where the objective is to rank a set of candidate answers for each question: Initialization: For each question $q\in Q$, we initialize a vertex set $V_{q}$, where each vertex corresponds to a candidate answer $ans_{i}$.
@@
 These graphs encapsulate the evaluators’ preferences and provide a structured representation of pairwise comparisons, facilitating further analysis.
-Algorithm 2 Construction of the Preference Graph for Response Selection
-1:Set of candidate answers $\{ans_{1},ans_{2},\ldots,ans_{n}\}$ for each question $q\in Q$, set of evaluators $\mathcal{A}=\{a_{1},a_{2},\ldots,a_{k}\}$
-2:Set of preference graph sets $\{G_{a}:a\in\mathcal{A}\}$ for each question $q\in Q$
-3:for each question $q\in Q$ do
-4:     for each evaluator $a\in\mathcal{A}$ do
-5:         Initialize vertex set $V_{q}=\{v_{1},v_{2},\ldots,v_{n}\}$, where each $v_{i}$ corresponds to $ans_{i}$
-6:         Initialize edge set $A_{a}=\emptyset$, and weight function $w_{a}:A_{a}\rightarrow\mathbb{R}^{+}$
-7:         for each pair of answers $(ans_{i},ans_{j})$ with $i\neq j$ do
-8:              if $a(ans_{i},ans_{j})>0$ then
-9:                  if $(v_{i},v_{j})\notin A_{a}$ then
-10:                       Add directed edge $(v_{i}\rightarrow v_{j})$ to $A_{a}$, set $w_{a}(v_{i},v_{j})=1$
-11:                  else
-12:                       Increment $w_{a}(v_{i},v_{j})$ by 1
-13:                  end if
-14:              else
-15:                  if $(v_{j},v_{i})\notin A_{a}$ then
-16:                       Add directed edge $(v_{j}\rightarrow v_{i})$ to $A_{a}$, set $w_{a}(v_{j},v_{i})=1$
-17:                  else
-18:                       Increment $w_{a}(v_{j},v_{i})$ by 1
-19:                  end if
-20:              end if
-21:         end for
-22:         Store the preference graph $G_{a}=(V_{q},A_{a},w_{a})$ for evaluator $a$
-23:     end for
-24:end for
-Appendix E Proof of Theorem 1
+Appendix F The Importance of Addressing Cyclic Inconsistency
+Cyclic inconsistencies in preference data introduce contradictions that undermine meaningful ranking and evaluation.
+Many applications, including alignment optimization (Ouyang et al., 2022; Song et al., 2024), recommendation systems (Hou et al., 2024), and model evaluation (Liu et al., 2024), rely on transitive preferences for stability and interpretability.
+For example, OpenAI’s RLHF pipelines (Ouyang et al., 2022) construct training data based on transitive preferences (e.g., A $\succ$ B $\succ$ C), which ensures consistency in model learning.
+However, if preference cycles exist (e.g., A $\succ$ B $\succ$ C $\succ$ A), the resulting contradictions lead to instability and degrade training performance.
+In our experiments, removing cycles and transforming preference graphs into directed acyclic graphs (DAGs) improved results across multiple tasks, including response selection, model ranking, and instruction tuning.
+These improvements were consistent across widely used benchmarks, suggesting that most observed cycles are artifacts of noise rather than meaningful violations of transitivity.
+Ensuring acyclicity enhances dataset reliability and improves model effectiveness in ranking and evaluation tasks.
+Appendix G Proof of Theorem 1
 See 1
-Proof E.1
+Proof.
 For brevity, we consider all edges in $G$ have weights equal to 1 and all weights in $\widehat{G}$ are divided by $N$.
@@
 $\mathbb{P}\left(G\subseteq\mathrm{MAS}(\widehat{G})\right)\geq$ $\mathbb{P}\left(\mathcal{E}_{1}\cap\mathcal{E}_{2}\right)\geq 1-2|A|\exp\left(-\frac{N\epsilon^{2}}{2}\right)-2U\exp\left(-\frac{N\epsilon^{2}}{6U^{2}\delta_{2}+2U\epsilon}\right).$
-Lemma 1
+Lemma 1.
 For a weighted directed graph $G=(V,A,w)$, if $(u\rightarrow v),(v\rightarrow u)\in A$, then $\mathrm{MAS}(G)$ contains exactly one of $(u\rightarrow v)$ and $(v\rightarrow u)$.
-Proof E.2
+Proof.
 Recall that $\mathrm{MAS}(G)$ gives an acyclic subgraph of $G$ with the maximum weight.
@@
 Therefore, $\mathrm{MAS}(G)$ must contain exactly one of $(u\rightarrow v)$ and $(v\rightarrow u)$.
-Appendix F Impact of Evaluator Weighting on GED Performance
+∎
+Appendix H Impact of Evaluator Weighting on GED Performance
 In our theoretical analysis, GED assumes equal weighting of edges in the preference graphs.
@@
 Systematically designing and optimizing these weighting schemes could lead to more robust and accurate evaluation frameworks.
-Appendix G Evaluating GED on Diverse Metrics
-To address the need for more comprehensive assessments beyond accuracy-based metrics, we expanded our evaluation of GED to include nuanced quality aspects of LLM-generated outputs, such as factuality, relevance, coherence, informativeness, helpfulness, and validity.
-This evaluation was motivated by the understanding that tasks involving LLMs often require subtle judgments beyond simple accuracy, and GED’s adaptability to these scenarios is crucial.
-Building upon this foundation, we conducted experiments following (Chen et al., 2023a).
-Specifically, we used Llama3-70B to generate ten candidate responses for each query, and GED was applied in the response selection setting with Llama3-8B, Mistral-7B, and Qwen2-7B serving as evaluators.
-Each metric—factuality, relevance, coherence, informativeness, helpfulness, and validity—was assessed to measure GED’s ability to enhance the overall quality of LLM outputs.
-The results, shown in Table 7, highlight GED’s effectiveness.
-GED consistently outperformed individual evaluators and random selection across all metrics.
-For instance, GED improved factuality by approximately 5 percentage points over the best individual evaluator.
-Similarly, it enhanced relevance and coherence, indicating better alignment with the query and logically consistent responses.
-GED also demonstrated notable improvements in informativeness and helpfulness, suggesting that the responses selected by GED provide more valuable and user-centric information.
-These findings confirm GED’s adaptability to tasks requiring nuanced quality assessments.
-By aggregating preferences from multiple evaluators, GED captures subtle qualities of generated content, enhancing not only accuracy but also the overall reliability and utility of LLM outputs.
-This adaptability reinforces GED’s value in scenarios demanding comprehensive evaluations of language models.
-Table 7: Performance comparison on nuanced quality metrics (in %).
-GED outperforms individual evaluators and random selection across Factuality, Relevance, Coherence, Inform., Helpful.
-and Validity metrics.
-Method Factuality Relevance Coherence Inform. Helpful. Validity Avg.
-Random 86.73 87.91 92.47 77.62 17.48 48.92 68.52
-Llama3-8B 88.59 89.91 94.41 79.77 18.48 50.92 70.35
-Mistral-7B 89.10 90.29 94.85 79.95 18.55 51.13 70.65
-Qwen2-7B 89.25 90.44 95.03 80.09 18.58 51.21 70.77
-GED 94.73 95.91 97.36 86.62 19.48 55.92 75.00
-Appendix H Expanded Related Work on Preference Denoising for LLMs
+Appendix I Expanded Related Work on Preference Denoising for LLMs
 This work situates itself within the broader field of denoising preference evaluations for LLMs, addressing inconsistencies and noise in preference graphs.
@@
 Additionally, GED distinguishes itself by providing theoretical guarantees for recovering the ground truth preference DAG under certain conditions.
-Furthermore, the ensemble mechanism in GED demonstrates the novel insight that combining weaker evaluators can surpass the performance of a single stronger evaluator, a feature not emphasized in the aforementioned methods.
-Appendix I Impact of Evaluator Selection on GED Performance
+Furthermore, the ensemble mechanism in GED demonstrates the novel insight that combining small evaluators can surpass the performance of a single stronger evaluator, a feature not emphasized in the aforementioned methods.
+Appendix J Impact of Evaluator Selection on GED Performance
 The selection of weak evaluators is a critical factor influencing the performance of GED.
-While our initial experiments demonstrated the benefits of combining weaker evaluators, further analysis is necessary to understand how their diversity and capabilities affect GED’s effectiveness.
+While our initial experiments demonstrated the benefits of combining weak evaluators, further analysis is necessary to understand how their diversity and capabilities affect GED’s effectiveness.
 To address this, we conducted additional experiments evaluating three different configurations of evaluator sets: the original set comprising Llama3-8B, Mistral-7B, and Qwen2-7B; a diverse set that adds Gemma-9B [^5] https://huggingface.co/google/gemma-2-9b-it to introduce more variation in model architecture and training data; and a higher-capability set that replaces smaller models with larger ones, specifically Llama3-70B, Mistral-8×7B, and Qwen2-72B.
@@
 Higher Capability 74.73 47.92 75.58 90.04 16.21 60.89
-Appendix J Cost Considerations and Active Learning with ActiveGED
-Using multiple evaluators and aggregating their preferences can be computationally expensive, especially when constructing dense preference graphs.
-This limitation becomes more pronounced in scenarios where preferences across all pairs of evaluators are required, which scales quadratically with the number of responses or models being compared.
-To address this, we clarify the computational trade-offs and propose an active learning-based approach to reduce the number of required pairwise evaluations.
-To further reduce the number of pairwise evaluations needed, we developed an active learning algorithm called ActiveGED.
-This algorithm strategically selects the most informative pairs for evaluation, effectively lowering the overall computational cost while maintaining high performance.
-ActiveGED combines random sampling with uncertainty-based selection to maximize information gain from each additional pairwise evaluation.
-We evaluated ActiveGED under budget constraints of 30% and 50% of the total possible pairwise comparisons, where we set the random budget ratio as 0.5.
-The results, presented in Table 9, demonstrate that ActiveGED achieves competitive performance with significantly fewer evaluations.
-For example, under a 50% budget, ActiveGED retains most of the performance benefits of full GED while cutting the number of pairwise comparisons in half.
-Table 9: Performance comparison of ActiveGED under different budget constraints. ActiveGED achieves competitive performance with significantly fewer pairwise evaluations.
-Evaluators Set of GED HumanEval AlpacaEval MATH GSM8k GAIA Avg.
-Random 57.93 29.58 72.75 84.67 11.52 51.29
-ActiveGED (30%) 67.28 30.96 74.65 85.73 11.39 54.00
-ActiveGED (50%) 68.62 31.91 74.87 87.06 12.08 54.91
-GED (Full Budget) 70.73 32.44 75.58 88.18 13.33 56.05
-The algorithm behind ActiveGED is outlined in Algorithm 3.
-It begins by randomly selecting a portion of the budget to initialize the preference graph and then iteratively selects the most informative edges based on uncertainty, as estimated using PageRank scores.
-This approach balances exploration (random sampling) and exploitation (uncertainty-based selection) to efficiently construct an accurate preference graph.
 Algorithm 3 ActiveGED
@@
 26:return $G^{*}=(V,A^{*},w^{*})$
+Appendix K Cost Considerations and Active Learning with ActiveGED
+Using multiple evaluators and aggregating their preferences can be computationally expensive, especially when constructing dense preference graphs.
+This limitation becomes more pronounced in scenarios where preferences across all pairs of evaluators are required, which scales quadratically with the number of responses or models being compared.
+To address this, we clarify the computational trade-offs and propose an active learning-based approach to reduce the number of required pairwise evaluations.
+To further reduce the number of pairwise evaluations needed, we developed an active learning algorithm called ActiveGED.
+This algorithm strategically selects the most informative pairs for evaluation, effectively lowering the overall computational cost while maintaining high performance.
+ActiveGED combines random sampling with uncertainty-based selection to maximize information gain from each additional pairwise evaluation.
+We evaluated ActiveGED under budget constraints of 30% and 50% of the total possible pairwise comparisons, where we set the random budget ratio as 0.5.
+The results, presented in Table 10, demonstrate that ActiveGED achieves competitive performance with significantly fewer evaluations.
+For example, under a 50% budget, ActiveGED retains most of the performance benefits of full GED while cutting the number of pairwise comparisons in half.
+Table 10: Performance comparison of ActiveGED under different budget constraints. ActiveGED achieves competitive performance with significantly fewer pairwise evaluations.
+Evaluators Set of GED HumanEval AlpacaEval MATH GSM8k GAIA Avg.
+Random 57.93 29.58 72.75 84.67 11.52 51.29
+ActiveGED (30%) 67.28 30.96 74.65 85.73 11.39 54.00
+ActiveGED (50%) 68.62 31.91 74.87 87.06 12.08 54.91
+GED (Full Budget) 70.73 32.44 75.58 88.18 13.33 56.05
+The algorithm behind ActiveGED is outlined in Algorithm 3.
+It begins by randomly selecting a portion of the budget to initialize the preference graph and then iteratively selects the most informative edges based on uncertainty, as estimated using PageRank scores.
+This approach balances exploration (random sampling) and exploitation (uncertainty-based selection) to efficiently construct an accurate preference graph.
 ActiveGED demonstrates that by carefully selecting the most informative pairs, it is possible to achieve competitive performance with significantly fewer evaluations.
 This makes GED more scalable and practical for real-world applications where computational resources are limited.
-Appendix K Mitigating Evaluator Biases in GED
+Appendix L Mitigating Evaluator Biases in GED
 Evaluator biases, such as position bias and token bias, can significantly impact the accuracy and fairness of the preference graphs used in GED.
@@
 By averaging the results across both orderings, any systematic preference for a particular label is effectively neutralized.
-Appendix L Denoising of Preference Graphs for GED
+Appendix M Denoising of Preference Graphs for GED
 In this section, we describe the denoising procedure used in the GED framework, specifically for transforming the aggregated preference graph into a DAG.
@@
 19:$s\leftarrow s_{1}\cdot s_{2}$
-Appendix M Rank ensemble method
+Appendix N Rank ensemble method
 Weight score (Adler et al., 2002)
@@
 In both methods, the objective is to maximize the (weighted) pairwise agreement between the consensus ranking and the input rankings.
-Appendix N Visualization
+Appendix O Visualization
 (a) Raw Preference Graph 1
@@
 GED effectively minimizes edge removals while maintaining the integrity of the original preference graph, making it a highly efficient solution for improving the consistency of preference evaluations.
-Appendix O Prompt Template
+Appendix P Prompt Template
 Prompt for Response Selection.

### v4

@@
 Despite the remarkable success of Large Language Models (LLMs), evaluating their outputs’ quality regarding preference remains a critical challenge.
-Existing works usually leverage an LLM as the judge for comparing LLMs’ output pairwisely, yet such model-based evaluator is weak evaluator due to conflicting preference, i.e., output A is better than B, B than C, but C than A, causing contradictory evaluation results.
-To address this, we introduce GED (Preference Graph Ensemble and Denoise), a novel approach that leverages multiple model-based evaluators to construct preference graphs, and then ensemble and denoise these graphs for better, non-contradictory evaluation results.
-In particular, our method consists of two primary stages: aggregating evaluations into a unified graph and applying a denoising process to eliminate cyclic inconsistencies, ensuring a directed acyclic graph (DAG) structure.
+While existing works usually leverage a strong LLM as the judge for comparing LLMs’ response pairwisely, such a single-evaluator approach is vulnerable to cyclic preference , i.e., output A is better than B, B than C, but C is better than A, causing contradictory evaluation results.
+To address this, we introduce PGED (Preference Graph Ensemble and Denoising , a novel approach that leverages multiple model-based evaluators to construct preference graphs, and then ensembles and denoises these graphs for acyclic, non-contradictory evaluation results.
 We provide theoretical guarantees for our framework, demonstrating its efficacy in recovering the ground truth preference structure.
-Extensive experiments on ten benchmarks demonstrate GED’s superiority in three applications: model ranking, response selection, and model alignment tasks.
-Notably, GED combines small LLM evaluators (e.g., Llama3-8B, Mistral-7B, Qwen2-7B) to outperform strong ones (e.g., Qwen2-72B), showcasing its effectiveness in enhancing evaluation reliability and improving model performance.
-Machine Learning, ICML
+Extensive experiments on ten benchmarks demonstrate PGED’s superiority in three applications: 1) model ranking for evaluation, 2) response selection for test-time scaling, and 3) data selection for model fine-tuning.
+Notably, PGED combines small LLM evaluators (e.g., Llama3-8B, Mistral-7B, Qwen2-7B) to outperform strong ones (e.g., Qwen2-72B), showcasing its effectiveness in enhancing evaluation reliability and improving model performance.
 1 Introduction
-Large Language Models (LLMs) have rapidly transformed various fields within artificial intelligence, particularly natural language processing (NLP) and decision-making systems (Wu et al., 2023; Li et al., 2023a).
-Despite the remarkable success of LLMs, the need for effective evaluation methods becomes paramount (Liu et al., 2023; Desmond et al., 2024; Siska et al., 2024; Boyeau et al., 2024; Chatzi et al., 2024).
-Preference evaluation, as one of the most important assessment methods, plays an indispensable role in evaluating and optimizing model performance (Rafailov et al., 2024; Yuan et al., 2024; Dubois et al., 2024b).
-Existing works usually leverage a powerful LLM (e.g., GPT4 (Achiam et al., 2023)) as the judge for comparing LLMs’ output pairwisely (Li et al., 2023b; Chen et al., 2023b; Wang et al., 2022).
+Large Language Models (LLMs) have rapidly advanced various areas of artificial intelligence, particularly natural language processing and decision-making (Wu et al.
+2023; Li et al.
+2023a).
+As LLMs become increasingly capable, effective evaluation becomes critical (Liu et al.
+2023; Desmond et al.
+2024; Siska et al.
+2024; Boyeau et al.
+2024; Chatzi et al.
+2024).
+Among evaluation methods, preference evaluation plays a critical role in both evaluating and aligning LLMs (Rafailov et al.
+2024; Yuan et al.
+2024; Dubois et al.
+2024b).
+A common practice is to rely on a strong LLM (e.g., GPT-4 (Achiam et al.
+2023)) as the judge to conduct pairwise comparisons (Li et al.
+2023b; Chen et al.
+2023b; Wang et al.
+2022).
 Figure 1: (a) A preference graph exhibiting cyclic inconsistencies (e.g., A $\succ$ B $\succ$ C $\succ$ A), which violate transitivity.
 (b) Empirical results showing that even advanced LLMs (e.g., GPT-4-o) exhibit significant noise in preference judgments, leading to inconsistent evaluations.
-(c) Overview of our proposed framework, GED, which ensembles multiple weak evaluators and applies denoising to recover a directed acyclic graph.
-However, while such model-based pairwise preference evaluations offer a flexible approach, they can lead to contradictory evaluations in the assessment process (Naresh et al., 2024; Zhang et al., 2024b).
-For example, an LLM might evaluate three responses and conclude that Response A is better than Response B, Response B is better than Response C, yet paradoxically also rank Response C as better than Response A.
-These cyclic patterns violate the transitivity assumption of preferences established in prior work (Ouyang et al., 2022; Song et al., 2024; Hou et al., 2024; Liu et al., 2024), thereby undermining the reliability of evaluation results.
-We model this conflicting preference via the preference graph.
-Specifically, a preference graph is constructed with each response as a node and directed edges indicating pairwise preferences—an edge from node A to node B shows that the evaluator preferred response A over B.
-The noise illustrated by cycles (A $\succ$ B $\succ$ C $\succ$ A) manifests as loops in the preference graph.
-This process is illustrated in Figure 1 (a).
-Ideally, a preference graph should be structured as a directed acyclic graph (DAG) to maintain consistency.
-In this paper, we define an evaluator as weak if it produces cycles into the preference graph.
-As shown in Figure  1 (b), we evaluated 10 Llama3-70B (AI@Meta, 2024) responses on HumanEval (Chen et al., 2021) and MATH (Hendrycks et al., 2021) using GPT-4-o, GPT-4-o-mini, GPT-3.5 (Achiam et al., 2023), Qwen2-72B (Yang et al., 2024a), and Llama3-8B (AI@Meta, 2024) as evaluators.
-Even with GPT-4-o, 64% of preference graphs in HumanEval and 38% in MATH contained cycles, highlighting persistent noise and the limitations of these models as weak evaluators.
-To address this, we propose a novel framework, GED (Preference Graph Ensemble and Denoise), to address the inconsistencies in preference graphs generated through pairwise evaluations.
-Our method involves two key steps: (1) ensembling multiple weak evaluators to mitigate noise introduced by individual evaluators and (2) applying a denoising process to the resulting preference graph.
-By aggregating evaluations from multiple weak evaluators, we ”average out” the noise and biases, resulting in a more robust approximation of the true preference structure.
+(c) Overview of our proposed framework, PGED, which ensembles multiple preference evaluators and applies denoising to recover a directed acyclic graph.
+However, such model-based evaluator setups often lead to cyclic preferences, where inconsistent rankings emerge, for instance, preferring A over B, B over C, yet C over A (Naresh, Tulabandhula et al.
+2024; Zhang, Yin, and Wan 2024).
+These cyclic patterns violate the transitivity assumption of preferences established in prior work (Ouyang et al.
+2022; Song et al.
+2024; Hou et al.
+2024; Liu et al.
+2024), thereby undermining the reliability of evaluation results.
+We model this conflicting preference using a preference graph, where nodes represent responses and directed edges indicate pairwise preferences.
+Cycles in such graphs (e.g., A $\succ$ B $\succ$ C $\succ$ A) reflect evaluation noise, as shown in Figure 1(a).
+Ideally, a preference graph should be a directed acyclic graph (DAG) to maintain consistency.
+Empirically, we evaluated 10 Llama3-70B (AI@Meta 2024) responses on HumanEval (Chen et al.
+2021) and MATH (Hendrycks et al.
+2021), using GPT-4-o, GPT-4-o-mini, GPT-3.5 (Achiam et al.
+2023), Qwen2-72B (Yang et al.
+2024), and Llama3-8B as judges.
+Even with GPT-4-o, 64% of preference graphs in HumanEval and 38% in MATH contained cycles, (Figure 1(b)), demonstrating the persistent noise in LLM-based preference evaluations and motivating the need for a more robust approach.
+To address this, we propose a novel framework, PGED (Preference Graph Ensemble and Denoising).
+Our method involves two key steps: (1) ensembling multiple preference evaluators to mitigate noise introduced by individual evaluators and (2) applying a denoising process to the resulting preference graph.
+By aggregating evaluations from multiple individual evaluators, we "average out" the noise and biases, resulting in a more robust approximation of the true preference structure.
 The denoising step further refines this aggregated graph by removing inconsistencies, ensuring the final preference graph is more reliable for downstream tasks.
-The overall process of GED is illustrated in Figure 1 (c).
-We provide a theoretical analysis demonstrating the soundness of GED, showing that by treating each individual preference graph as a random perturbation of a ground truth DAG, our ensemble and denoising framework can recover the ground truth DAG with high probability.
-To validate the practical efficacy of GED, we conduct extensive experiments across model ranking, response selection, and model alignment tasks, utilizing ten widely recognized benchmark datasets, including HumanEval (Chen et al., 2021), AlpacaEval (Li et al., 2023b), MATH (Hendrycks et al., 2021), GSM8k (Chen et al., 2021), GAIA (Mialon et al., 2023), LIMA (Zhou et al., 2023), Vicuna (Chiang et al., 2023), Koala (Vu et al., 2023), WizardLM (Xu et al., 2023), and Self-Instruct (Wang et al., 2022).
-In these experiments, GED consistently outperformed baseline methods.
-For instance, in the response selection task, GED achieved an average performance gain of 4.51% over baseline methods across multiple benchmarks.
-Additionally, GED demonstrated substantial gains in scenarios where combining preference graphs from small evaluators surpassed the performance of even stronger individual evaluators.
-Specifically, when using using Llama3-8B, Mistral-7B, and Qwen2-7B as evaluators, GED exceeded the performance of using the Qwen2-72B in response selection task.
-These results highlight GED’s ability to mitigate preference noise, improve consistency, and enhance model performance across diverse evaluation settings.
+The overall process of PGED is illustrated in Figure 1 (c).
+We provide a theoretical analysis demonstrating the soundness of PGED, showing that by treating each individual preference graph as a random perturbation of a ground truth DAG, our ensemble and denoising framework can recover the ground truth DAG with high probability.
+To validate the practical efficacy of PGED, we conduct ten tasks ranging from response selection, model ranking to model alignment tasks, utilizing ten widely recognized benchmark datasets.
+In these experiments, PGED consistently outperformed baseline methods.
+Additionally, PGED demonstrated substantial gains in scenarios where combining preference graphs from small evaluators surpassed the performance of even stronger individual evaluators.
+Specifically, when using using Llama3-8B, Mistral-7B, and Qwen2-7B as evaluators, PGED exceeded the performance of using the Qwen2-72B in response selection task.
+These results highlight PGED’s ability to mitigate preference noise, improve consistency, and enhance model performance across diverse evaluation settings.
+Our contributions are summarized as follows:
+  • We propose PGED, a novel framework that ensembles multiple preference evaluators and denoises the resulting preference graphs to produce acyclic and reliable evaluation outputs.
+  • We provide theoretical guarantees that PGED can recover the ground-truth preference DAG under a reasonable noise model, thereby offering provable consistency.
+  • We conduct extensive experiments on ten benchmark datasets and three key tasks, model ranking, response selection, and data selection for fine-tuning, demonstrating that PGED significantly improves evaluation robustness.
+Notably, PGED using small models outperforms even strong single evaluators like Qwen2-72B.
 2 Related Work
-Preference evaluation of LLMs.
-Reference-free evaluation metrics have a long history (Louis & Nenkova, 2013; Boyeau et al., 2024; Chatzi et al., 2024; Shankar et al., 2024; Naresh et al., 2024), which evaluates the generated text based on intrinsic properties and coherence with the context.
-Although they achieve high accuracy on matching inner-evaluator, the achievement suffers from spurious correlations such as perplexity and length (Durmus et al., 2022).
-Recently, people have started using a strong model (e.g., GPT-4) as an evaluator to perform a zero-shot reference-free evaluation on the weak models (Shen et al., 2023; Dubois et al., 2024b; Chen et al., 2023b).
-However, using LLM-based preference evaluations can introduce inconsistencies in preference graphs, often resulting in cyclic preferences or contradictions when comparing multiple outputs.
+Preference evaluation.
+Classical preference aggregation includes Elo, which iteratively updates item ratings from pairwise outcomes (Jiang, Ren, and Lin 2023); Bradley–Terry (BT/BTL) models, which assign latent utilities to explain win probabilities (Bradley and Terry 1952; Lu and Boutilier 2011); and MergeSort-style procedures, which sort items via pairwise comparisons with $O(n\log n)$ queries (Shyamasundar 1996).
+LLM-based preference evaluation increasingly relies on a strong model (e.g., GPT-4) as a zero-shot, reference-free judge of weaker models’ outputs (Shen et al.
+2023; Dubois et al.
+2024b; Chen et al.
+2023b).
+PRD combines peer ranking over pairwise preferences with peer discussion between LLMs to reach consensus (Li, Patel, and Du 2023a).
+ChatEval forms a multi-agent referee team that debates and evaluates responses for more reliable assessments (Chan et al.
+2023a).
+However, LLM judges can induce contradictory preferences and cycles when aggregating across outputs, yielding noisy and inconsistent preference graphs (Naresh, Tulabandhula et al.
+2024; Zhang, Yin, and Wan 2024).
+SPA (Kadekodi, McTavish, and Ustun 2025) produces partial-order rankings that abstain on disputed pairs to balance comparability and disagreement, while ContraSolver (Zhang, Yin, and Wan 2024) builds a preference graph with self-annotations and flags edges likely responsible for contradictions.
+In contrast, our approach ensembles multiple evaluator-induced graphs and denoises them via a weighted feedback-arc-set objective to recover an acyclic structure suitable for robust ranking.
 Weak supervision.
-The concept of weak-to-strong supervision originates from the need to leverage noisy or partial labels in machine learning tasks, enabling the development of more robust models from imperfect data (Ratner et al., 2016; Zhang et al., 2023b, 2022).
-In LLMs, weak-to-strong supervision aids AI alignment by allowing weaker models to improve strong ones, enhancing performance without extensive data and supporting scalable oversight (Zheng et al., 2024a; Guo & Yang, 2024; Tong et al., 2024).
-Similarly, in task-oriented LLMs, weak-to-strong learning improves LLM’s ability by enabling strong models to refine their data autonomously, boosting performance without extensive high-quality input (Zhang et al., 2023a; Yang et al., 2024b).
-Through weak-to-strong supervision, LLM performance can be significantly improved by iteratively transforming low-quality labels into more reliable ones, leading to more effective model training and robust outputs (Zakershahrak & Ghodratnama, 2024; Lang et al., 2024).
-3 Methodologies
-In this section, we begin by defining a preference graph, which serves as the foundation for representing pairwise preferences among candidates (Section 3.1).
-Building on this foundation, we introduce GED structured into three key stages (Section 3.2): (1) graph ensemble, where we aggregate individual preference graphs into a unified structure, (2) graph denoising, which removes cycles and inconsistencies to ensure the preference graph is acyclic, and (3) graph-to-ranking, where we extract a reliable ranking of candidates from the denoised graph.
-Below, we provide detailed descriptions of each step.
-3.1 Preference Graph
+The concept of weak supervision originates from the need to leverage noisy or partial labels in machine learning tasks, enabling the development of more robust models from imperfect data (Ratner et al.
+2016; Zhang, Song, and Ratner 2023; Zhang et al.
+2022).
+In LLMs, weak-to-strong supervision aids AI alignment by allowing weaker models to improve strong ones, enhancing performance without extensive data and supporting scalable oversight (Zheng et al.
+2024a; Guo and Yang 2024; Tong et al.
+2024).
+Similarly, in task-oriented LLMs, weak-to-strong learning improves LLM’s ability by enabling strong models to refine their data autonomously, boosting performance without extensive high-quality input (Zhang et al.
+2023; Yang, Ma, and Liu 2024).
+Through weak-to-strong supervision, LLM performance can be significantly improved by iteratively transforming low-quality labels into more reliable ones, leading to more effective model training and robust outputs (Zakershahrak and Ghodratnama 2024; Lang, Sontag, and Vijayaraghavan 2024).
+Recent work also shows weak LLMs can rival human feedback quality, enabling scalable and cost-efficient alignment (Tao and Li 2024; Li et al.
+2025).
+3 PGED: Preference Graph Ensemble and Denoising
+In this section, we first introduce key definitions and assumptions underlying our approach, including the formal definition of a preference graph and the assumption of preference transitivity (Section 3.1).
+Building on these preliminaries, we then present our proposed framework, PGED, which consists of three main steps: (1) Graph Ensemble (Section 3.2), aggregating multiple evaluators’ preference graphs into a single unified structure; (2) Graph Denoising (Section 3.3), removing cycles to ensure an acyclic, consistent preference structure; and (3) Graph-to-Ranking (Section 3.4), converting the denoised graph into a reliable candidate ranking.
+The overall process of PGED is illustrated in Figure 1(c).
+3.1 Preliminarily
+Transitivity of Preferences.
+In preference evaluation, we assume that the ground-truth preference relation is transitive. That is, for any distinct $u,v,w\in V$, if $u\succ v$ and $v\succ w$, then it must hold that $u\succ w$:
+$(u\succ v)\wedge(v\succ w)\Longrightarrow(u\succ w).$ (1)
+This transitivity assumption ensures that pairwise comparisons can be consistently embedded into a global ranking.
+Preference Graph.
 A preference graph is a directed graph $G_{P}=(V,A,w)$, where $V=\{v_{1},v_{2},\ldots,v_{n}\}$ represents $n$ candidates, $A\subseteq V\times V$ is a set of directed arcs indicating pairwise preferences, and $w:A\rightarrow\mathbb{R}^{+}$ assigns weights to arcs, representing preference strength.
-For distinct $u,v\in V$, an arc $(u,v)\in A$ exists if $w(u,v)>0$, where the weight $w(u,v)$ aggregates individual preferences:
-$w(u,v)=\sum_{i=1}^{k}(s_{i}(u,v)-s_{i}(v,u)),$
-with $s_{i}(u,v)$ representing the score from the $i$-th source. The preference graph encapsulates aggregate preferences, with arc weights reflecting cumulative strength.
-3.2 GED: Preference Graph Ensemble and Denoise
-Our method, GED (Preference Graph Ensemble and Denoise), begins by performing graph ensemble to aggregate a set of preference graphs.
-It then applies graph denoising to ensure acyclicity, followed by graph-to-ranking to derive the final node ranking.
-The detailed steps are as follows:
-Graph ensemble.
-Given $k$ weighted graphs $G_{1},\dots,G_{k}$ with shared vertex set $V$, the ensemble graph $G_{E}=(V,A_{E},w_{E})$ is constructed by setting $A_{E}=\bigcup_{i=1}^{k}A_{i}$ and defining the weight $w_{E}(u,v)=\sum_{i=1}^{k}w_{i}(u,v)$ for each arc $(u,v)\in A_{E}$.
-Graph denoising.
-Graph denoising involves transforming the original graph $G=(V,A,w)$ into a DAG.
-This transformation is achieved by identifying and removing a set of arcs known as the Feedback Arc Set (FAS) (Gabow, 1995), which is a set of arcs whose removal makes the graph acyclic.
-The goal is to find a minimum FAS, denoted as $R^{*}(G)$, which is a set of arcs with the smallest total weights that needs to be removed to eliminate all cycles in $G$.
-To find this minimum FAS, we can order the vertices of $G$ in a specific sequence $s=\{v_{1},v_{2},\ldots,v_{n}\}$.
-This vertex sequence induces a FAS $R(s)$, consisting of all arcs that point against the direction of the sequence, i.e., arcs $v_{j}\rightarrow v_{i}$ where $j>i$.
-The graph denoising problem is thus reframed as finding an optimal vertex sequence $s^{*}$ that induces the minimal FAS, such that $R(s^{*})=R^{*}(G)$.
-This optimal sequence $s^{*}$ ensures that the total weights of arcs eliminated to achieve a DAG is minimized.
-Finding a minimum FAS in general is known to be an NP-complete problem, whose computational complexity can be exponential (Karp, 2010; Bodlaender et al., 2012).
-Therefore, in our experiment, we apply the well-established approximation algorithm proposed in Eades et al.
-(1993).
-Details can be found in Appendix M.
-Graph to ranking.
+For distinct $u,v\in V$, an arc $(u,v)\in A$ exists if $w(u,v)>0$.
+3.2 Graph Ensemble
+Given $k$ weighted preference graphs $G_{i}=(V,A_{i},w_{i})$ on a common vertex set $V$, define
+$A_{E}$ $\triangleq\Big\{(u,v)\in V\times V:\ \sum_{i=1}^{k}w_{i}(u,v)>0\Big\},$ (2)
+$w_{E}(u,v)$ $\triangleq\sum_{i=1}^{k}w_{i}(u,v),\quad(u,v)\in A_{E},$
+where $k$ is the total number of preference sources and $w_{i}(u,v)$ is the preference result from the $i$-th source and $w_{i}(u,v)=0$ if $(u,v)\notin A_{i}$.
+The ensemble graph is $G_{E}=(V,A_{E},w_{E})$.
+3.3 Graph Denoising
+To ensure consistency in the aggregated preference graph, we apply a denoising step that transforms a potentially cyclic weighted digraph $G=(V,A,w)$ into a directed acyclic graph (DAG).
+We cast this as the classical weighted feedback arc set (WFAS) problem (Gabow 1995), which seeks a minimum-total-weight set of arcs whose removal renders the graph acyclic.
+Formally,
+$R^{*}(G)\;=\;\arg\min_{R\subseteq A}\sum_{(u,v)\in R}w(u,v)$ (3)
+$\text{s.t.}\quad(V,A\setminus R)\ \text{is acyclic}.$
+A convenient formulation uses vertex sequencing.
+Given an ordering $s=\{v_{1},\ldots,v_{n}\}$, the induced feedback arc set $R(s)$ comprises all arcs that violate the order, i.e., $(v_{j}\to v_{i})$ with $j>i$.
+The denoising problem then reduces to finding $s^{*}$ that minimizes the total weight of backward edges.
+Computing the optimal ordering is NP-hard (Karp 2010), so we adopt an efficient greedy procedure tailored to weighted graphs.
+Our algorithm iteratively builds a total order by removing structurally informative vertices.
+At each step, we first peel all sinks (zero out-degree) and place them at the end of the order, then peel all sources (zero in-degree) and place them at the beginning.
+When neither exists, we choose a vertex $u$ maximizing the weighted degree difference
+$\delta(u)\;=\;d^{+}(u)-d^{-}(u),$ (4)
+where $d^{+}(u)=\sum_{(u,v)\in A}w(u,v)$ and $d^{-}(u)=\sum_{(v,u)\in A}w(v,u)$.
+Repeating until all vertices are removed yields a total order; the backward arcs under this order form an approximate feedback arc set $R(s)$.
+We then construct and return only the denoised graph $G^{\prime}=(V,A\setminus R(s),w)$, which is a subgraph of the input and is typically sparse rather than fully connected.
+The full procedure is summarized in Algorithm 1.
+Its theoretical guarantee can be seen in Appendix M.
+Algorithm 1 Preference Graph Denoising for PGED
+1:Input: Weighted digraph $G=(V,A,w)$
+2:Output: Denoised graph $G^{\prime}=(V,A^{\prime},w)$
+3:Let $A_{0}\leftarrow A$; initialize $s_{1}\leftarrow[\ ],\ s_{2}\leftarrow[\ ]$
+4:while $V\neq\emptyset$ do
+5:  while $\exists$ sink $u$ in $G$ do
+6:   Prepend $u$ to $s_{2}$
+7:   Remove $u$ and its incident arcs from $G$
+8:  end while
+9:  while $\exists$ source $u$ in $G$ do
+10:   Append $u$ to $s_{1}$
+11:   Remove $u$ and its incident arcs from $G$
+12:  end while
+13:  if $V=\emptyset$ then break
+14:  end if
+15:  Select $u=\arg\max_{v\in V}\big(d^{+}(v)-d^{-}(v)\big)$
+16:  Append $u$ to $s_{1}$
+17:  Remove $u$ and its incident arcs from $G$
+18:end while
+19:Concatenate $s=s_{1}\|s_{2}$
+20:Compute $R(s)=\{(v_{j}\to v_{i})\in A_{0}:j>i\text{ in }s\}$
+21:Set $A^{\prime}=A_{0}\setminus R(s)$
+22:return $G^{\prime}=(V,A^{\prime},w)$
+3.4 Graph to ranking
 Given a DAG $G=(V,A,w)$, we derive a ranking by computing the descendant count $\text{desc}(v)$ for each vertex $v$, defined as the number of vertices reachable from $v$:
-$\text{desc}(v)=\left|\{u\in V:v\rightarrow u\}\right|,$
+$\text{desc}(v)=\left|\{u\in V:v\rightarrow u\}\right|,$ (5)
 where $v\rightarrow u$ denotes a directed path. Vertices are ranked based on $\text{desc}(v)$, with ties broken lexicographically:
-$v_{1}\succ v_{2}\succ\cdots\succ v_{n}.$
+$v_{1}\succ v_{2}\succ\cdots\succ v_{n}.$ (6)
 This ranking reflects both individual preferences and their relative strengths in the graph.
-3.3 Applications
-We apply GED to three tasks: Response Selection (selecting the best response from LLM-generated candidates), Model Ranking (ranking models based on task performance), and Model Alignment (identifying the best instruction-response pairs for training).
-Response selection.
-In this task, a model $\mathcal{M}$ generates $n$ candidate answers $\{ans_{1},\ldots,ans_{n}\}$ for each question $q\in Q$, and the goal is to identify the optimal answer $ans_{q}^{*}$.
-Multiple evaluators $\mathcal{A}=\{a_{1},\ldots,a_{k}\}$ provide pairwise preferences among these candidates, constructing a set of preference graphs $\{G_{a}:a\in\mathcal{A}\}$.
-Each graph $G_{a}=(V_{q},A_{a},w_{a})$ represents the preferences of evaluator $a$, where $V_{q}$ corresponds to the candidates and $A_{a}$ indicates pairwise preferences with weights $w_{a}$ reflecting preference strength.
-This process is detailed in Appendix E.
-To aggregate these graphs, GED first merges them into a unified graph $G_{q}=(V_{q},A_{q},w_{q})$, removes cycles to obtain a DAG, and derives the ranking $\mathcal{R}_{q}=\{v_{1}\succ v_{2}\succ\ldots\succ v_{n}\}$.
-The top-ranked answer is selected as $ans_{q}^{*}$.
-Repeating this for all $q\in Q$ yields the final set $ans^{*}=\{ans_{1}^{*},\ldots,ans_{t}^{*}\}$, representing consensus from multiple evaluators and ensuring high-quality responses.
-Model Ranking.
-The goal of model ranking is to rank a set of models $M=\{\mathcal{M}_{1},\ldots,\mathcal{M}_{n}\}$ based on their responses to questions $Q=\{q_{1},\ldots,q_{t}\}$.
-Evaluators $\mathcal{A}=\{a_{1},\ldots,a_{k}\}$ provide pairwise preferences for model outputs, constructing preference graphs $\{G_{a}:a\in\mathcal{A}\}$ for each question $q$.
-Here, each graph $G_{a}=(V_{q},A_{a},w_{a})$ represents preferences, where $V_{q}$ corresponds to the models and $w_{a}$ reflects preference strength.
-The detailed procedure is outlined in Appendix E.
-Using GED, we aggregate these graphs into a unified graph $G_{q}$, transform it into a DAG by removing cycles, and derive the ranking $\mathcal{R}_{q}$.
-Repeating this for all $q\in Q$ yields a set of rankings $\{\mathcal{R}_{q}:q\in Q\}$.
-Finally, a ranking ensemble is applied to compute the overall ranking $\mathcal{R}^{*}$ across questions, reflecting model performance as assessed by evaluators, as detailed in Appendix N
-Model Alignment.
-In the model alignment task, the objective is to identify the best response $y^{*}$ for each instruction $x$ from candidate responses $\{y_{1},\ldots,y_{n}\}$.
-Evaluators $\mathcal{A}$ provide pairwise preferences, forming preference graphs $\{G_{a}:a\in\mathcal{A}\}$, where $G_{a}=(V_{x},A_{a},w_{a})$ represents preferences over $\{y_{1},\ldots,y_{n}\}$.
-GED aggregates these graphs into a unified graph $G_{x}$, removes cycles to produce a DAG, and derives a ranking $\mathcal{R}_{x}$.
-The highest-ranked response is selected as $y^{*}$ for $x$.
-Repeating this for all $t$ instructions yields the final training set $\{(x_{1},y_{1}^{*}),\ldots,(x_{t},y_{t}^{*})\}$, reflecting consensus among evaluators.
+4 Downstream Tasks
+We apply PGED to three tasks: Response Selection (selecting the best response from LLM-generated candidates), Model Ranking (ranking models based on task performance), and Model Alignment (identifying the best instruction-response pairs for training).
+Details on how preference graphs are constructed for each of the three settings can be found in Appendix H.
+4.1 Response Selection
+For each question $q\in Q$, a model $\mathcal{M}$ generates $n$ candidate answers $C_{q}=\{ans_{1},\ldots,ans_{n}\}$.
+A set of evaluators $\mathcal{A}=\{a_{1},\ldots,a_{k}\}$ provides pairwise preferences over $C_{q}$.
+For each evaluator $a$, we construct a weighted preference graph $G_{a}^{q}=(V_{q},A_{a},w_{a})$, where $V_{q}=\{v_{1},\ldots,v_{n}\}$ indexes candidates and $w_{a}(u,v)$ accumulates the number of wins of $u$ over $v$ (ties are ignored).
+We then apply PGED: (i) Graph Ensemble aggregates $\{G_{a}^{q}\}_{a\in\mathcal{A}}$ into a single graph $G_{E}^{q}$ (Section 3.2); (ii) Graph Denoising removes a minimum-weight set of feedback arcs to produce a DAG $G^{\prime}_{q}$ (Section 3.3); and (iii) Graph-to-Ranking converts $G^{\prime}_{q}$ into a total order $\mathcal{R}_{q}=\{v_{1}\succ\cdots\succ v_{n}\}$ (Section 3.4).
+The top-ranked candidate is selected as $ans_{q}^{*}$.
+Repeating this for all $q\in Q$ yields $ans^{*}=\{ans_{1}^{*},\ldots,ans_{t}^{*}\}$.
+4.2 Model Ranking
+Given a set of models $M=\{\mathcal{M}_{1},\ldots,\mathcal{M}_{n}\}$ and questions $Q=\{q_{1},\ldots,q_{t}\}$, our goal is to produce a global ranking of $M$.
+For each question $q\in Q$ and evaluator $a\in\mathcal{A}$, we construct a weighted preference graph $G_{a}^{q}=(V_{q},A_{a},w_{a})$, where $V_{q}=\{m_{1},\ldots,m_{n}\}$ indexes model outputs $\mathcal{M}_{i}(q)$ and $w_{a}(u,v)$ accumulates wins of $u$ over $v$ (ties are ignored).
+We then apply PGED per question: (i) Graph Ensemble aggregates $\{G_{a}^{q}\}_{a\in\mathcal{A}}$ into $G_{E}^{q}$ (Section 3.2); (ii) Graph Denoising yields a DAG $G^{\prime}_{q}$ (Section 3.3); and (iii) Graph-to-Ranking returns a total order $\mathcal{R}_{q}$ over $V_{q}$ (Section 3.4).
+Finally, we aggregate $\{\mathcal{R}_{q}:q\in Q\}$ into an overall model ranking $\mathcal{R}^{*}$ using a ranking-ensemble procedure (Appendix Q).
+4.3 Model Alignment
+For each instruction $x\in X$, let the candidate set be $Y_{x}=\{y_{1},\ldots,y_{n}\}$.
+Evaluators $\mathcal{A}$ provide pairwise preferences over $Y_{x}$.
+For each $a\in\mathcal{A}$, we build a weighted preference graph $G_{a}^{x}=(V_{x},A_{a},w_{a})$, where $V_{x}=\{v_{1},\ldots,v_{n}\}$ indexes responses and $w_{a}(u,v)$ aggregates wins of $u$ over $v$ (ties ignored ).
+Applying PGED, ensemble, denoising, and graph-to-ranking, produces a total order $\mathcal{R}_{x}$; we select the top response $y_{x}^{*}$ for instruction $x$.
+Repeating this for all $x\in X$ yields the aligned training set $\{(x,y_{x}^{*}):x\in X\}$.
 5 Theoretical Analysis
-In this section, we provide a theoretical foundation for our method, showing that by modeling preference graphs as random perturbations of a ground truth DAG, GED can reliably recover the true structure through graph ensemble and denoising with high probability, demonstrating its robustness in handling noisy evaluations.
+In this section, we provide a theoretical foundation for our method, showing that by modeling preference graphs as random perturbations of a ground truth DAG, PGED can reliably recover the true structure through graph ensemble and denoising with high probability, demonstrating its robustness in handling noisy evaluations.
 Theoretically, we treat each of our preference graph as a random perturbation of some ground truth DAG $G=(V,A)$.
@@
 That is, each edge in $E$ has probability $\delta_{1}$ of being flipped and each pair of unconnected nodes has probability $\delta_{2}$ of being connected with a random direction.
+Now, given that $G_{1},\dots,G_{N}\overset{\text{i.i.d.}}{\sim}\mathcal{G}(G,\delta_{1},\delta_{2})$, we will show that to some extent our combination of graph ensemble and graph denoising can indeed provably recover the ground truth DAG $G$.
+For simplicity, all edges in $G_{1},\dots,G_{N}$ and $G$ are considered equal weighted.
+Meanwhile, we use $\text{MAS}(\cdot)$ to denote the graph obtained by denoising, which stands for the maximum acyclic subgraph (MAS).
+Then, we have the following theorem.
+Theorem 1
+Suppose $G_{1},\dots,G_{N}\overset{\text{i.i.d.}}{\sim}\mathcal{G}(G,\delta_{1},\delta_{2})$ for some ground truth $G=(V,A)$.
+Let $\widehat{G}$ be the graph ensembled from $G_{1},\dots,G_{N}$ by operations defined in Section 3.
+Then, as long as $\delta_{1}=0.5-\epsilon$ for some $\epsilon>0$, we have
+$\begin{split}\mathbb{P}\left(G\subseteq\text{MAS}(\widehat{G})\right)&\geq 1-2|A|\exp\left(-\frac{N\epsilon^{2}}{2}\right)\\ &\quad-2U\exp\left(-\frac{N\epsilon^{2}}{6U^{2}\delta_{2}+2U\epsilon}\right),\end{split}$
 […]
-Theorem 1.
-[…]
-$\begin{split}\mathbb{P}\left(G\subseteq\text{MAS}(\widehat{G})\right)&\geq 1-2|A|\exp\left(-\frac{N\epsilon^{2}}{2}\right)\\ &\quad-2U\exp\left(-\frac{N\epsilon^{2}}{6U^{2}\delta_{2}+2U\epsilon}\right),\end{split}$
-where $G\subseteq\text{MAS}(\widehat{G})$ represents that $G$ is a subgraph of $\text{MAS}(\widehat{G})$ and $U=\frac{|V|(|V|-1)}{2}-|A|$ is the number of pairs of unconnected nodes in $G$.
-The full proof is given in Appendix G.
+The full proof is given in Appendix J.
 From the theorem, we can see that the probability of failure decreases exponentially as the number of samples $N$ increases.
 Meanwhile, this guarantee only requires $\delta_{1}<0.5$ and does not place restrictions on $\delta_{2}$, which are very mild conditions.
-Table 1: Performance comparison of response selection methods across five benchmarks.
-GED consistently outperforms baseline methods, demonstrating the effectiveness of graph denoising and the aggregation of multiple evaluators.
 Method HumanEval AlpacaEval MATH GSM8k GAIA Avg
@@
 Qwen2-72B 57.93 29.58 72.75 84.67 11.52 51.29
-ContraSolver 65.42 31.12 74.95 86.84 12.22 54.11
+ContraSolver (Qwen2-72B) 65.42 31.12 74.95 86.84 12.22 54.11
 ListPreference 61.52 31.67 71.75 85.0 10.90 52.16
 Single model Self-consistency 60.98 29.33 73.58 84.91 8.86 51.53
+Elo 62.21 31.34 75.01 87.21 12.36 53.63
+Bradley-Terry 62.83 32.91 74.97 86.64 12.07 53.88
+Merge Sort 62.81 31.88 74.83 86.90 12.13 53.71
 Llama3-8B 62.19 29.31 74.27 83.16 11.31 52.04
@@
 Single evaluator with graph denoising 68.90 31.17 75.33 87.45 12.26 55.02
-Multi-MV 66.18 29.57 74.77 86.42 11.72 53.73
-GED (w/o denoising) 69.25 30.98 74.29 87.17 12.68 54.87
-Multiple evaluator GED 70.73 32.44 75.58 88.18 13.33 56.05
-Figure 2: Comparison of GED with GPT-3.5, GPT-4-o-mini, and GPT-4-o on 100 randomly selected tasks.
-GED consistently outperforms GPT-3.5 across all tasks and surpasses GPT-4-o-mini on challenging tasks like HumanEval and GSM8k, showcasing the effectiveness of weak evaluator aggregation with graph denoising.
-5 Experiments on Response Selection
+Majority Voting 66.18 29.57 74.77 86.42 11.72 53.73
+Weight Majority Voting 66.43 30.12 74.79 86.68 11.84 53.97
+PRD 66.53 30.33 74.89 86.75 11.83 54.07
+ChatEval 66.51 30.26 74.79 86.85 11.92 54.07
+SPA 66.32 30.05 74.82 86.57 11.81 53.91
+PGED(w/o denoising) 69.25 30.98 74.29 87.17 12.68 54.87
+Multiple evaluator PGED 71.86 32.95 76.57 89.76 13.74 56.98
+Table 1: Performance comparison of response selection methods across five benchmarks.
+PGED consistently outperforms baseline methods, highlighting the effectiveness of graph denoising and multi-evaluator aggregation.
+Single model denotes directly using the evaluator to answer the question.
+Single evaluator and Multi evaluator refer to selecting the best response from ten Qwen2-72B-generated candidates using a single or multiple evaluators, respectively.
+6 Response Selection for Test-time Scaling
 Experiment Setup.
-In this section, we evaluate the performance of GED on five benchmarks: HumanEval (Chen et al., 2021), AlpacaEval (Li et al., 2023b), MATH (Hendrycks et al., 2021), GSM8k (Chen et al., 2021), and GAIA (Mialon et al., 2023).
-The Qwen2-72B (Yang et al., 2024a) model ($\mathcal{M}$) generates ten candidate responses per question, and we assess the effectiveness of different methods in selecting the best response.
-For further implementation details, see Appendix A.
+In this section, we evaluate the performance of PGED on five benchmarks: HumanEval (Chen et al.
+2021), AlpacaEval (Li et al.
+2023b), MATH (Hendrycks et al.
+2021), GSM8k (Chen et al.
+2021), and GAIA (Mialon et al.
+2023).
+The Qwen2-72B (Yang et al.
+2024) model ($\mathcal{M}$) generates ten candidate responses per question, and we assess the effectiveness of different methods in selecting the best response.
 We evaluate performance using three setups.
-First, in the single model setting, the baselines include ContraSolver(Zhang et al., 2024b), Self-consistency(Wang et al., 2022), and direct evaluation with models (Llama3-8B, Mistral-7B, Qwen2-7B and Qwen2-72B).
+First, in the single model setting, the baselines include ContraSolver(Zhang, Yin, and Wan 2024), Self-consistency(Wang et al.
+2022), and direct evaluation with models (Llama3-8B, Mistral-7B, Qwen2-7B and Qwen2-72B).
 Additionally, we include a baseline called ListPreference, where instead of pairwise comparisons, all candidate responses are input into Qwen2-72B for selecting the most appropriate response.
-Then, in the single evaluator setting, individual evaluators (Llama3-8B, Mistral-7B, Qwen2-7B, Qwen2-72B) select the best response from $\mathcal{M}$’s outputs, with and without applying GED’s graph denoising.
-Finally, in the multiple evaluators setup, we combine three small evaluators (Llama3-8B, Qwen2-7B, Mistral-7B) to select responses from Qwen2-72B with GED.
-We also introduce a baseline, Multi-MV, which selects the response that receives the most votes from evaluators in pairwise comparisons.
-We present the results of GED and its variant (w/o denoising), which ensembles the preference graphs without the denoising step.
-Table 2: Results of the model ranking task, evaluated using Ranking Correction. Higher correlation values indicate a stronger alignment with the ground truth rankings.
+Then, in the single evaluator setting, individual evaluators (Llama3-8B, Mistral-7B, Qwen2-7B, Qwen2-72B) select the best response from $\mathcal{M}$’s outputs, with and without applying PGED’s graph denoising.
+We additionally report results for three classical baselines in this setting: Elo (Jiang, Ren, and Lin 2023), Bradley–Terry (BTL) (Bradley and Terry 1952; Lu and Boutilier 2011), and MergeSort (Shyamasundar 1996).
+Finally, in the multiple evaluators setup, we combine three small evaluators (Llama3-8B, Qwen2-7B, Mistral-7B) to select responses from Qwen2-72B with PGED.
+We additionally include four aggregation baselines: Majority Voting, Weighted Majority Voting, PRD (Li, Patel, and Du 2023a), ChatEval (Chan et al.
+2023a), and SPA (Kadekodi, McTavish, and Ustun 2025).
+We present the results of PGED and its variant (w/o denoising), which ensembles the preference graphs without the denoising step.
+Note that, among the baselines we report: (i) in the single model setting, ContraSolver, ListPreference, and Self-consistency use Qwen2-72B as the base model; (ii) in the single evaluator setting, Elo, Bradley–Terry, and MergeSort use Qwen2-72B as the base model; and (iii) in the multiple evaluators setting, Majority Voting, Weighted Majority Voting, PRD, ChatEval and SPA use Llama3-8B, Mistral-7B, and Qwen2-7B as the base model set.
+For details on the datasets and baselines, please refer to Appendix A.
+Figure 2: Comparison of PGED with GPT-3.5, GPT-4-o-mini, and GPT-4-o on 100 randomly selected tasks.
+PGED consistently outperforms GPT-3.5 across all tasks and surpasses GPT-4-o-mini on challenging tasks like HumanEval and GSM8k, showcasing the effectiveness of weak evaluator aggregation with graph denoising.
+Main results.
+Table 1 presents the results of the response selection task across five benchmarks.
+PGED consistently outperforms all baseline methods, including single model evaluations (single model), direct response selection by individual models (single evaluator) and Multiple evaluator.
+This demonstrates the strength of aggregating preference evaluators with PGED, particularly when coupled with graph denoising, which enhances response quality by filtering out noise and biases.
+Furthermore, by combining preference graphs derived from smaller models (Llama3-8B, Mistral-7B, Qwen2-7B), PGED outperforms a much larger evaluator (Qwen2-72B).
+This underscores the value of ensemble methods in mitigating the limitations of individual evaluators.
+Then, the denoising process proves to be crucial for improving consistency and overall response quality.
+The substantial performance gains observed when using PGED with denoising, compared to both the single evaluator setup and the ensemble without denoising, highlight its importance in refining response selection.
+For Majority Voting, while it improves upon individual evaluators, it still underperforms PGED, highlighting PGED’s ability to capture nuanced evaluation signals and reduce inconsistencies.
+Additionally, we observed that the ListPreference baseline performed worse than Qwen2-72B as single evaluator, likely due to LLM limitations in handling long-text.
+Lastly, to further evaluate PGED, we compared its performance with GPT-3.5, GPT-4-o-mini, and GPT-4-o.
+Due to computational and API cost constraints, we limited the evaluation to 100 data points for each task.
+As shown in Figure 2, PGED consistently outperformed GPT-3.5 across all tasks and surpassed GPT-4-o-mini on challenging benchmarks like HumanEval and GSM8k.
+These results highlight the superiority of PGED, particularly in leveraging multi-preference evaluators and graph denoising to outperform individual state-of-the-art models.
+More discussion on cost is provided in Appendix R.
 Model Weight Score Kemeny
@@
 Single evaluator with graph denoising 65.47 70.06 69.92 73.39 73.21 70.41
-GED (w/o ensemble) 62.82 68.44 68.44 69.34 67.34 67.27
-GED (w/o denoising) 64.84 69.23 69.81 75.35 74.37 70.72
-Multiple evaluator GED 66.59 71.14 71.14 77.17 76.46 72.50
+PGED(w/o ensemble) 62.82 68.44 68.44 69.34 67.34 67.27
+PGED(w/o denoising) 64.84 69.23 69.81 75.35 74.37 70.72
+Multiple evaluator PGED 66.59 71.14 71.14 77.17 76.46 72.50
+Table 2: Results of the model ranking task, evaluated using Ranking Correction. Higher correlation values indicate a stronger alignment with the ground truth rankings.
+7 Model Ranking for Evaluation
+Experiment Setup.
+[…]
 Main results.
-Table 1 presents the results of the response selection task across five benchmarks.
-GED consistently outperforms baseline methods, including both single model evaluations (single model) and direct response selection by individual models (single evaluator).
-This demonstrates the strength of aggregating weak evaluators with GED, particularly when coupled with graph denoising, which enhances response quality by filtering out noise and biases.
-Furthermore, by combining preference graphs derived from smaller models (Llama3-8B, Mistral-7B, Qwen2-7B), GED outperforms a much larger evaluator (Qwen2-72B).
-This underscores the value of ensemble methods in mitigating the limitations of individual evaluators.
-Then, the denoising process proves to be crucial for improving consistency and overall response quality.
-The substantial performance gains observed when using GED with denoising, compared to both the single evaluator setup and the ensemble without denoising, highlight its importance in refining response selection.
-For Multi-MV, while it improves upon individual evaluators, it still underperforms GED, highlighting GED’s ability to capture nuanced evaluation signals and reduce inconsistencies.
-Additionally, we observed that the ListPreference baseline performed worse than Qwen2-72B as single evaluator, likely due to LLM limitations in handling long-text.
-Lastly, to further evaluate GED, we compared its performance with GPT-3.5, GPT-4-o-mini, and GPT-4-o.
-Due to computational and API cost constraints, we limited the evaluation to 100 data points for each task.
-As shown in Figure 2, GED consistently outperformed GPT-3.5 across all tasks and surpassed GPT-4-o-mini on challenging benchmarks like HumanEval and GSM8k.
-These results highlight the superiority of GED, particularly in leveraging multi-weak evaluators and graph denoising to outperform individual state-of-the-art models.
-Ablation study.
-We evaluate the impact of removing the ensembling step in GED, referred to as the (w/o ensemble) variant.
-In this case, individual evaluators’ preference graphs are denoised and converted to rankings, which are then aggregated using methods such as Weight Score, Kemeny, Weighted Kemeny, Pairwise Majority, and Weighted Pairwise Majority (detailed in Appendix N).
-For simplicity of presentation, we use Weight Score to represent GED (w/o ensemble) (Weight Score).
-As shown in Figure 3, all (w/o ensemble) methods consistently underperform compared to GED.
-This performance gap arises because converting graphs to ranks before aggregation leads to information loss.
-In contrast, GED ensembles the graphs directly, preserving more detailed preference information and resulting in better final rankings.
-6 Experiments on Model Ranking
-Experiment Setup.
-In this section, we evaluate the effectiveness of GED in the model ranking task within a human preference setting, using the AlpacaEval benchmark (Li et al., 2023b).
-We employ 30 widely used models from the AlpacaEval dataset as our model set $\mathcal{M}$, while the benchmark’s questions form the question set $Q$.
-The rankings provided by the AlpacaEval benchmark serve as ground truth for evaluating the accuracy of various ranking methods.
-This is justified by AlpacaEval’s strong correlation with Chatbot Arena rankings, making it a reasonable proxy for human judgments (Dubois et al., 2024a).
-We adopt Ranking Correction, measured by the Spearman rank correlation coefficient, to evaluate the similarity.
-To generate rankings, we utilize outputs from the open-source models Llama3-70B, Qwen2-72B, Mistral-8$\times$7B, and Qwen1.5-72B as our evaluators.
-For further implementation details, see Appendix A.
-We investigate two variants of GED: (w/o ensemble) denoises the preference graphs from different evaluators for the same question, converts each into a ranking, and then ensembles these rankings to produce the final output, while (w/o denoising) directly ensembles the preference graphs to obtain the final ranking without denoising.
-Figure 3: Comparison of GED and (w/o ensemble) variants.
-GED outperforms due to preserving more information by directly ensembling preference graphs, while rank aggregation in the (w/o ensemble) methods leads to performance loss.
-Main results.
-The results, presented in Table 2, show that GED outperforms all single-model baselines, highlighting the significant improvement in ranking accuracy achieved by leveraging preference information from multiple evaluators.
-Moreover, GED surpasses the (w/o ensemble) variant, indicating that generating rankings through graph ensemble first prevents information loss compared to converting individual graphs into rankings.
+The results, presented in Table 2, show that PGED outperforms all single-model baselines, highlighting the significant improvement in ranking accuracy achieved by leveraging preference information from multiple evaluators.
+Moreover, PGED surpasses the (w/o ensemble) variant, indicating that generating rankings through graph ensemble first prevents information loss compared to converting individual graphs into rankings.
 When the ensemble graph is not denoised (w/o denoising), residual noise can adversely affect the final ranking quality.
 Additionally, our denoising method also enhances results in single-model settings.
-Figure 4: Performance comparison of different methods (Random, Longest, ContraSolver, and GED) across multiple benchmarks, including LIMA, Vicuna, Koala, WizardLM and Self-Instruct.
-The results show GED effectively filters low-quality responses, improving performance and model alignment over baselines.
-Table 3: Performance comparison of different methods (Random, Longest, ContraSolver, and GED) on model alignment task across the HH-RLHF benchmark.
-The results demonstrate the superiority of GED in consistently selecting high-quality responses, leading to improved model performance compared to baseline methods.
-Base Model Harmless (base) Helpful (base) Helpful (online) Helpful (rejection) Avg.
-Origin 69.67 61.12 65.41 64.06 65.07
-Random 69.38 62.87 66.75 65.57 66.14
-Longest 69.65 63.54 66.99 66.43 66.65
-ContraSolver 69.57 63.61 66.87 66.59 66.66
-Llama-2-7B GED 69.71 64.10 67.87 67.01 67.17
-Origin 61.59 59.51 65.21 63.17 62.37
-Random 59.15 59.61 64.06 62.38 61.30
-Longest 61.81 60.53 64.52 63.22 62.52
-ContraSolver 61.48 59.85 64.66 63.41 62.85
-Mistral-7B GED 61.96 60.71 65.49 63.82 63.50
-7 Experiments on Instruct Tuning
+Figure 3: Performance comparison of different methods (Random, Longest, ContraSolver, and PGED) across multiple benchmarks.
+The results show PGED effectively filters low-quality responses, improving performance and model alignment over baselines.
+8 Data Selection for Model Fine-tuning
 Experiment Setup.
-In this section, we explore the effects of various data selection methods for model alignment on Llama-2-7B (Touvron et al., 2023) and Mistral-7B (Jiang et al., 2023) through instruct tuning.
-Specifically, we randomly sampled 5,000 data points from UltraFeedback (Cui et al., 2023) and used Qwen1.5-14B (Yang et al., 2024a) to generate eight responses per data point as instruct data.
-We then applied four different methods—Random, Longest (Zhao et al., 2024), ContraSolver (using Qwen2-72B as the evaluator) (Zhang et al., 2024b), and our proposed GED, which leverages Llama3-8B, Mistral-7B, and Qwen2-7B as evaluators—to select a subset of these responses for model alignment training.
+Base Model Harmless (base) Helpful (base) Helpful (online) Helpful (rej.) Avg.
+Origin 69.68 61.65 64.84 63.37 64.89
+Random 69.85 63.53 67.00 65.85 66.56
+Longest 69.32 64.15 66.56 66.30 66.58
+ContraSolver 69.75 63.28 66.71 65.86 66.40
+Llama-2-7B
+Ours 72.27 65.56 69.44 67.60 68.72
+Origin 61.30 59.95 65.21 63.75 62.55
+Random 59.85 59.34 63.38 61.87 61.11
+Longest 61.66 59.73 64.92 62.92 62.31
+ContraSolver 60.87 59.83 64.23 63.26 62.05
+Mistral-7B
+Ours 64.57 64.48 69.66 67.73 66.61
+Table 3: Performance comparison of different methods on model alignment task across the HH-RLHF benchmark.
+In this section, we explore the effects of various data selection methods for model alignment on Llama-2-7B (Touvron et al.
+2023) and Mistral-7B (Jiang et al.
+2023) through instruct tuning.
+Specifically, we randomly sampled 5,000 data points from UltraFeedback (Cui et al.
+2023) and used Qwen1.5-14B (Yang et al.
+2024) to generate eight responses per data point as instruct data.
+We then applied four different methods—Random, Longest (Zhao et al.
+2024), ContraSolver (using Qwen2-72B as the evaluator) (Zhang, Yin, and Wan 2024), and our proposed PGED, which leverages Llama3-8B, Mistral-7B, and Qwen2-7B as evaluators—to select a subset of these responses for model alignment training.
 The Origin refers to the performance of the base model without alignment.
-The models were evaluated on the HH-RLHF (Bai et al., 2022) benchmark, which comprises four subsets: Harmless (base), Helpful (base), Helpful (online), and Helpful (rejection).
-For evaluation, we employed the same Reward model as in prior work (Song et al., 2024; Yu et al., 2023) to quantify human preference levels.
+The models were evaluated on the HH-RLHF (Bai et al.
+2022) benchmark, which comprises four subsets: Harmless (base), Helpful (base), Helpful (online), and Helpful (rejection).
+For evaluation, we employed the same Reward model as in prior work (Song et al.
+2024; Yu et al.
+2023) to quantify human preference levels.
 The results are presented in Table 3.
-For further implementation details, see Appendix A.
-To ensure a comprehensive assessment, we further evaluated the models—using the Llama-2-7B backbone—on additional benchmarks, including LIMA (Zhou et al., 2023), Vicuna (Chiang et al., 2023), Koala (Vu et al., 2023), WizardLM (Xu et al., 2023), and Self-Instruct (Wang et al., 2022), in accordance with recent studies (Chen et al., 2023b; Zhang et al., 2024a; Hu et al., 2024).
+To ensure a comprehensive assessment, we further evaluated the models—using the Llama-2-7B backbone—on additional benchmarks, including AlpacaEval 2.0 (Li et al.
+2023b), LIMA (Zhou et al.
+2023), Koala (Vu et al.
+2023), and Self-Instruct (Wang et al.
+2022), in accordance with recent studies (Chen et al.
+2023b; Zhang et al.
+2024; Hu et al.
+2024).
 The corresponding results are summarized in Figure 3.
+For details on the datasets and baselines, please refer to Appendix A.
 Main results.
-From Table 3, we observe that GED consistently outperforms all baseline methods, demonstrating its effectiveness in selecting high-quality responses when multiple answers are available for a given instruction.
+From Table 3, we observe that PGED consistently outperforms all baseline methods, demonstrating its effectiveness in selecting high-quality responses when multiple answers are available for a given instruction.
 When faced with multiple responses $y_{1},y_{2},\ldots,y_{n}$ for a given instruction $x$, the Random selection method can have a detrimental impact, especially when the quality of the responses is inconsistent.
@@
 While longer responses may provide more detailed answers, they are not necessarily better in terms of quality, particularly when both high-quality and low-quality answers exist for the same question.
-This is reflected in the results where the Longest method underperforms compared to both ContraSolver and GED, emphasizing that response length alone is not always a reliable criterion.
+This is reflected in the results where the Longest method underperforms compared to both ContraSolver and PGED, emphasizing that response length alone is not always a reliable criterion.
 From Figure 3, we observe similar trends as in Table 3.
-GED consistently outperforms all baselines across various datasets, demonstrating its effectiveness in selecting high-quality responses.
+PGED consistently outperforms all baselines across various datasets, demonstrating its effectiveness in selecting high-quality responses.
 Notably, in AlpacaEval and Self-Instruct, the Random baseline performs worse than the Origin model, highlighting that when response quality varies significantly, poor selection can degrade model performance.
-In contrast, GED leverages preference graphs and denoising techniques to filter out low-quality responses, ensuring more robust and reliable performance, particularly in settings with inconsistent responses, as it removes evaluation noise and leads to more robust performance.
-Table 4: Performance comparison on nuanced quality metrics (%).
-GED outperforms individual evaluators and random selection across Factuality, Relevance, Coherence, Inform., Helpful.
-and Validity metrics.
-Method Factuality Relevance Coherence Inform. Helpful. Validity Avg.
-Random 86.73 87.91 92.47 77.62 17.48 48.92 68.52
-Llama3-8B 88.59 89.91 94.41 79.77 18.48 50.92 70.35
-Mistral-7B 89.10 90.29 94.85 79.95 18.55 51.13 70.65
-Qwen2-7B 89.25 90.44 95.03 80.09 18.58 51.21 70.77
-GED 94.73 95.91 97.36 86.62 19.48 55.92 75.00
-8 Evaluating GED on More Metrics
-To assess GED on more metrics, we evaluated its impact on nuanced quality aspects of LLM outputs, including factuality, relevance, coherence, informativeness, helpfulness, and validity.
-Since LLM tasks often require subtle judgments, GED’s adaptability is crucial.
-Following (Chen et al., 2023a), we used Llama3-70B to generate ten candidate responses per query and applied GED for response selection, with Llama3-8B, Mistral-7B, and Qwen2-7B as evaluators.
-Table 4 shows that GED consistently outperformed individual evaluators and random selection across all metrics.
-It improved factuality by  5 % points over the best evaluator, while also enhancing relevance, coherence, informativeness, and helpfulness.
-These results highlight GED’s ability to aggregate preferences effectively, capturing subtle qualities of generated content and improving overall reliability and utility in LLM assessments.
+In contrast, PGED leverages preference graphs and denoising techniques to filter out low-quality responses, ensuring more robust and reliable performance, particularly in settings with inconsistent responses, as it removes evaluation noise and leads to more robust performance.
+9 Case Study
+(a) Raw Preference Graph 1
+(b) Denoised Preference Graph 1
+(c) Raw Preference Graph 2
+(d) Denoised Preference Graph 2
+Figure 4: Case studies showcasing the raw and denoised preference graphs.
+In this section, we present two HumanEval case studies illustrating PGED on evaluator-aggregated graphs.
+Raw graphs are built from Llama3-8B, Mistral-7B, and Qwen2-7B judging ten Qwen2-72B responses (nodes 0–9).
+Figure 4 contrasts raw graphs (a, c) with denoised DAGs (b, d); dashed edges indicate removals.
+In Case 1, the raw graph contains cycles (e.g., $9\succ 4\succ 3\succ 9$, $9\succ 4\succ 3\succ 2\succ 9$).
+Removing a single edge $4\to 3$ breaks all cycles, yielding the DAG in Figure 4(b).
+In Case 2, cycles such as $7\succ 0\succ 6\succ 7$ and $7\succ 1\succ 0\succ 6\succ 7$ are eliminated by deleting $0\to 6$, producing the DAG in Figure 4(d).
+In both cases, alternative single-edge deletions fail to restore acyclicity without further removals.
+PGED thus identifies a minimal feedback arc set, preserving information while ensuring consistency.
 10 Conclusion
-In this paper, we presented GED, a framework designed to address inconsistencies in pairwise preference evaluations by LLMs.
-By employing graph ensemble techniques and denoising, GED reduces cyclic patterns and enhances the reliability of evaluation outcomes.
-Our theoretical analysis shows that GED can recover the ground truth DAG under reasonable conditions, improving consistency in preference rankings.
+In this paper, we presented PGED, a framework designed to address inconsistencies in pairwise preference evaluations by LLMs.
+By employing graph ensemble techniques and denoising, PGED reduces cyclic patterns and enhances the reliability of evaluation outcomes.
+Our theoretical analysis shows that PGED can recover the ground truth DAG under reasonable conditions, improving consistency in preference rankings.
 Extensive experiments across response selection, model ranking, and instruct tuning demonstrate the efficacy of our method.
-GED consistently outperformed baseline methods in both single-evaluator and multi-evaluator settings, particularly in scenarios where combining small evaluators led to superior results over larger individual evaluators.
-Future work will explore extending GED to broader evaluation frameworks and applying its principles to more complex decision-making tasks, including multi-agent systems and human-AI interaction.
-Impact Statements
-This work enhances the reliability of model-based preference evaluation, reducing inconsistencies in AI judgments.
-It improves fairness and transparency in AI evaluation, benefiting NLP applications and ethical decision-making frameworks.
+PGED consistently outperformed baseline methods in both single-evaluator and multi-evaluator settings, particularly in scenarios where combining small evaluators led to superior results over larger individual evaluators.
+Future work will explore extending PGED to broader evaluation frameworks and applying its principles to more complex decision-making tasks, including multi-agent systems and human-AI interaction.
 Appendix A Implementation Details
@@
 The references to Llama-2-7B, Llama3-70B, Llama3-8B, Mistral-7B, Mistral-8×7B, Qwen2-7B, and Qwen2-72B in the main text refer to the specific models: Llama-2-7b-chat-hf, Meta-Llama-3-70B-Instruct, Meta-Llama-3-8B-Instruct, Mixtral-7B-Instruct-v0.3, Mixtral-8x7B-Instruct-v0.1, Qwen2-7B-Instruct, and Qwen2-72B-Instruct.
-We utilized the reward model oasst-rm-2-pythia-6.9b-epoch-1 following prior works (Song et al., 2024; Yu et al., 2023).
+We utilized the reward model oasst-rm-2-pythia-6.9b-epoch-1 following prior works (Song et al.
+2024; Yu et al.
+2023).
 Each experiment was repeated three times, and the average performance was reported as the final result.
-Our training script was adapted from the example provided in LlamaFactory (Zheng et al., 2024b)  [^1] https://github.com/hiyouga/LLaMA-Factory/blob/main/examples/train_full/llama3_full_sft_ds3.yaml.
-The training was configured with a batch size of 1 per device, gradient accumulation steps of 4, a learning rate of 1e-5, and he model was trained for 3 epochs, with warmup over 20 steps and a cosine learning rate scheduler.
-For generating diverse responses from LLMs, we followed the configuration in Yuan et al.
-(2024), setting $T=0.7$ and $p=0.9$.
-For tasks such as AlpacaEval (Dubois et al., 2024b), we used GPT-4-o unless stated otherwise.
+Our training script was adapted from the example provided in LlamaFactory  (Zheng et al.
+2024b).
+The training was configured with a batch size of 1 per device, gradient accumulation steps of 4, a learning rate of 1e-5, and the model was trained for 3 epochs, with warmup over 20 steps and a cosine learning rate scheduler.
+For generating diverse responses from LLMs, we followed the configuration in (Yuan et al.
+2024), setting $T=0.7$ and $p=0.9$.
+For tasks such as AlpacaEval (Dubois et al.
+2024b), we used GPT-4-o unless stated otherwise.
 A.2 Details of Evaluator Selection Across Different Tasks
@@
 Response Selection.
-In the response selection task, we evaluated both single models and ensembles of evaluators using the GED method.
+In the response selection task, we evaluated both single models and ensembles of evaluators using the PGEDmethod.
 For single model evaluation, we assessed the standalone performance of Llama3-8B, Mistral-7B, Qwen2-7B, and Qwen2-72B on benchmarks such as HumanEval, AlpacaEval, MATH, GSM8k, and GAIA.
-This served two purposes: first, to establish the baseline performance of smaller models (Llama3-8B, Mistral-7B, Qwen2-7B) that are later combined using GED, and second, to compare against Qwen2-72B, where we tested a baseline approach by randomly selecting one response from the 10 it generated for each question.
+This served two purposes: first, to establish the baseline performance of smaller models (Llama3-8B, Mistral-7B, Qwen2-7B) that are later combined using PGED, and second, to compare against Qwen2-72B, where we tested a baseline approach by randomly selecting one response from the 10 it generated for each question.
 For single evaluator evaluation, each model (Llama3-8B, Mistral-7B, Qwen2-7B, and Qwen2-72B) was also used as an evaluator to rank responses generated by Qwen2-72B.
 Notably, Qwen2-72B acted as a self-evaluator, selecting the best response from its own generated outputs.
-Finally, for multiple evaluator evaluation, our GED method combined the evaluations of Llama3-8B, Mistral-7B, and Qwen2-7B to rank responses.
+Finally, for multiple evaluator evaluation, our PGEDmethod combined the evaluations of Llama3-8B, Mistral-7B, and Qwen2-7B to rank responses.
 Model Ranking.
@@
 Larger models generally produce more reliable rankings, closely aligning with GPT-4, and the AlpacaEval benchmark, containing 805 tasks, makes the computational cost of using larger models acceptable.
-In the multiple evaluator setting, our GED method aggregated the evaluations from these four larger models to produce robust and consistent rankings.
+In the multiple evaluator setting, our PGEDmethod aggregated the evaluations from these four larger models to produce robust and consistent rankings.
 The combination of these high-capacity models ensures that our approach yields rankings that are both accurate and consistent across tasks.
@@
 These models were selected because they balance computational efficiency and evaluation performance, making them suitable for iterative instruction tuning processes.
-The evaluators’ pairwise preferences were aggregated using GED to identify the most appropriate responses for training.
+The evaluators’ pairwise preferences were aggregated using PGEDto identify the most appropriate responses for training.
 This approach ensured that the selected data pairs reflected a consensus among the evaluators while keeping computational costs manageable.
-A.3 Definition of Weak Evaluators
-In this work, ”weak evaluators” are defined by their tendency to produce noisy or inconsistent pairwise preferences, not by their overall model capacity.
-Even advanced models like GPT-4-o can generate preference graphs with cycles, indicating that preference inconsistency is a universal challenge, regardless of model scale.
-Evaluators such as Llama3-8B, Mistral-7B, Qwen2-7B, and even GPT-4-o are considered weak if their pairwise evaluations exhibit significant noise or conflicts.
-A.4 Evaluation Settings: Single Model vs. Single Evaluator
-The ”single model” setting evaluates each model’s (Llama3-8B, Mistral-7B, Qwen2-7B, Qwen2-72B) outputs directly on benchmarks like HumanEval, AlpacaEval, MATH, GSM8k, and GAIA, without selection or modification.
-In contrast, the ”single evaluator” setting uses these models to select the best response from ten candidates generated by Qwen2-72B, assessing their evaluation capability.
+A.3 Evaluation Settings: Single Model vs. Single Evaluator
+The "single model" setting evaluates each model’s (Llama3-8B, Mistral-7B, Qwen2-7B, Qwen2-72B) outputs directly on benchmarks like HumanEval, AlpacaEval, MATH, GSM8k, and GAIA, without selection or modification.
+In contrast, the "single evaluator" setting uses these models to select the best response from ten candidates generated by Qwen2-72B, assessing their evaluation capability.
 The key difference is that the single model setting focuses on generation quality, while the single evaluator setting assesses evaluation ability.
-A.5 Definition of Cycle Rate
+A.4 Definition of Cycle Rate
 The Cycle Rate is the percentage of preference graphs with at least one cycle, indicating inconsistency in pairwise comparisons.
@@
 This metric helps assess the reliability of evaluators, such as GPT-4-o and GPT-4-o-mini, across different datasets.
-A.6 Models Used for Ranking
+A.5 Models Used for Ranking
 We evaluate model ranking using 30 widely used models from the AlpacaEval dataset, covering diverse architectures and capabilities.
-Our model set $\mathcal{M}$ includes families such as Llama (Touvron et al., 2023), Vicuna (Zheng et al., 2023), GPT (Achiam et al., 2023), Claude (Anthropic, 2024), Qwen (Yang et al., 2024a), Mistral (Jiang et al., 2023), Yi (Young et al., 2024), and WizardLM (Xu et al., 2023), ensuring a broad assessment.
+Our model set $\mathcal{M}$ includes families such as Llama (Touvron et al.
+2023), Vicuna (Zheng et al.
+2023), GPT (Achiam et al.
+2023), Claude (Anthropic 2024), Qwen (Yang et al.
+2024), Mistral (Jiang et al.
+2023), Yi (Young et al.
+2024), and WizardLM (Xu et al.
+2023), ensuring a broad assessment.
 Proprietary models include OpenAI’s GPT series (e.g., gpt-3.5-turbo-0301, gpt-4o-2024-05-13) and Anthropic’s Claude models (e.g., claude-2, claude-3-opus-20240229), serving as strong baselines.
@@
 This diverse model set enables a comprehensive evaluation of ranking accuracy using AlpacaEval’s ground-truth rankings.
-A.7 Dataset
+A.6 Dataset
 In this appendix, we provide detailed information about the datasets used in main text.
-  • UltraFeedback (Cui et al., 2023): UltraFeedback is a large-scale, fine-grained, diverse preference dataset, used for training powerful reward models and critic models.
+  • UltraFeedback (Cui et al.
+2023): UltraFeedback is a large-scale, fine-grained, diverse preference dataset, used for training powerful reward models and critic models.
 We collect about 64k prompts from diverse resources (including UltraChat, ShareGPT, Evol-Instruct, TruthfulQA, FalseQA, and FLAN).
 We then use these prompts to query multiple LLMs (see Table for model lists) and generate 4 different responses for each prompt, resulting in a total of 256k samples.
-  • HH-RLHF (Bai et al., 2022): The HH-RLHF dataset contains human preference data for training language models to be helpful and harmless, as well as red teaming data to identify harmful model outputs.
+  • HH-RLHF (Bai et al.
+2022): The HH-RLHF dataset contains human preference data for training language models to be helpful and harmless, as well as red teaming data to identify harmful model outputs.
 The preference data includes pairs of chosen and rejected responses, while the red teaming data includes transcripts of adversarial interactions with AI assistants, rated for harmfulness.
-We strictly follow prior works (Song et al., 2024; Yu et al., 2023) and used the code from this repository  [^2] https://github.com/AlibabaResearch/DAMO-ConvAI/tree/main/PRO for testing.
-  • MATH (Hendrycks et al., 2021): The MATH dataset consists of 12,500 challenging competition-level math problems, each with a detailed step-by-step solution.
+We strictly follow prior works (Song et al.
+2024; Yu et al.
+2023) for testing.
+  • MATH (Hendrycks et al.
+2021): The MATH dataset consists of 12,500 challenging competition-level math problems, each with a detailed step-by-step solution.
 It is designed to teach models to generate answer derivations and explanations, aiding in mathematical reasoning.
@@
 Due to the high resource cost of using the full test set, we randomly sampled 400 problems from the test set for evaluation.
-  • GSM8k (Chen et al., 2021): GSM8K (Grade School Math 8K) is a collection of 8.5K high-quality math word problems designed for grade school students.
+  • GSM8k (Chen et al.
+2021): GSM8K (Grade School Math 8K) is a collection of 8.5K high-quality math word problems designed for grade school students.
 It supports the task of multi-step reasoning and question answering in basic math.
@@
 Due to the high computational cost of using the entire test set, we randomly sampled 400 data points from the test set for our evaluation.
-  • GAIA (Mialon et al., 2023): The GAIA dataset is a benchmark designed to evaluate next-generation LLMs with augmented capabilities like tooling and search access.
+  • GAIA (Mialon et al.
+2023): The GAIA dataset is a benchmark designed to evaluate next-generation LLMs with augmented capabilities like tooling and search access.
 It consists of over 450 complex questions with unambiguous answers, requiring various levels of autonomy and tooling.
@@
 We used the entire test set for our evaluation.
-  • HumanEval (Chen et al., 2021): The OpenAI HumanEval dataset contains 164 programming problems, each with a function signature, docstring, body, and unit tests.
+  • HumanEval (Chen et al.
+2021): The OpenAI HumanEval dataset contains 164 programming problems, each with a function signature, docstring, body, and unit tests.
 These problems are handwritten to ensure they were not included in the training sets of code generation models.
@@
   • AlpacaEval.
-(Dubois et al., 2024b): AlpacaEval consists of 805 instructions, including 252 from the self-instruct test set (Wang et al., 2022), 188 from the Open Assistant (OASST) test set, 129 from Anthropic’s helpful test set (Zhou et al., 2023), 80 from the Vicuna test set (Chiang et al., 2023), and 156 from the Koala test set (Vu et al., 2023).
+(Dubois et al.
+2024b): AlpacaEval consists of 805 instructions, including 252 from the self-instruct test set (Wang et al.
+2022), 188 from the Open Assistant (OASST) test set, 129 from Anthropic’s helpful test set (Zhou et al.
+2023), 80 from the Vicuna test set (Chiang et al.
+2023), and 156 from the Koala test set (Vu et al.
+2023).
   • LIMA.
-(Zhou et al., 2023): LIMA collects a training dataset of 1000 prompts and responses, curated to ensure stylistic consistency while accommodating diverse input types.
+(Zhou et al.
+2023): LIMA collects a training dataset of 1000 prompts and responses, curated to ensure stylistic consistency while accommodating diverse input types.
 It also includes an open-source test set of 300 prompts and a development set of 50.
-The data is primarily sourced from community-driven Q&A platforms like Stack Exchange, wikiHow, and the Pushshift Reddit Dataset (Baumgartner et al., 2020), along with manually curated examples.
+The data is primarily sourced from community-driven Q&A platforms like Stack Exchange, wikiHow, and the Pushshift Reddit Dataset (Baumgartner et al.
+2020), along with manually curated examples.
 The inclusion of human-authored examples further increases dataset diversity.
@@
   • Vicuna.
-(Chiang et al., 2023): Vicuna organizes its 80 test instructions into eight distinct categories: Fermi problems, commonsense, roleplaying, coding/math/writing tasks, counterfactuals, knowledge, and general questions.
+(Chiang et al.
+2023): Vicuna organizes its 80 test instructions into eight distinct categories: Fermi problems, commonsense, roleplaying, coding/math/writing tasks, counterfactuals, knowledge, and general questions.
 This categorization aims to comprehensively assess different facets of chatbot performance.
-Prior work suggests that Vicuna’s instructions are generally of lower complexity and difficulty (Xu et al., 2023).
+Prior work suggests that Vicuna’s instructions are generally of lower complexity and difficulty (Xu et al.
+2023).
 We utilize the Vicuna test set to assess the performance of large language models across these varied categories of instructions.
   • Self-Instruct.
-(Wang et al., 2022): Self-Instruct contains 252 human-authored test instructions, each paired with a well-constructed output.
+(Wang et al.
+2022): Self-Instruct contains 252 human-authored test instructions, each paired with a well-constructed output.
 This dataset is curated to simulate real-world use cases of instruction-following models, spanning various domains such as email composition, social media, productivity tools, and coding tasks.
@@
   • Wizardlm.
-(Xu et al., 2023): Wizardlm consists of a training set of 70k examples derived from 52k instructions initially provided by Alpaca.
+(Xu et al.
+2023): Wizardlm consists of a training set of 70k examples derived from 52k instructions initially provided by Alpaca.
 The test set includes 218 instructions sourced from various open-source projects and online communities, covering 29 distinct skills derived from real-world tasks.
@@
   • Koala.
-(Vu et al., 2023): Koala comprises 180 real-world user queries sourced from the web, spanning diverse topics and typically reflecting a conversational tone.
+(Vu et al.
+2023): Koala comprises 180 real-world user queries sourced from the web, spanning diverse topics and typically reflecting a conversational tone.
 These queries are especially relevant for evaluating models intended for chat-based applications.
@@
 We exclusively used the Koala test set to gauge our model’s proficiency in handling authentic conversational queries.
-A.8 Aggregation Process in GED Across Different Tasks
-The GED implementation differs between response selection and model ranking tasks due to their objectives.
-In response selection, GED aggregates preference graphs from multiple evaluators for each question into a single DAG.
+A.7 Aggregation Process in PGEDAcross Different Tasks
+The PGEDimplementation differs between response selection and model ranking tasks due to their objectives.
+In response selection, PGEDaggregates preference graphs from multiple evaluators for each question into a single DAG.
 From this DAG, a final ranking is derived, and the top-ranked answer is selected as the output.
 Notably, aggregation is performed only at the per-question level, without a rank aggregation across questions.
-In model ranking, GED involves two steps: first, aggregating the evaluators’ preference graphs into a DAG for each question to rank the models, and second, employing a rank ensemble method to aggregate these per-question rankings into a final overall model ranking across all questions.
-For details on the aggregation modeling, see Section 3.3, covering Response Selection, Model Ranking, and Model Alignment.
-Appendix B Alternative Evaluator Configurations for Model Ranking
-To address concerns about the computational cost and fairness of using 70B-level models as weak evaluators in the model ranking task, we conducted additional experiments using smaller, more comparable models.
-Table 5 summarizes the performance of these models, both individually and when combined using GED.
-The results show that GED outperforms individual evaluators even when using smaller 7B-scale models, achieving an average score of 62.70 compared to the best individual performance of 57.92 (Qwen2-7B with graph denoising).
-This demonstrates that the aggregation of smaller models through GED effectively enhances performance while reducing computational costs.
-These findings validate the versatility of GED, showing that it can provide robust and accurate rankings without relying solely on large-scale models.
+In model ranking, PGEDinvolves two steps: first, aggregating the evaluators’ preference graphs into a DAG for each question to rank the models, and second, employing a rank ensemble method to aggregate these per-question rankings into a final overall model ranking across all questions.
+For details on the aggregation modeling, see Section 4, covering Response Selection, Model Ranking, and Model Alignment.
+Appendix B Ablation study for Test Time scaling
+We evaluate the impact of removing the ensembling step in PGED, referred to as the (w/o ensemble) variant.
+In this case, individual evaluators’ preference graphs are denoised and converted to rankings, which are then aggregated using methods such as Weight Score, Kemeny, Weighted Kemeny, Pairwise Majority, and Weighted Pairwise Majority (detailed in Appendix Q).
+For simplicity of presentation, we use Weight Score to represent PGED(w/o ensemble) (Weight Score).
+As shown in Figure 5, all (w/o ensemble) methods consistently underperform compared to PGED.
+This performance gap arises because converting graphs to ranks before aggregation leads to information loss.
+In contrast, PGED ensembles the graphs directly, preserving more detailed preference information and resulting in better final rankings.
+Figure 5: Comparison of PGED and (w/o ensemble) variants.
+PGED outperforms due to preserving more information by directly ensembling preference graphs, while rank aggregation in the (w/o ensemble) methods leads to performance loss.
+Appendix C Limitation
+Our framework improves consistency in preference evaluations, but it implicitly promotes a "majority view" by aggregating judgments from multiple LLM-based evaluators.
+This may inadvertently suppress minority or dissenting preferences, especially in cases where subjectivity, cultural variation, or ethical nuance plays a critical role.
+While the goal is to reduce noise, there is a risk that important human disagreement is treated as inconsistency to be "denoised." In socially sensitive domains, such as education, healthcare, or content moderation, this could limit the diversity of acceptable outputs or reinforce dominant norms embedded in current LLMs.
+Table 4: Performance comparison on nuanced quality metrics (%).
+PGEDoutperforms individual evaluators and random selection across Factuality, Relevance, Coherence, Inform., Helpful.
+and Validity metrics.
+Method Fact. Rel. Coh. Info. Help. Valid. Avg.
+Random 86.73 87.91 92.47 77.62 17.48 48.92 68.52
+Llama3-8B 88.59 89.91 94.41 79.77 18.48 50.92 70.35
+Mistral-7B 89.10 90.29 94.85 79.95 18.55 51.13 70.65
+Qwen2-7B 89.25 90.44 95.03 80.09 18.58 51.21 70.77
+PGED 94.73 95.91 97.36 86.62 19.48 55.92 75.00
+Appendix D Evaluating PGED on More Metrics
+To assess PGEDon more metrics, we evaluated its impact on nuanced quality aspects of LLM outputs, including factuality, relevance, coherence, informativeness, helpfulness, and validity.
+Since LLM tasks often require subtle judgments, PGED’s adaptability is crucial.
+Following (Chen et al.
+2023a), we used Llama3-70B to generate ten candidate responses per query and applied PGEDfor response selection, with Llama3-8B, Mistral-7B, and Qwen2-7B as evaluators.
+Table 4 shows that PGEDconsistently outperformed individual evaluators and random selection across all metrics.
+It improved factuality by  5 % points over the best evaluator, while also enhancing relevance, coherence, informativeness, and helpfulness.
+These results highlight PGED’s ability to aggregate preferences effectively, capturing subtle qualities of generated content and improving overall reliability and utility in LLM assessments.
+Appendix E Alternative Evaluator Configurations for Model Ranking
+To address concerns about the computational cost and fairness of using 70B-level models as preference evaluators in the model ranking task, we conducted additional experiments using smaller, more comparable models.
+Table 5 summarizes the performance of these models, both individually and when combined using PGED.
+The results show that PGEDoutperforms individual evaluators even when using smaller 7B-scale models, achieving an average score of 62.70 compared to the best individual performance of 57.92 (Qwen2-7B with graph denoising).
+This demonstrates that the aggregation of smaller models through PGEDeffectively enhances performance while reducing computational costs.
+These findings validate the versatility of PGED, showing that it can provide robust and accurate rankings without relying solely on large-scale models.
 Table 5: Performance comparison in the model ranking task using 7B-scale models as evaluators on the AlpacaEval dataset.
 Higher values indicate better performance.
-GED achieves robust performance even with smaller evaluators.
+PGEDachieves robust performance even with smaller evaluators.
 Model Weight Score Kemeny
@@
 Single evaluator with graph denoising 50.47 55.06 54.92 61.39 61.21 56.61
-Multiple evaluator GED 57.59 61.14 61.14 67.17 66.46 62.70
-Appendix C Comparison with Point-Wise Scoring Methods
-To further investigate the difference between point-wise and preference-based scoring methods, we conducted additional experiments under the response selection setting.
-Specifically, we evaluated a point-wise baseline where evaluators assigned scores to responses on a scale of 1 to 5, selecting the highest-rated response.
-The results are shown in Table 6.
-Table 6: Comparison of point-wise methods and GED on response selection tasks.
-Model HumanEval AlpacaEval MATH GSM8k GAIA Avg
-Point-wise - Llama3-8B 60.97 28.77 74.06 82.65 10.97 51.48
-Point-wise - Mistral-7B 62.83 26.93 74.02 83.21 10.22 51.44
-Point-wise - Qwen2-7B 60.41 28.30 74.32 84.27 10.82 51.62
-Point-wise - Majority Voting 63.72 29.42 74.71 84.36 11.32 52.70
-GED (Llama3-8B, Mistral-7B, Qwen2-7B) 70.73 32.44 75.58 88.18 13.33 56.05
-From the results, we observe that point-wise methods consistently underperform compared to GED across all tasks.
-This can be attributed to the fact that point-wise scoring assesses each response independently, without considering the relative quality of different responses.
-As a result, it lacks the global information available in preference-based ranking, where responses are directly compared.
-While majority voting improves point-wise performance slightly by aggregating scores from multiple evaluators, it remains inferior to GED, which leverages pairwise preferences to construct a more informed global ranking.
-Despite its limitations, the point-wise approach has an advantage in terms of computational efficiency, as it requires fewer comparisons than pairwise methods.
-This trade-off suggests that the choice between point-wise and preference-based scoring should depend on the specific application: point-wise scoring may be preferable in large-scale settings where efficiency is critical, while GED is more effective for achieving higher-quality rankings.
-Appendix D Impact of Evaluator Quantity on Denoising Quality
-We investigated how the number of evaluators affects the denoising quality in GED by conducting experiments using different numbers of evaluators in the response selection setting.
-Specifically, we evaluated the performance of GED when aggregating preferences from two, three, and four evaluators.
-The evaluators used were Llama3-8B, Mistral-7B, Qwen2-7B, and Gemma-9B [^3] https://huggingface.co/google/gemma-2-9b-it.
-Table 7: Performance comparison of GED with varying numbers of evaluators across five benchmarks.
-Increasing the number of evaluators enhances denoising quality, reflected in improved performance metrics.
-Evaluators Set HumanEval AlpacaEval MATH GSM8k GAIA Avg.
-Llama3-8B, Mistral-7B 69.21 31.87 74.97 86.92 12.51 55.10
-Llama3-8B, Mistral-7B, Qwen2-7B 70.73 32.44 75.58 88.18 13.33 56.05
-Llama3-8B, Mistral-7B, Qwen2-7B, Gemma-9B 70.98 32.87 75.91 88.75 13.46 56.39
-As shown in Table 7, increasing the number of evaluators consistently improves the denoising quality of GED, as reflected by higher performance across all benchmarks.
-For instance, the average performance improves from 55.10% with two evaluators to 56.39% with four evaluators.
-This enhancement can be attributed to the diverse perspectives and complementary strengths of multiple evaluators, contributing to a more robust and accurate aggregation of preferences.
-These results highlight the importance of selecting a diverse and capable set of evaluators to enhance GED’s effectiveness.
-Incorporating more evaluators allows the denoising process to better identify and mitigate inconsistencies in the preference graphs, leading to improved overall performance in response selection tasks.
+Multiple evaluator PGED 57.59 61.14 61.14 67.17 66.46 62.70
 Algorithm 2 Construction of the Preference Graph for Model Ranking
@@
 27:end for
+Appendix F Comparison with Point-Wise Scoring Methods
+To further investigate the difference between point-wise and preference-based scoring methods, we conducted additional experiments under the response selection setting.
+Specifically, we evaluated a point-wise baseline where evaluators assigned scores to responses on a scale of 1 to 5, selecting the highest-rated response.
+The results are shown in Table 6.
+Table 6: Comparison of point-wise methods and PGEDon response selection tasks.
+Model HumanEval AlpacaEval MATH GSM8k GAIA Avg
+Point-wise - Llama3-8B 60.97 28.77 74.06 82.65 10.97 51.48
+Point-wise - Mistral-7B 62.83 26.93 74.02 83.21 10.22 51.44
+Point-wise - Qwen2-7B 60.41 28.30 74.32 84.27 10.82 51.62
+Point-wise - Majority Voting 63.72 29.42 74.71 84.36 11.32 52.70
+PGED(Llama3-8B, Mistral-7B, Qwen2-7B) 70.73 32.44 75.58 88.18 13.33 56.05
+From the results, we observe that point-wise methods consistently underperform compared to PGEDacross all tasks.
+This can be attributed to the fact that point-wise scoring assesses each response independently, without considering the relative quality of different responses.
+As a result, it lacks the global information available in preference-based ranking, where responses are directly compared.
+While majority voting improves point-wise performance slightly by aggregating scores from multiple evaluators, it remains inferior to PGED, which leverages pairwise preferences to construct a more informed global ranking.
+Despite its limitations, the point-wise approach has an advantage in terms of computational efficiency, as it requires fewer comparisons than pairwise methods.
+This trade-off suggests that the choice between point-wise and preference-based scoring should depend on the specific application: point-wise scoring may be preferable in large-scale settings where efficiency is critical, while PGEDis more effective for achieving higher-quality rankings.
+Appendix G Impact of Evaluator Quantity on Denoising Quality
+We investigated how the number of evaluators affects the denoising quality in PGEDby conducting experiments using different numbers of evaluators in the response selection setting.
+Specifically, we evaluated the performance of PGEDwhen aggregating preferences from two, three, and four evaluators.
+The evaluators used were Llama3-8B, Mistral-7B, Qwen2-7B, and Gemma-9B.
+Table 7: Performance comparison of PGEDwith varying numbers of evaluators across five benchmarks.
+Increasing the number of evaluators enhances denoising quality, reflected in improved performance metrics.
+Evaluators Set HumanEval AlpacaEval MATH GSM8k GAIA Avg.
+Llama3-8B, Mistral-7B 69.21 31.87 74.97 86.92 12.51 55.10
+Llama3-8B, Mistral-7B, Qwen2-7B 70.73 32.44 75.58 88.18 13.33 56.05
+Llama3-8B, Mistral-7B, Qwen2-7B, Gemma-9B 70.98 32.87 75.91 88.75 13.46 56.39
+As shown in Table 7, increasing the number of evaluators consistently improves the denoising quality of PGED, as reflected by higher performance across all benchmarks.
+For instance, the average performance improves from 55.10% with two evaluators to 56.39% with four evaluators.
+This enhancement can be attributed to the diverse perspectives and complementary strengths of multiple evaluators, contributing to a more robust and accurate aggregation of preferences.
+These results highlight the importance of selecting a diverse and capable set of evaluators to enhance PGED’s effectiveness.
+Incorporating more evaluators allows the denoising process to better identify and mitigate inconsistencies in the preference graphs, leading to improved overall performance in response selection tasks.
+Appendix H Construction of the Preference Graph
+In this section, we provide a detailed explanation of the process used to construct the preference graph sets for both the model ranking and response selection tasks, as outlined in Algorithms 2 and 3.
+These algorithms form the backbone of our method, enabling the representation of pairwise preferences as directed graphs, which are essential for downstream aggregation and ranking.
+Algorithm 2 describes the construction process for generating a set of preference graphs for the model ranking task.
+The procedure is as follows: Initialization: For each question $q\in Q$, we begin by initializing a vertex set $V_{q}$, where each vertex $v_{i}$ corresponds to a model $\mathcal{M}_{i}$ in the set of models $M$.
+We also initialize an empty set of edges $A_{a}$ and a weight function $w_{a}$, which will be used to track the strength of the preferences between model pairs.
+Pairwise Comparisons: For each pair of models $\mathcal{M}_{i}$ and $\mathcal{M}_{j}$, the assigned evaluator $a\in\mathcal{A}$ assesses their responses to the given question $q$.
+If evaluator $a$ prefers $\mathcal{M}_{i}$ over $\mathcal{M}_{j}$, a directed edge $(v_{i}\rightarrow v_{j})$ is added to the edge set $A_{a}$, and its corresponding weight is incremented.
+Conversely, if $\mathcal{M}_{j}$ is preferred, the edge $(v_{j}\rightarrow v_{i})$ is added or its weight is updated.
+Graph Storage: Once all pairwise comparisons have been processed for a given evaluator, the resulting graph $G_{a}=(V_{q},A_{a},w_{a})$ is stored for that evaluator.
+This process is repeated for all evaluators in $\mathcal{A}$ and for all questions in $Q$, generating a set of preference graphs for each evaluator and question.
+Algorithm 3 follows a similar structure but applies to the response selection task, where the objective is to rank a set of candidate answers for each question: Initialization: For each question $q\in Q$, we initialize a vertex set $V_{q}$, where each vertex corresponds to a candidate answer $ans_{i}$.
+As in the model ranking task, we also initialize an edge set $A_{a}$ and a weight function $w_{a}$ for each evaluator $a\in\mathcal{A}$.
+Pairwise Comparisons: Evaluators compare the quality of pairs of candidate answers $ans_{i}$ and $ans_{j}$ for each question.
+A directed edge is added based on the evaluator’s preference, with the weight reflecting the strength of preference.
+As before, if evaluator $a$ prefers $ans_{i}$ over $ans_{j}$, an edge $(v_{i}\rightarrow v_{j})$ is added or its weight incremented, and vice versa.
+Graph Storage: After all pairwise comparisons are complete, the preference graph $G_{a}=(V_{q},A_{a},w_{a})$ for evaluator $a$ is stored.
+This procedure is repeated for all evaluators and questions, resulting in a set of preference graphs for each evaluator and each question.
+Both algorithms ensure that the preference graphs are constructed in a consistent manner, forming the basis for the aggregation and denoising processes used later in our framework.
+These graphs encapsulate the evaluators’ preferences and provide a structured representation of pairwise comparisons, facilitating further analysis.
 Algorithm 3 Construction of the Preference Graph for Response Selection
@@
 end for
-Appendix E Construction of the Preference Graph
-In this section, we provide a detailed explanation of the process used to construct the preference graph sets for both the model ranking and response selection tasks, as outlined in Algorithms 1 and 2.
-These algorithms form the backbone of our method, enabling the representation of pairwise preferences as directed graphs, which are essential for downstream aggregation and ranking.
-Algorithm 1 describes the construction process for generating a set of preference graphs for the model ranking task.
-The procedure is as follows: Initialization: For each question $q\in Q$, we begin by initializing a vertex set $V_{q}$, where each vertex $v_{i}$ corresponds to a model $\mathcal{M}_{i}$ in the set of models $M$.
-We also initialize an empty set of edges $A_{a}$ and a weight function $w_{a}$, which will be used to track the strength of the preferences between model pairs.
-Pairwise Comparisons: For each pair of models $\mathcal{M}_{i}$ and $\mathcal{M}_{j}$, the assigned evaluator $a\in\mathcal{A}$ assesses their responses to the given question $q$.
-If evaluator $a$ prefers $\mathcal{M}_{i}$ over $\mathcal{M}_{j}$, a directed edge $(v_{i}\rightarrow v_{j})$ is added to the edge set $A_{a}$, and its corresponding weight is incremented.
-Conversely, if $\mathcal{M}_{j}$ is preferred, the edge $(v_{j}\rightarrow v_{i})$ is added or its weight is updated.
-Graph Storage: Once all pairwise comparisons have been processed for a given evaluator, the resulting graph $G_{a}=(V_{q},A_{a},w_{a})$ is stored for that evaluator.
-This process is repeated for all evaluators in $\mathcal{A}$ and for all questions in $Q$, generating a set of preference graphs for each evaluator and question.
-Algorithm 2 follows a similar structure but applies to the response selection task, where the objective is to rank a set of candidate answers for each question: Initialization: For each question $q\in Q$, we initialize a vertex set $V_{q}$, where each vertex corresponds to a candidate answer $ans_{i}$.
-As in the model ranking task, we also initialize an edge set $A_{a}$ and a weight function $w_{a}$ for each evaluator $a\in\mathcal{A}$.
-Pairwise Comparisons: Evaluators compare the quality of pairs of candidate answers $ans_{i}$ and $ans_{j}$ for each question.
-A directed edge is added based on the evaluator’s preference, with the weight reflecting the strength of preference.
-As before, if evaluator $a$ prefers $ans_{i}$ over $ans_{j}$, an edge $(v_{i}\rightarrow v_{j})$ is added or its weight incremented, and vice versa.
-Graph Storage: After all pairwise comparisons are complete, the preference graph $G_{a}=(V_{q},A_{a},w_{a})$ for evaluator $a$ is stored.
-This procedure is repeated for all evaluators and questions, resulting in a set of preference graphs for each evaluator and each question.
-Both algorithms ensure that the preference graphs are constructed in a consistent manner, forming the basis for the aggregation and denoising processes used later in our framework.
-These graphs encapsulate the evaluators’ preferences and provide a structured representation of pairwise comparisons, facilitating further analysis.
-Appendix F The Importance of Addressing Cyclic Inconsistency
+Appendix I The Importance of Addressing Cyclic Inconsistency
 Cyclic inconsistencies in preference data introduce contradictions that undermine meaningful ranking and evaluation.
-Many applications, including alignment optimization (Ouyang et al., 2022; Song et al., 2024), recommendation systems (Hou et al., 2024), and model evaluation (Liu et al., 2024), rely on transitive preferences for stability and interpretability.
-For example, OpenAI’s RLHF pipelines (Ouyang et al., 2022) construct training data based on transitive preferences (e.g., A $\succ$ B $\succ$ C), which ensures consistency in model learning.
+Many applications, including alignment optimization (Ouyang et al.
+2022; Song et al.
+2024), recommendation systems (Hou et al.
+2024), and model evaluation (Liu et al.
+2024), rely on transitive preferences for stability and interpretability.
+For example, OpenAI’s RLHF pipelines (Ouyang et al.
+2022) construct training data based on transitive preferences (e.g., A $\succ$ B $\succ$ C), which ensures consistency in model learning.
 However, if preference cycles exist (e.g., A $\succ$ B $\succ$ C $\succ$ A), the resulting contradictions lead to instability and degrade training performance.
@@
 Ensuring acyclicity enhances dataset reliability and improves model effectiveness in ranking and evaluation tasks.
-Appendix G Proof of Theorem 1
+Appendix J Proof of Theorem 1
 See 1
-Proof.
-For brevity, we consider all edges in $G$ have weights equal to 1 and all weights in $\widehat{G}$ are divided by $N$.
-By construction, we can notice that for each $(u\rightarrow v)\in A$, the weight $w_{\widehat{G}}(v\rightarrow u)$ can be viewed as an empirical estimate of $\delta_{1}$.
-Then, we claim that the following two events can imply $G\subseteq\mathrm{MAS}(\widehat{G})$:
+Proof J.1
+[…]
   • $(\mathcal{E}_{1})$ For any $(u\rightarrow v)\in A$, it holds $\left|w_{\widehat{G}}(v\rightarrow u)-\delta_{1}\right|\leq\frac{\epsilon}{2}$.
   • $(\mathcal{E}_{2})$ For any pair of nodes $(u,v)$ such that $(u\rightarrow v),(v\rightarrow u)\notin A$, it holds $\left|w_{\widehat{G}}(u\rightarrow v)-w_{\widehat{G}}(v\rightarrow u)\right|\leq\frac{\epsilon}{U}$.
-To see this, first by Lemma 1, we know that for any pair of nodes $(u,v)$, $\mathrm{MAS}(\widehat{G})$ will contain exactly one of $(u\rightarrow v)$ and $(v\rightarrow u)$.
-[^4] There is non-zero probability that some edges in $\widehat{G}$ will have zero weight, but we treat them as existing for the ease of argument.
-That is, we allow only $\widehat{G}$ to contain zero-weight edges.
-Therefore, for any $(u\rightarrow v)\in A$, $\mathrm{MAS}(\widehat{G})$ will contain exactly one of $(u\rightarrow v)$ and $(v\rightarrow u)$.
-Then, since $\mathcal{E}_{1}$ holds, $\delta_{1}=0.5-\epsilon<0.5$ and $w_{\widehat{G}}(u\rightarrow v)+w_{\widehat{G}}(v\rightarrow u)=1$, we have $w_{\widehat{G}}(u\rightarrow v)-w_{\widehat{G}}(v\rightarrow u)\geq\epsilon$.
-Furthermore, since $\mathcal{E}_{2}$ holds, for $(u,v)$ such that $(u\rightarrow v),(v\rightarrow u)\notin A$, arbitrary way of edge removing among these nodes can influence the total edge weights by at most $\epsilon$.
-Therefore, when applying the denoising operation to $\widehat{G}$, for any $(u\rightarrow v)\in A$, only $(u\rightarrow v)$ will be kept in $\mathrm{MAS}(\widehat{G})$, which makes $G\subseteq\mathrm{MAS}(\widehat{G})$.
-As a result, we have $\mathbb{P}(\mathcal{E}_{1}\cap\mathcal{E}_{2})\leq\mathbb{P}\left(G\subseteq\mathrm{MAS}(\widehat{G})\right)$.
+[…]
 Then, we can now bound the probability of $\mathcal{E}_{1}\cap\mathcal{E}_{2}$.
@@
 Therefore, by Bernstein’s inequality, we have
-$\mathbb{P}\left(\left|w_{\widehat{G}}(u\rightarrow v)-w_{\widehat{G}}(v\rightarrow u)\right|\leq\frac{\epsilon}{U}\right)\geq$ $1-2\exp\left(-\frac{N\epsilon^{2}}{6U^{2}\delta_{2}+2U\epsilon}\right)$
-$\implies\mathbb{P}(\mathcal{E}_{2})\geq$ $1-2U\exp\left(-\frac{N\epsilon^{2}}{6U^{2}\delta_{2}+2U\epsilon}\right),$
+$\mathbb{P}\left(\left|w_{\widehat{G}}(u\rightarrow v)-w_{\widehat{G}}(v\rightarrow u)\right|\leq\frac{\epsilon}{U}\right)$
+$\geq 1-2\exp\left(-\frac{N\epsilon^{2}}{6U^{2}\delta_{2}+2U\epsilon}\right)$
+$\implies\mathbb{P}(\mathcal{E}_{2})$ $\geq 1-2U\exp\left(-\frac{N\epsilon^{2}}{6U^{2}\delta_{2}+2U\epsilon}\right),$
 where the second inequality is an union bound over all unconnected node pairs in $G$. As a result, we eventually have
-$\mathbb{P}\left(G\subseteq\mathrm{MAS}(\widehat{G})\right)\geq$ $\mathbb{P}\left(\mathcal{E}_{1}\cap\mathcal{E}_{2}\right)\geq 1-2|A|\exp\left(-\frac{N\epsilon^{2}}{2}\right)-2U\exp\left(-\frac{N\epsilon^{2}}{6U^{2}\delta_{2}+2U\epsilon}\right).$
-Lemma 1.
+$\mathbb{P}\left(G\subseteq\mathrm{MAS}(\widehat{G})\right)\geq\mathbb{P}\left(\mathcal{E}_{1}\cap\mathcal{E}_{2}\right)$
+$\geq 1-2|A|\exp\left(-\frac{N\epsilon^{2}}{2}\right)-2U\exp\left(-\frac{N\epsilon^{2}}{6U^{2}\delta_{2}+2U\epsilon}\right).$
+Lemma 1
 For a weighted directed graph $G=(V,A,w)$, if $(u\rightarrow v),(v\rightarrow u)\in A$, then $\mathrm{MAS}(G)$ contains exactly one of $(u\rightarrow v)$ and $(v\rightarrow u)$.
-Proof.
+Proof J.2
 Recall that $\mathrm{MAS}(G)$ gives an acyclic subgraph of $G$ with the maximum weight.
@@
 Therefore, $\mathrm{MAS}(G)$ must contain exactly one of $(u\rightarrow v)$ and $(v\rightarrow u)$.
-∎
-Appendix H Impact of Evaluator Weighting on GED Performance
-In our theoretical analysis, GED assumes equal weighting of edges in the preference graphs.
+Appendix K Impact of Evaluator Weighting on PGED
+In our theoretical analysis, PGEDassumes equal weighting of edges in the preference graphs.
 However, in practical scenarios, evaluators may have varying levels of reliability or expertise.
-Incorporating evaluator-specific confidence scores or performance metrics could enhance the effectiveness of GED.
-To investigate this, we conducted experiments using a weighted version of GED, referred to as WeightGED, under the Response Selection setting.
+Incorporating evaluator-specific confidence scores or performance metrics could enhance the effectiveness of PGED.
+To investigate this, we conducted experiments using a weighted version of PGED, referred to as WeightGED, under the Response Selection setting.
 In WeightGED, we assigned weights to evaluators based on their individual performance on specific datasets.
@@
 These accuracies were normalized to compute evaluator weights:
-[…]
-[…]
-[…]
+$\text{weight}(\text{Llama3-8B})$ $=\frac{62.19}{62.19+67.24+61.58}=0.326,$
+$\text{weight}(\text{Mistral-7B})$ $=\frac{67.24}{62.19+67.24+61.58}=0.352,$
+$\text{weight}(\text{Qwen2-7B})$ $=\frac{61.58}{62.19+67.24+61.58}=0.322.$
 These weights were used to scale the contributions of each evaluator’s preferences during graph construction.
-We compared the performance of GED and WeightGED across multiple benchmarks, as presented in Table 8.
-Table 8: Performance comparison of GED and WeightGED using 7B-scale evaluators across five benchmarks.
-WeightGED marginally outperforms GED, demonstrating the potential benefits of incorporating evaluator-specific weights.
+We compared the performance of PGEDand WeightGED across multiple benchmarks, as presented in Table 8.
+Table 8: Performance comparison of PGEDand WeightGED using 7B-scale evaluators across five benchmarks.
+WeightGED marginally outperforms PGED, demonstrating the potential benefits of incorporating evaluator-specific weights.
 Method HumanEval AlpacaEval MATH GSM8k GAIA Avg.
-GED 70.73 32.44 75.58 88.18 13.33 56.05
+PGED 70.73 32.44 75.58 88.18 13.33 56.05
 7B Evaluators WeightGED 70.97 32.67 75.71 88.24 13.56 56.23
-GED 73.21 59.87 82.49 86.43 16.27 63.65
+PGED 73.21 59.87 82.49 86.43 16.27 63.65
 GPT Evaluators WeightGED 74.52 61.71 83.93 87.84 17.32 65.06
-As shown in Table 8, WeightGED achieves marginal but consistent improvements over GED across all benchmarks.
+As shown in Table 8, WeightGED achieves marginal but consistent improvements over PGEDacross all benchmarks.
 For instance, using 7B-scale evaluators, WeightGED improves the average performance from 56.05% to 56.23%.
 Similarly, with GPT evaluators, the average performance increases from 63.65% to 65.06%.
-These results suggest that incorporating evaluator-specific weights based on performance metrics can enhance the effectiveness of GED.
+These results suggest that incorporating evaluator-specific weights based on performance metrics can enhance the effectiveness of PGED.
 Furthermore, we conducted additional experiments using stronger evaluators such as GPT-3.5, GPT-4-o-mini, and GPT-4-o.
 The weights were computed based on their respective performance accuracies, following the same normalization procedure.
-The improvements observed with these evaluators reinforce the potential of weighting schemes in enhancing GED.
+The improvements observed with these evaluators reinforce the potential of weighting schemes in enhancing PGED.
 In summary, our preliminary findings indicate that integrating evaluator-specific confidence scores or performance metrics is a promising direction for future work.
 Systematically designing and optimizing these weighting schemes could lead to more robust and accurate evaluation frameworks.
-Appendix I Expanded Related Work on Preference Denoising for LLMs
+Appendix L Expanded Related Work on Preference Denoising for LLMs
 This work situates itself within the broader field of denoising preference evaluations for LLMs, addressing inconsistencies and noise in preference graphs.
 We acknowledge that existing literature has explored two primary approaches to this challenge: within-model preference modeling and modular pre-processing.
-Below, we provide a detailed comparison of GED with representative methods from these approaches.
-Within-model approaches, such as robust DPO (rDPO) (Chowdhury et al., 2024) and conservative DPO (cDPO) (Mitchell, 2023), focus on integrating denoising mechanisms directly within the preference modeling process.
+Below, we provide a detailed comparison of PGED with representative methods from these approaches.
+Within-model approaches, such as robust DPO (rDPO) (Chowdhury, Kini, and Natarajan 2024) and conservative DPO (cDPO) (Mitchell 2023), focus on integrating denoising mechanisms directly within the preference modeling process.
 These methods incorporate regularization techniques to mitigate the effects of noisy or adversarial preference data during model alignment.
 While these approaches are effective in refining the preference modeling pipeline, they are tightly coupled with specific models and tasks, limiting their versatility.
-In contrast, GED is a modular framework that operates as a pre-processing step, denoising preference graphs before downstream tasks, making it adaptable to a wide range of applications and models.
-Modular approaches like CURATRON (Naresh et al., 2024) address noise and missing comparisons in preference datasets using techniques such as low-rank matrix completion.
-While CURATRON effectively mitigates certain types of noise, it does not explicitly target cyclic inconsistencies (e.g., A $\succ$ B, B $\succ$ C, C $\succ$ A), which are a critical focus of GED.
-By leveraging a graph-based framework, GED detects and removes such cycles through its denoising process, ensuring that the resulting preference graph is acyclic and thus more consistent and reliable for downstream use.
-Additionally, GED distinguishes itself by providing theoretical guarantees for recovering the ground truth preference DAG under certain conditions.
-Furthermore, the ensemble mechanism in GED demonstrates the novel insight that combining small evaluators can surpass the performance of a single stronger evaluator, a feature not emphasized in the aforementioned methods.
-Appendix J Impact of Evaluator Selection on GED Performance
-The selection of weak evaluators is a critical factor influencing the performance of GED.
-While our initial experiments demonstrated the benefits of combining weak evaluators, further analysis is necessary to understand how their diversity and capabilities affect GED’s effectiveness.
-To address this, we conducted additional experiments evaluating three different configurations of evaluator sets: the original set comprising Llama3-8B, Mistral-7B, and Qwen2-7B; a diverse set that adds Gemma-9B [^5] https://huggingface.co/google/gemma-2-9b-it to introduce more variation in model architecture and training data; and a higher-capability set that replaces smaller models with larger ones, specifically Llama3-70B, Mistral-8×7B, and Qwen2-72B.
+In contrast, PGED is a modular framework that operates as a pre-processing step, denoising preference graphs before downstream tasks, making it adaptable to a wide range of applications and models.
+Modular approaches like CURATRON (Naresh, Tulabandhula et al.
+2024) address noise and missing comparisons in preference datasets using techniques such as low-rank matrix completion.
+While CURATRON effectively mitigates certain types of noise, it does not explicitly target cyclic inconsistencies (e.g., A $\succ$ B, B $\succ$ C, C $\succ$ A), which are a critical focus of PGED.
+By leveraging a graph-based framework, PGED detects and removes such cycles through its denoising process, ensuring that the resulting preference graph is acyclic and thus more consistent and reliable for downstream use.
+Additionally, PGED distinguishes itself by providing theoretical guarantees for recovering the ground truth preference DAG under certain conditions.
+Furthermore, the ensemble mechanism in PGED demonstrates the novel insight that combining small evaluators can surpass the performance of a single stronger evaluator, a feature not emphasized in the aforementioned methods.
+Appendix M Theoretical Analysis of Graph Denoising
+Let $G=(V,A,w)$ be the aggregated digraph, with $w:A\to\mathbb{R}_{\geq 0}$. Write $m\triangleq|\{(u,v)\in A:w(u,v)>0\}|$. For $u\in V$, define
+$d^{+}(u)$ $\triangleq\sum_{(u,v)\in A}w(u,v),d^{-}(u)\triangleq\sum_{(v,u)\in A}w(v,u),$
+$d(u)$ $\triangleq d^{+}(u)+d^{-}(u),\delta(u)\triangleq d^{+}(u)-d^{-}(u).$
+Let $s=s_{1}\|s_{2}$ be the order returned by Algorithm 1, and
+$R(s)$ $\triangleq\{(v_{j}\to v_{i})\in A:\ j>i\text{ in }s\},$
+$G^{\prime}$ $\triangleq(V,A\setminus R(s),w).$
+Proposition 1 (Acyclicity)
+$G^{\prime}$ is a directed acyclic graph.
+Proof M.1
+We prove by contradiction.
+Assume $G^{\prime}$ contains a directed cycle $C=(u_{1}\!\to u_{2}\!\to\cdots\!\to u_{\ell}\!\to u_{1})$.
+Since $G^{\prime}$ keeps exactly the forward arcs w.r.t.
+the total order $s$, every arc $(u_{t}\!\to u_{t+1})$ in $C$ must satisfy $\mathrm{pos}_{s}(u_{t})<\mathrm{pos}_{s}(u_{t+1})$, where $\mathrm{pos}_{s}(\cdot)$ is the index in $s$.
+Summing these strict inequalities around the cycle yields
+$\mathrm{pos}_{s}(u_{1})<\mathrm{pos}_{s}(u_{2})<\cdots<\mathrm{pos}_{s}(u_{\ell})<\mathrm{pos}_{s}(u_{1}),$
+which is impossible. Therefore $G^{\prime}$ has no directed cycle.
+Assumption 1 (Bounded integral weights)
+There exists a fixed integer $k\geq 1$ (the number of evaluators) such that $w(u,v)\in\{0,1,\dots,k\}$ for all $(u,v)\in A$.
+Assumption 1 holds in our ensemble setting because each evaluator contributes at most one unit to any arc weight.
+Data structures.
+We maintain: (i) a queue $Q^{-}$ of sources ($d^{-}(u)=0$); (ii) a queue $Q^{+}$ of sinks ($d^{+}(u)=0$); (iii) $\delta$-buckets $\{B_{\Delta}\}_{\Delta=-M}^{M}$, where $M\triangleq\max_{u\in V}d(u)$.
+Each bucket is a doubly linked list of vertices with current $\delta(u)=\Delta$ and $d^{\pm}(u)>0$.
+We also keep an integer pointer $p$ storing the maximum index $\Delta$ with $B_{\Delta}\neq\emptyset$.
+All vertices are in exactly one of $^{-}, Q+Q^{+}$, or some $B_{\Delta}$.
+Lemma 2 (Range and update magnitude)
+Under Assumption 1: (a) $M=\mathcal{O}(m)$; (b) when removing a vertex $u$, for any neighbor $v$ we have
+$\delta_{\text{new}}(v)$ $=\delta_{\text{old}}(v)+w(u,v)-w(v,u),$
+$\Rightarrow\quad|\delta_{\text{new}}(v)-\delta_{\text{old}}(v)|$ $\leq k.$
+Proof M.2
+(a) Since each incident arc contributes at most $k$ to $d^{\pm}(\cdot)$,
+$d(u)=\sum_{(u,x)}w(u,x)+\sum_{(x,u)}w(x,u)\;\leq\;k\cdot\deg(u),$
+hence $M=\max_{u}d(u)\leq k\cdot\max_{u}\deg(u)\leq k\cdot m=\mathcal{O}(m)$.
+(b) When $u$ is removed, exactly the arcs $(u,v)$ and $(v,u)$ (if present) are deleted among $v$’s incidences.
+Thus $d^{-}(v)$ decreases by $w(u,v)$, $d^{+}(v)$ decreases by $w(v,u)$, and $\delta(v)=d^{+}(v)-d^{-}(v)$ increases by $w(u,v)$ and decreases by $w(v,u)$, as claimed.
+By Assumption 1, $w(u,v),w(v,u)\leq k$, hence the $|\cdot|$ bound.
+Lemma 3 (Bucket maintenance in $\mathcal{O}(1)$ amortized time)
+Under Assumption 1 , the following hold.
+  1. 1.
+    Initialization. Computing $d^{\pm}(\cdot)$ and placing vertices into $^{-}, Q+Q^{+}$, or $B_{\Delta}$ takes $\mathcal{O}(m)$ time and $\mathcal{O}(m)$ space.
+  2. 2.
+    Updates.
+Removing a vertex $u$ and processing all its incident arcs take time $\mathcal{O}(1+\deg(u))$: every neighbor $v$ is updated once, and $v$ moves from $B_{\Delta}$ to $B_{\Delta^{\prime}}$ with $|\Delta^{\prime}-\Delta|\leq k$; enqueuing/dequeuing from $Q^{\pm}$ or $B_{\Delta}$ is $\mathcal{O}(1)$ via doubly linked lists.
+  3. 3.
+    Selecting $\max\delta$. Maintaining pointer $p$ to the largest index with $B_{p}\neq\emptyset$ costs $\mathcal{O}(1)$ amortized per removal.
+Proof M.3
+(1) We make one pass over adjacency lists: for $(x,y)$ add $w(x,y)$ to $d^{+}(x)$ and to $d^{-}(y)$.
+This is $\mathcal{O}(m)$.
+Classification into $Q^{\pm}$ or some $B_{\Delta}$ is then $\mathcal{O}(n)$, dominated by $\mathcal{O}(m)$ since a connected digraph has $m\geq n-1$.
+(2) When removing $u$, we unlink $u$ from its current container in $\mathcal{O}(1)$.
+For each incident arc touching $u$, we visit the unique neighbor $v$ once, adjust $d^{\pm}(v)$ by Lemma 2, recompute $\delta(v)$, and move $v$:
+[…]
+Each move is pointer manipulation in a doubly linked list, costing $\mathcal{O}(1)$.
+Hence the work per removal is $\mathcal{O}(1+\deg(u))$ and each arc is processed exactly once (when the first endpoint is removed).
+(3) Amortized bound for $p$.
+Let $p_{t}$ be the pointer after the $t$-th removal.
+There are two kinds of motion:
+Downward motion.
+Whenever $B_{p_{t}}$ becomes empty, we decrement $p_{t}$ until finding the next nonempty bucket.
+Since $p$ never moves below M-M and never increases during this downward scan, the total number of decrements over the entire execution is at most $2M=\mathcal{O}(m)$.
+Upward motion. Deleting $u$ can increase some $\delta(v)$, but by Lemma 2 the new maximum bucket index can exceed the previous maximum by at most $k$:
+$\max_{x\in V}\delta_{t}(x)$ $\leq\max\Bigl\{\max_{x\in V}\delta_{t-1}(x),\ \max_{v\in N(u)}\delta_{t-1}(v)+k\Bigr\}$
+$\leq p_{t-1}+k.$
+We therefore scan upward by at most $k$ to reestablish $p_{t}$.
+Since $k$ is a fixed constant and there are $|V|$ removals, the total upward scans cost $\mathcal{O}(k|V|)=\mathcal{O}(|V|)$.
+For connected digraphs, $|V|\leq m+1$, hence $\mathcal{O}(|V|)=\mathcal{O}(m)$.
+Combining both motions gives $\mathcal{O}(1)$ amortized per removal.
+Theorem 2 (Linear time and space)
+Under Assumption 1, Algorithm 1 runs in $\mathcal{O}(m)$ time and uses $\mathcal{O}(m)$ space.
+Proof M.4
+Time.
+Initialization is $\mathcal{O}(m)$ by Lemma 3(1).
+Across the execution, each vertex is removed once.
+By Lemma 3(2), the cost of removing $u$ is $\mathcal{O}(1+\deg(u))$.
+Summing over all $u$ and noting $\sum_{u}\deg(u)=2m$, we obtain
+$\sum_{u\in V}\mathcal{O}(1+\deg(u))$ $=\mathcal{O}\bigl(|V|+\sum_{u}\deg(u)\bigr)$
+$=\mathcal{O}(|V|+m)$
+$=\mathcal{O}(m),$
+since a connected digraph has $|V|\leq m+1$.
+The amortized $\mathcal{O}(1)$ selection of the maximum-$\delta$ bucket follows from Lemma 3(3), and is already charged in the removal step.
+Therefore the overall running time is $\mathcal{O}(m)$.
+Space.
+Adjacency lists and degree/weight arrays take $\mathcal{O}(m)$.
+The queues $Q^{\pm}$ store disjoint subsets of $V$; the buckets $\{B_{\Delta}\}$ occupy an array of size $2M+1=\mathcal{O}(m)$ with total stored vertices $|V|$.
+All pointers and auxiliary counters are $\mathcal{O}(1)$ each.
+Hence the working memory is $\mathcal{O}(m)$.
+Remark 1 (Arbitrary real weights)
+If $w$ is real-valued without a constant bound on per-arc increments, the bucket update in Lemma 3 no longer yields $\mathcal{O}(1)$.
+Replacing buckets by a binary max-heap over current $\delta(u)$ preserves correctness and yields $\mathcal{O}(m\log n)$ time and $\mathcal{O}(m)$ space.
+Appendix N Impact of Evaluator Selection on PGED
+The selection of preference evaluators is a critical factor influencing the performance of PGED.
+While our initial experiments demonstrated the benefits of combining preference evaluators, further analysis is necessary to understand how their diversity and capabilities affect PGED’s effectiveness.
+To address this, we conducted additional experiments evaluating three different configurations of evaluator sets: the original set comprising Llama3-8B, Mistral-7B, and Qwen2-7B; a diverse set that adds Gemma-9B to introduce more variation in model architecture and training data; and a higher-capability set that replaces smaller models with larger ones, specifically Llama3-70B, Mistral-8×7B, and Qwen2-72B.
 The results, shown in Table 9, reveal two key insights.
-First, incorporating diversity by adding Gemma-9B improves GED’s performance slightly, increasing the average score from 56.05% to 56.39%.
+First, incorporating diversity by adding Gemma-9B improves PGED’s performance slightly, increasing the average score from 56.05% to 56.39%.
 This suggests that models with diverse training data and architectures can contribute to more robust aggregated evaluations.
@@
 Diversity provides marginal gains by bringing varied perspectives, while higher-capacity models contribute to more significant improvements by enhancing the overall quality of evaluations.
-This suggests that, when computational resources allow, incorporating advanced models into the evaluator set can meaningfully boost GED’s performance.
-Table 9: Performance comparison with different evaluator sets across five benchmarks. The results highlight the impact of evaluator diversity and capability on GED’s effectiveness.
+This suggests that, when computational resources allow, incorporating advanced models into the evaluator set can meaningfully boost PGED’s performance.
+Table 9: Performance comparison with different evaluator sets across five benchmarks. The results highlight the impact of evaluator diversity and capability on PGED’s effectiveness.
 Evaluators Set HumanEval AlpacaEval MATH GSM8k GAIA Avg
@@
 26:return $G^{*}=(V,A^{*},w^{*})$
-Appendix K Cost Considerations and Active Learning with ActiveGED
+Appendix O Cost Considerations and Active Learning with ActiveGED
 Using multiple evaluators and aggregating their preferences can be computationally expensive, especially when constructing dense preference graphs.
@@
 The results, presented in Table 10, demonstrate that ActiveGED achieves competitive performance with significantly fewer evaluations.
-For example, under a 50% budget, ActiveGED retains most of the performance benefits of full GED while cutting the number of pairwise comparisons in half.
+For example, under a 50% budget, ActiveGED retains most of the performance benefits of full PGEDwhile cutting the number of pairwise comparisons in half.
 Table 10: Performance comparison of ActiveGED under different budget constraints. ActiveGED achieves competitive performance with significantly fewer pairwise evaluations.
-Evaluators Set of GED HumanEval AlpacaEval MATH GSM8k GAIA Avg.
+Evaluators Set of PGED HumanEval AlpacaEval MATH GSM8k GAIA Avg.
 Random 57.93 29.58 72.75 84.67 11.52 51.29
@@
 ActiveGED (50%) 68.62 31.91 74.87 87.06 12.08 54.91
-GED (Full Budget) 70.73 32.44 75.58 88.18 13.33 56.05
+PGED(Full Budget) 70.73 32.44 75.58 88.18 13.33 56.05
 The algorithm behind ActiveGED is outlined in Algorithm 4.
@@
 ActiveGED demonstrates that by carefully selecting the most informative pairs, it is possible to achieve competitive performance with significantly fewer evaluations.
-This makes GED more scalable and practical for real-world applications where computational resources are limited.
-Appendix L Mitigating Evaluator Biases in GED
-Evaluator biases, such as position bias and token bias, can significantly impact the accuracy and fairness of the preference graphs used in GED.
+This makes PGEDmore scalable and practical for real-world applications where computational resources are limited.
+Appendix P Mitigating Evaluator Biases in PGED
+Evaluator biases, such as position bias and token bias, can significantly impact the accuracy and fairness of the preference graphs used in PGED.
 Addressing these biases is crucial for ensuring reliable evaluations and robust performance.
@@
 This approach minimizes the impact of position bias on the overall rankings.
-Token bias occurs when an evaluator exhibits a latent preference for a specific option or label (e.g., consistently favoring ”Option A” over ”Option B”).
+Token bias occurs when an evaluator exhibits a latent preference for a specific option or label (e.g., consistently favoring "Option A" over "Option B").
 Our framework implicitly mitigates token bias through the aforementioned position-swapping strategy, which prevents evaluators from associating a fixed label with a specific position.
 By averaging the results across both orderings, any systematic preference for a particular label is effectively neutralized.
-Appendix M Denoising of Preference Graphs for GED
-In this section, we describe the denoising procedure used in the GED framework, specifically for transforming the aggregated preference graph into a DAG.
-Let $G=(V,A)$ be a simple connected directed graph, where $n=|V|$ and $m=|A|$, with $n-1\leq m\leq\binom{n}{2}$.
-A feedback arc set (FAS) of $G$, denoted as $R(G)$, is a set of arcs whose removal transforms $G$ into a DAG.
-The minimum feedback arc set $R^{*}(G)$ is the FAS of minimum cardinality, and finding $R^{*}(G)$ is the well-known FAS problem.
-Consider a scenario where the vertices of $G$ are arranged sequentially along a horizontal line and labeled as $v_{1},v_{2},\ldots,v_{n}$ from left to right.
-This arrangement is referred to as a vertex sequence and is denoted by $s=v_{1}v_{2}\ldots v_{n}$.
-Each vertex sequence $s$ induces a feedback arc set $R(s)$, consisting of all leftward arcs $v_{j}\rightarrow v_{i}$ for $j>i$.
-The FAS problem can therefore be reformulated as finding a vertex sequence $s^{*}$ such that $R(s^{*})=R^{*}(G)$.
-Our proposed algorithm, GED, computes a vertex sequence $s$ that corresponds to a minimal feedback arc set $R(s)$.
-The algorithm iteratively removes vertices (and their incident arcs) from $G$, focusing on sinks, sources, and vertices that maximize a specific property.
-For any vertex $u\in V$, let $d(u)$ denote its degree, $d^{+}(u)$ its outdegree, and $d^{-}(u)$ its indegree, such that $d(u)=d^{+}(u)+d^{-}(u)$.
-At each step, after removing sinks and sources, the algorithm selects a vertex $u$ for which $\delta(u)=d^{+}(u)-d^{-}(u)$ is maximized.
-If the removed vertex $u$ is a sink, it is concatenated with a vertex sequence $s_{2}$; otherwise, it is concatenated with $s_{1}$.
-Once $G$ is reduced to an empty graph, the final vertex sequence $s$ is obtained by concatenating $s_{1}$ and $s_{2}$.
-The detailed steps of the algorithm are shown in Algorithm 4.
-In our GED framework, this denoising step is essential for ensuring that the aggregated preference graph becomes acyclic, thus enabling a reliable ranking to be extracted.
-By iteratively removing vertices based on their structural properties, we minimize the feedback arc set and ensure that the remaining graph is a DAG, which can be directly used to generate rankings in subsequent steps.
-Algorithm 4 Preference Graphs Denoising for GED
-1:$G$: Directed graph, var $s$: Vertex sequence
-2:$s_{1}\leftarrow\emptyset$
-3:$s_{2}\leftarrow\emptyset$
-4:while $G\neq\emptyset$ do
-5:     while $G$ contains a sink do
-6:         Choose a sink $u$
-7:         $s_{2}\leftarrow u\cdot s_{2}$
-8:         $G\leftarrow G-u$
-9:     end while
-10:     while $G$ contains a source do
-11:         Choose a source $u$
-12:         $s_{1}\leftarrow s_{1}\cdot u$
-13:         $G\leftarrow G-u$
-14:     end while
-15:     Choose a vertex $u$ with maximum $\delta(u)$
-16:     $s_{1}\leftarrow s_{1}\cdot u$
-17:     $G\leftarrow G-u$
-18:end while
-19:$s\leftarrow s_{1}\cdot s_{2}$
-Appendix N Rank ensemble method
-Weight score (Adler et al., 2002)
+Appendix Q Rank ensemble method
+Weight score (Adler, Friedman, and Sinuany-Stern 2002)
 : The Weight Score method assigns a score to each vertex $v$ based on its position in each ranking $\mathcal{R}_{i}$. For a vertex $v$ in ranking $\mathcal{R}_{i}$, the score is given by:
@@
 The final consensus ranking $\mathcal{R}^{*}$ is obtained by sorting the vertices $v$ in descending order of $T(v)$.
-Kemeny and weighted Kemeny (Kemeny, 1959)
+Kemeny and weighted Kemeny (Kemeny 1959)
 : The Kemeny method seeks a consensus ranking $\mathcal{R}^{*}$ that minimizes the total pairwise disagreements between $\mathcal{R}^{*}$ and the input rankings, measured using the Kendall $\tau$-distance:
@@
 Here, the goal is to minimize the weighted Kendall tau distance, emphasizing rankings with higher weights.
-Pairwise majority and weighted pairwise majority (Caragiannis et al., 2016)
-: The Pairwise Majority (PM) method determines a consensus ranking $\mathcal{R}^{*}$ by maximizing the number of pairwise agreements with the input rankings.
+Pairwise majority and weighted pairwise majority
+(Caragiannis, Procaccia, and Shah 2016): The Pairwise Majority (PM) method determines a consensus ranking $\mathcal{R}^{*}$ by maximizing the number of pairwise agreements with the input rankings.
 For each pair of vertices $(v_{i},v_{j})$, the goal is to ensure that the majority of rankings agree with their relative order in $\mathcal{R}^{*}$:
-$\mathcal{R}^{*}=\arg\max_{\mathcal{R}}\sum_{i<j}\left(\sum_{p=1}^{k}\mathds{1}(\mathcal{R}_{p}(v_{i})<\mathcal{R}_{p}(v_{j}))\right)\cdot\mathds{1}(\mathcal{R}(v_{i})<\mathcal{R}(v_{j}))$ (5)
+$\mathcal{R}^{*}=\arg\max_{\mathcal{R}}\sum_{i<j}\left(\sum_{p=1}^{k}\mathds{1}(\mathcal{R}_{p}(v_{i})<\mathcal{R}_{p}(v_{j}))\right)\cdot\mathds{1}(\mathcal{R}(v_{i})<\mathcal{R}(v_{j}))$
 The Weighted Pairwise Majority method incorporates weights $\alpha_{p}$ to account for the reliability of each ranking $\mathcal{R}_{p}$:
-$\mathcal{R}^{*}=\arg\max_{\mathcal{R}}\sum_{i<j}\left(\sum_{p=1}^{k}\alpha_{p}\cdot\mathds{1}(\mathcal{R}_{p}(v_{i})<\mathcal{R}_{p}(v_{j}))\right)\cdot\mathds{1}(\mathcal{R}(v_{i})<\mathcal{R}(v_{j}))$ (6)
+$\mathcal{R}^{*}=\arg\max_{\mathcal{R}}\sum_{i<j}\left(\sum_{p=1}^{k}\alpha_{p}\cdot\mathds{1}(\mathcal{R}_{p}(v_{i})<\mathcal{R}_{p}(v_{j}))\right)\cdot\mathds{1}(\mathcal{R}(v_{i})<\mathcal{R}(v_{j}))$
 In both methods, the objective is to maximize the (weighted) pairwise agreement between the consensus ranking and the input rankings.
-Appendix O Visualization
-(a) Raw Preference Graph 1
-(b) Denoised Preference Graph 1
-(c) Raw Preference Graph 2
-(d) Denoised Preference Graph 2
-Figure 5: Case studies showcasing the raw and denoised preference graphs.
-In both Case 1 and Case 2, the raw preference graphs (a, c) contain cyclic inconsistencies, which are resolved by GED into directed acyclic graphs (b, d).
-The dashed lines in the denoised graphs represent the edges that were removed by GED to eliminate cycles.
-The nodes labeled 0-9 correspond to the ten generated responses by Qwen2-72B.
-These examples illustrate the effectiveness of GED in eliminating noise and restoring consistency in preference evaluations.
-In this section, we present two case studies that demonstrate the effectiveness of our proposed GED method in denoising preference graphs.
-Figure 5 illustrates the raw preference graphs, which are generated by multiple evaluators, i.e., Llama3-8B, Mistral-7B, and Qwen2-7B, through the responses produced by Qwen2-72B on the HumanEval benchmark.
-The nodes labeled 0-9 in the graphs correspond to the ten generated responses.
-The comparison between the raw (left) and denoised (right) graphs shows how our method successfully resolves cyclic inconsistencies, transforming noisy graphs into DAGs.
-In the denoised graphs, the dashed lines represent the edges that were removed by GED to eliminate cycles.
-Case Study 1.
-In the first case study, we showcase the impact of our denoising approach.
-The raw graph (Figure 5 (a)) contains multiple cyclic inconsistencies, such as $9\succ 4\succ 3\succ 9$, $9\succ 4\succ 3\succ 2\succ 9$, and $9\succ 4\succ 3\succ 7\succ 9$.
-By applying GED, we identify that removing a single edge (4 $\succ$ 3) eliminates all cycles, converting the noisy preference graph into a consistent DAG, as shown in Figure 5 (b).
-Case Study 2.
-The second case study (Figure 5 (c)) presents another scenario with conflicting preferences, where cycles like $7\succ 0\succ 6\succ 7$ and $7\succ 1\succ 0\succ 6\succ 7$ indicate noisy judgments.
-Here, removing just one edge (0 $\succ$ 6) using GED is sufficient to eliminate all cycles and convert the graph into a DAG, as depicted in Figure 5 (d).
-Conclusion from Case Studies.
-In both case studies, removing any other edge would not fully resolve all cyclic inconsistencies without requiring additional deletions, which would result in more information loss.
-GED effectively minimizes edge removals while maintaining the integrity of the original preference graph, making it a highly efficient solution for improving the consistency of preference evaluations.
-Appendix P Prompt Template
+Appendix R Cost-Performance Analysis of PGED
+We conducted a detailed cost-performance analysis based on publicly available API pricing from Amazon Bedrock.
+API Cost Comparison: For our PGED method, we utilize three relatively lightweight evaluators: Llama3-8B, Mistral-7B, and Qwen2-7B.
+Since pricing information for Qwen2 is not currently available on Amazon Bedrock, we conservatively adopt the pricing of Llama3 as a proxy.
+Specifically, the cost per 1 million input tokens is \$0.09, and the cost per 1 million output tokens is \$0.175.
+In contrast, using a single strong evaluator such as Qwen2-72B is significantly more expensive, with costs of \$0.23 per 1 million input tokens and \$0.40 per 1 million output tokens.
+This comparison shows that our PGED method can achieve superior evaluation performance at only approximately 39.13% of the input cost and 43.75% of the output cost.
+Performance vs.
+Budget Trade-off: As demonstrated in Appendix O, PGED is capable of outperforming strong evaluators like Qwen2-72B even when utilizing only 30–50% of the evaluation budget.
+This efficiency arises from effectively aggregating and denoising multiple weaker signals.
+The computational overhead involved in graph fusion and denoising steps within PGED is minimal compared to the inference time of large language models, making PGED a cost-effective and practical solution for scalable preference evaluations.
+Table 11: Performance comparison of PGEDwith multi-evaluator baselines PRD and ChatEval. PGEDachieves the highest score on each dataset and the best overall average.
+Method HumanEval AlpacaEval MATH GSM8k GAIA Avg.
+PRD 66.53 30.33 74.89 86.75 11.83 54.07
+ChatEval 66.42 29.71 74.62 86.66 11.53 53.79
+PGED 70.73 32.44 75.58 88.18 13.33 56.05
+Appendix S Comparison with More Baselines
+We extend our experimental analysis to include two recent multi-evaluator baselines: PRD (Li, Patel, and Du 2023b) and ChatEval (Chan et al.
+2023b).
+Both methods are designed to improve evaluation reliability by aggregating feedback from multiple evaluators.
+PRD focuses on pairwise ranking distillation, while ChatEval employs evaluator voting mechanisms.
+For a fair comparison, we evaluate all methods using the same set of models: Llama3-8B, Mistral-7B, and Qwen2-7B.
+Table 11 summarizes the results across five standard benchmark datasets.
+PGEDconsistently outperforms both PRD and ChatEval across all tasks.
+These results highlight the effectiveness of PGED’s structural approach.
+Unlike debate-style or voting-based methods, PGEDconstructs and denoises preference graphs to enforce global consistency, resulting in both improved empirical performance and stronger theoretical guarantees.
+Appendix T PGED with Strong Evaluators
+To examine whether PGED can benefit from stronger evaluators, we conducted additional experiments using three high-capacity language models: GPT-4-o-mini, GPT-4-o, and Gemini 2.0.
+These evaluators were used to construct individual and ensemble preference graphs under the PGED framework.
+The results, presented in Figure 6, demonstrate clear improvements of PGED over each individual evaluator across multiple tasks.
+This confirms that PGED not only performs well with lightweight models but also enhances the evaluation quality when applied to strong LLMs such as GPT-4 and Gemini.
+These findings support the generality and scalability of our framework.
+By aggregating and denoising even high-quality but noisy judgments, PGED consistently improves robustness and overall evaluation accuracy.
+Figure 6: Performance of PGED using strong evaluators: GPT-4-o-mini, GPT-4-o, and Gemini 2.0. PGED ensemble improves upon each individual model across tasks.
+Appendix U Comparison with Weighted Voting Baselines
+We further evaluate PGED against two intuitive multi-evaluator aggregation baselines: Weighted Majority Voting and Weighted Average of Ratings.
+Both baselines utilize the same set of evaluators (Llama3-8B, Mistral-7B, and Qwen2-7B) and apply weight-based schemes to either majority votes or numerical scores.
+Weighted Majority Voting assigns each evaluator a weight and determines the preferred candidate based on weighted voting outcomes across all pairwise comparisons.
+Weighted Average of Ratings aggregates normalized point-wise scores by evaluator weights to rank responses.
+As shown in Figure 7, PGED consistently outperforms both baselines across all evaluated tasks.
+This demonstrates that leveraging the structure of preference graphs and enforcing global consistency through denoising leads to more accurate and robust evaluations compared to simple weighted heuristics.
+Figure 7: Performance comparison of GED with Weighted Majority Voting and Weighted Average of Ratings. GED outperforms both across tasks.
+Appendix V Prompt Template
 Prompt for Response Selection.
-In this section, we provide detailed prompt templates used for response selection across five datasets: HumanEval (Chen et al., 2021), AlpacaEval (Dubois et al., 2024b), MATH (Hendrycks et al., 2021), GSM8k (Chen et al., 2021), and GAIA (Mialon et al., 2023).
+In this section, we provide detailed prompt templates used for response selection across five datasets: HumanEval (Chen et al.
+2021), AlpacaEval (Dubois et al.
+2024b), MATH (Hendrycks et al.
+2021), GSM8k (Chen et al.
+2021), and GAIA (Mialon et al.
+2023).
 These prompts include both a system prompt to establish the evaluator’s context and a user prompt tailored to the specific task requirements.
@@
 The templates are shown in Table 12, Table 13, Table 14, Table 15, and Table 16.
+Prompt for Model Ranking.
+This section presents the prompt template used in the Model Ranking task.
+The template is designed to evaluate and compare responses generated by different models for a given instruction, based on criteria such as accuracy, clarity, completeness, and helpfulness.
+The evaluation process involves analyzing two candidate responses and identifying which one better fulfills the requirements of the instruction.
+The detailed prompt for the Model Ranking is provided in Table 17.
+Prompt for Instruct Tuning.
+In this section, we provide the prompt template used for data selection in the Instruct Tuning task.
+The goal is to select the most appropriate response for each instruction from multiple candidates, ensuring that the selected responses are helpful, harmless, and relevant.
+We use the HH-RLHF dataset (Bai et al.
+2022), which contains human preference data aimed at training language models to be both helpful and harmless.
+The detailed prompt used by evaluators to assess and compare candidate responses is presented in Table 18.
 Table 12: Prompt template for evaluating programming solutions on the HumanEval dataset.
@@
 Provide a detailed comparison focusing on the criteria mentioned above.
-Prompt for Model Ranking.
-This section presents the prompt template used in the Model Ranking task.
-The template is designed to evaluate and compare responses generated by different models for a given instruction, based on criteria such as accuracy, clarity, completeness, and helpfulness.
-The evaluation process involves analyzing two candidate responses and identifying which one better fulfills the requirements of the instruction.
-The detailed prompt for the Model Ranking is provided in Table 16.
 Table 17: Prompt template in the Model Ranking.
@@
 Provide a detailed comparison focusing on the criteria mentioned above.
-Prompt for Instruct Tuning.
-In this section, we provide the prompt template used for data selection in the Instruct Tuning task.
-The goal is to select the most appropriate response for each instruction from multiple candidates, ensuring that the selected responses are helpful, harmless, and relevant.
-We use the HH-RLHF dataset (Bai et al., 2022), which contains human preference data aimed at training language models to be both helpful and harmless.
-The detailed prompt used by evaluators to assess and compare candidate responses is presented in Table 17.
 Table 18: Prompt template for evaluating responses in the Instruct Tuning.